### v1

# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

Abstract
Explainable AI (XAI) refers to techniques that provide human-understandable insights into the workings of AI models. Recently, the focus of XAI is being extended towards Large Language Models (LLMs) which are often criticized for their lack of transparency. This extension calls for a significant transformation in XAI methodologies because of two reasons. First, many existing XAI methods cannot be directly applied to LLMs due to their complexity and advanced capabilities. Second, as LLMs are increasingly deployed across diverse industry applications, the role of XAI shifts from merely opening the “black box” to actively enhancing the productivity and applicability of LLMs in real-world settings. Meanwhile, unlike traditional machine learning models that are passive recipients of XAI insights, the distinct abilities of LLMs can reciprocally enhance XAI. Therefore, in this paper, we introduce Usable XAI in the context of LLMs by analyzing (1) how XAI can benefit LLMs and AI systems, and (2) how LLMs can contribute to the advancement of XAI. We introduce 10 strategies, introducing the key techniques for each and discussing their associated challenges. We also provide case studies to demonstrate how to obtain and leverage explanations. The code used in this paper can be found at: https://github.com/JacksonWuxs/UsableXAI_LLM.
**footnotetext: Equal contribution
1 Introduction
Explainability holds great promise in understanding machine learning models and providing directions for improvement. In practice, users have high expectations for model explainability:
1. Through explanation, can we know if a model works properly?
2. Does explainability help developing better models?
First, explanations are expected to illuminate whether a model operates in accordance with human expectations. For example, does the model leverage reliable evidence and domain knowledge in its decision-making? Does the model contain bias and discrimination? Does the model show any vulnerabilities to potential attacks? Will the model output harmful information? Second, in recognition of model imperfections, we aspire for explainability to inform the development of better models. For example, how to adjust the behaviors of a model if we find it is using unreliable or unreasonable features in making predictions? Can we improve the performance of a model by aligning its behavior with human preferences?
Therefore, the question arises: Have these expectations been met? In recent years, the body of literature on Explainable AI (XAI) has expanded rapidly to improve model transparency (Du et al., 2019a; Murdoch et al., 2019; Tjoa & Guan, 2020; Došilović et al., 2018; Rudin et al., 2022), encompassing a wide array of methods customized for different data modalities, including visual (Zhang & Zhu, 2018), textual (Danilevsky et al., 2020), graph (Yuan et al., 2022), and time-series data (Zhao et al., 2023c). Some literature delves into specific techniques, such as attention methods, generalized additive models, and causal models. Additionally, some offer reviews on general principles and categorizations or initiate discussions on evaluating the faithfulness of explanations (Yang et al., 2019). Despite the progress, the last mile of XAI – making use of explanations – has not received enough attention. In many cases, we seem to be satisfied with just acquiring explanations and their associated visualizations, sometimes followed by qualitative analysis of the model’s strengths and weaknesses. While these explanations can reveal a model’s imperfections, quantifying model properties (e.g., fairness, security, rationality) or taking the next concrete steps toward model improvement remains a difficult task.
The challenges in achieving usable explainability are twofold. First, there is an inherent conflict between AI automation and human engagement in XAI. On one hand, humans need to define explainability that the model should follow, or scrutinize explanation to identify if any vulnerabilities exist in the model. On the other hand, the requirement for human oversight introduces substantial costs, posing challenges to the scalability and practical implementation of model debugging and improvement in AI workflows. Second, many of the current approaches view explainability as a purely technical matter, ignoring the needs of practitioners and non-technical stakeholders. Existing XAI methods are mainly developed as statistical and mathematical tools. However, there exists a noticeable disparity between the objectives of these tools and the expectations of practitioners across various application domains (Malizia & Paternò, 2023). An explanation that satisfies a technical audience might offer little value to a non-technical audience.
While the opacity issues have not yet been fully resolved for traditional deep models (e.g., multi-layer perceptrons, convolutional and recurrent neural network), the recent advancements of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023b; Chiang et al., 2023) appear to have exacerbated the challenge we are facing. Firstly, LLMs typically possess a significantly larger model size and a greater number of parameters. This increased model complexity intensifies the difficulty of explaining their inner workings. Second, different from traditional ML models that primarily focus on low-level pattern recognition tasks such as classification and parsing, LLMs can handle more complex tasks such as generation, reasoning and question answering. Understanding the exclusive abilities of LLMs presents novel challenges for XAI techniques. Considering the transformative impact of LLMs across various applications, ensuring the explainability and ethical use of LLMs has become an imminent and pressing need. Meanwhile, the emergent capabilities of LLMs also present new opportunities for XAI research. Their human-like communication and commonsense reasoning skills offer prospects for achieving explainability in ways that could potentially augment or replace human involvement.
Figure 1: The contributions and outline of this paper. We define Usable XAI in the context of LLMs with seven strategies of enhancing LLMs with XAI, and three strategies of enhancing XAI with LLMs.
Defining “Usable XAI”. In light of the above considerations, in the context of LLMs, we define Usable XAI which includes two aspects as follows. (1) Utilizing Explainability to Enhance LLM and AI Systems. Beyond just producing explanations or enhancing the transparency of LLMs, we explore whether these explanations can pinpoint issues for model debugging or improve the overall performance of LLMs or AI models at large, such as accuracy, controllability, fairness, and truthfulness. (2) Utilizing LLMs to Enhance XAI Frameworks. The human-like communication ability of LLMs can enhance model explanations in terms of user-friendliness, by converting the numerical values into understandable language. Also, the commonsense knowledge stored in LLMs can significantly boost the practicality of existing XAI frameworks, by playing the role of humans and alleviating the need for real human involvement in AI workflows.
Contribution of this paper. In this paper, we investigate 10 strategies towards usable XAI techniques in the context of LLMs. These strategies are organized into two major categories: (1) Usable XAI for LLMs; (2) LLM for Usable XAI, as shown in Figure 1. Additionally, we conduct case studies to substantiate the discussion on selected techniques. For each strategy, we also explore the open challenges and areas that require further investigation in future work.
  • Usable XAI for LLMs. We introduce how interpretation can be utilized to enhance AI pipelines, including LLMs and small models. First, we investigate how explanations could be utilized to diagnose and enhance LLMs in terms of utility. We study three types of post-hoc explanation methods, targeting LLM predictions (Section 2), LLM components (Section 3), and training samples (Section 4), respectively. Second, we focus on how explanations could be leveraged to scrutinize and boost model trustworthiness (Section 5), including security, fairness, toxicity, and truthfulness, which is crucial to achieving human alignment. Third, we discuss how explainability could guide the augmentation of data, including both inference data (i.e., prompts) and training data. Specifically, we discuss two strategies of crafting explainable prompts for LLMs: Chain-of-Thought prompts (Section 6) and knowledge-enhanced prompts (Section 7). Furthermore, we introduce leveraging LLM explanations to augment training data for improving small models (Section 8).
  • LLM for Usable XAI. In this part, we investigate strategies for leveraging the advanced capabilities of LLMs to address the challenges in traditional XAI domains, thus enhancing the usability of XAI in practice. First, we examine ways to enhance the user-friendliness of explanations through the generative capabilities of LLMs (Section 9). Second, we introduce how to automate the design of interpretable AI workflows by leveraging the planning abilities of LLMs (Section 10). Third, we introduce how to facilitate the evaluation of XAI methods by utilizing the unique property of LLMs in emulating human cognition processes (Section 11).
Differences between this paper and existing surveys. Many surveys have been conducted to examine Explainable AI (Du et al., 2019a; Tjoa & Guan, 2020; Došilović et al., 2018) or Interpretable Machine Learning (Murdoch et al., 2019). This paper differs from existing work as we focus on explanation methods for large language models. Meanwhile, different from the existing survey (Zhao et al., 2023b) that mainly reviews explanation methods for LLMs, our paper puts an emphasis on the XAI usability in LLM studies. To the best of our knowledge, the most related paper to our survey is (Luo & Specia, 2024), which also discusses several aspects where explanations can improve LLM performance. Nevertheless, this light-weight investigation lacks a thorough examination of XAI methods (e.g., sample-based explanation, interpretable workflows, explainable prompts) and how LLMs can benefit existing XAI frameworks (e.g., data augmentation, improving user-friendliness, XAI evaluation). Finally, our paper contributes further by providing detailed case studies and open-sourced codes, fostering future research in applying explanations effectively within the LLM context.
2 LLM Diagnosis via Attribution Methods
This section introduces attribution methods as post-hoc explanations for LLMs, and how we can discover model defects with attribution scores. We start with revisiting existing attribution methods, and then discuss which methods are still suitable for explaining LLMs. Since LLMs widely serve both classification and generation tasks, our discussion categorizes the attribution methods accordingly. After that, we explore case studies of applying attribution methods to assess LLM-generated output quality. Finally, we discuss future work of designing novel post-hoc explanation methods for LLMs.
2.1 Literature Review of Attribution Methods
The attribution-based explanation quantifies the importance of each input feature that contributes to making predictions. Given a language model $f$ with a prediction $\hat{y}=f(x)$ according to the $N$-words input prompt $x$, the explainer $g$ assesses the influence of input words in $x$ as $\mathbf{a}=g(x,\hat{y},f)\in\mathbb{R}^{N}$. Typically, the sign of $a_{n}\in\mathbf{a}$ indicates word $x_{n}$ positively or negatively influences $\hat{y}$, and a greater value of $|a_{n}|$ indicates a stronger impact. In text classification, $\hat{y}$ denotes a specific class label. In text generation, $\hat{y}$ represents a varying length of generated text.
Many existing attribution-based explanation methods focus on classification tasks and cannot be directly applied to the generation task. The primary distinction between them is that: classification is limited to a specific set of predictions, while generation encompasses an endless array of possibilities. For instance, in sentiment analysis, a language model can be instructed to output a number between 0 and 1 that indicates the positivity of input text by adding a linear layer and a sigmoid function on top of the language model. However, in the generative setting, the model can express this positivity in numerous expressions, such as “the reviewer definitely loves this movie” and “it is a strong positive movie review”. This distinction poses a unique challenge in adapting explanation methods from classification to generation tasks. In the following, we review related works based on the scenarios they are applicable to.
2.1.1 Attributing Inputs for Label Classification
Common attribution methods (Du et al., 2019a; Murdoch et al., 2019) developed for traditional deep models include gradient-based methods, perturbation-based methods, surrogate methods, and decomposition methods. We introduce the general idea and representative examples for each category, followed by the analysis of their suitability for explaining large language models.
Perturbation-based Explanation. Perturbation-based methods assess the importance of input features by perturbing them and monitoring changes in prediction confidence, i.e., $a_{n}=p(\hat{y}|x)-p(\hat{y}|\widetilde{x}_{n})$, where $\widetilde{x}_{n}$ refers to the input sequence with the $n$-th feature being perturbed. Each feature could refer to a word (Li et al., 2016a), a phrase (Wu et al., 2020b), or a word embedding (Li et al., 2016b). The underlying principle is that perturbing a more important feature should result in a more pronounced alteration in the model’s prediction confidence. However, this method has limitations, particularly in its assumption that features are independent, which is not always the case with textual data due to word inter-dependencies. Additionally, it is computationally intensive for explaining LLMs, requiring $N$ inferences for an input of $N$ words.
Gradient-based Explanation. Gradient-based methods offer a computationally efficient approach for estimating model sensitivity to input features based on gradients $\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}$, where $\mathbf{x}_{n}$ refers to the embedding of word $x_{n}$. Some methods employ the $L_{2}$-norm of gradients to assess word importance (Li et al., 2016a), i.e., $a_{n}=\|\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}\|_{2}$. This approach only requires a single inference and one backpropagation pass. Some extended methods multiply the gradient with the word embedding (Kindermans et al., ; Ebrahimi et al., 2018; Mohebbi et al., 2021), i.e., $a_{n}=\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}\cdot\mathbf{x}_{n}$. These methods may yield explanations with limited faithfulness for deep models (Shrikumar et al., 2017), as gradients only reflect the local relationship between input variation and output variation. To address this, Integrated Gradients (IG) has been proposed (Sundararajan et al., 2017; Sikdar et al., 2021; Sanyal & Ren, 2021; Enguehard, 2023), which accumulates gradients as input transitions from a reference point to the actual data point. Nevertheless, IG entails multiple rounds of inference and backpropagation, thus significantly increasing computational demands.
Surrogate-based Explanation. Surrogate-based explanation methods understand complex models by constructing a simpler model $g$ trained on $\mathcal{D}(x,\hat{y})=\{(\widetilde{x}_{k},\widetilde{y}_{k})\}_{k=1}^{K}$, where $\mathcal{D}(x,\hat{y})$ denotes a dataset constructed for the target instance $(x,\hat{y})$; $\widetilde{x}_{k}$ is usually obtained by perturbing $x$, and $\widetilde{y}_{k}=f(\widetilde{x}_{k})$. The surrogate model $g$, ranging from basic linear models to sophisticated decision trees, serves as a proxy to approximate the decision boundary of the target model $f$ for a specific instance $(x,\hat{y})$. Notable examples include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017), and TransSHAP (Kokalj et al., 2021), where the first two are designed for general deep neural networks and the last one is tailored for Transformer-based language models. Nevertheless, a significant limitation of them is their intensive reliance on repeated interactions with the target model, a process that is impractical for LLMs.
Decomposition-based Explanation. Decomposition-based methods assign linearly additive relevance scores to inputs, effectively breaking down the model’s prediction. Layer-wise Relevance Propagation (Montavon et al., 2019) and Taylor-type Decomposition (Montavon et al., 2017) are well-known techniques for computing these relevance scores. These methods have been adapted for Transformer-based language models in various research (Voita et al., 2019; 2020; Wu & Ong, 2021). However, a primary challenge in implementing decomposition-based explanations is the need for tailored decomposition strategies to accommodate different model architectures. Although many large language models are based on the Transformer framework, there are key variations between them, such as LLaMA (Touvron et al., 2023a) and GPT (OpenAI, 2023), particularly in aspects like positional encoding strategy and feed-forward network design. This challenge poses a limitation on the universal applicability of decomposition methods for general-purpose interpretation.
Table 1: Time complexity analysis on different attribution methods for the generative task.
Method Forward Backward Notes
Mask Perturbation $\mathcal{O}(N)$ 0 -
Gradient×$\times$×Input $\mathcal{O}(1)$ $\mathcal{O}(M)$
Integrated Gradients $\mathcal{O}(N_{step})$ $\mathcal{O}(N_{step}\cdot M)$ $N_{step}$ is the number of steps for integrating gradients.
LIME $\mathcal{O}(N_{aug})$ 0 $N_{aug}$ is the number of augmented samples.
SHAP $\mathcal{O}(2^{N})$ 0 -
To summarize, the traditional explanation methods are not always suitable for LLMs. In particular, the perturbation-based and gradient-based explanations are relatively easy to extend for attributing LLM responses to the input prompts, while the surrogate-based and decomposition-based methods become significantly challenging to do so. Specifically, surrogate-based methods suppose that an explainable small model could approximate the decision boundary of the target model around a local example, but there are limited explainable models for the text generation task. Meanwhile, decomposition-based methods require designing decomposition strategies for different layers, which is challenging for big LLM architectures. Another primary concern is their significant demand for computing resources. Given an $N$-words input prompt and an $M$-words output response, we present the time complexity of several representative explanation methods in Table 1. It demonstrates that existing methods either require a large number of forward operations or backward operations. Therefore, improving the efficiency of the attribution-based explanation is an important direction for future research and development.
2.1.2 Attributing Inputs for Text Generation
Figure 2: An example of attribution saliency map between input and output tokens.
The explanation of generative models can be defined as attributing the overall confidence $p(\hat{y}|x)$ to the input $x$, where $\hat{y}$ denotes the generated response $\hat{y}=[\hat{y}_{1},...,\hat{y}_{M}]$ with $M$ words. One method to achieve this is by treating the text generation process as a sequence of word-level classification tasks. This perspective allows for the application of existing classification-based explanation techniques to assess the influence of each input word $x_{n}$ in relation to each output word $\hat{y}_{m}$, resulting in a corresponding attribution score $a_{n,m}$. After gathering the attributions $a_{n,m}$ for $m=1,...,M$, we perform an aggregation to determine the overall contribution of each input word $x_{n}$. This is accomplished by aggregating the individual attributions for all output words corresponding to the input word, denoted as $a_{n}=\mathrm{Aggregate}([a_{n,1},...,a_{n,M}])$. The simplest approach for this aggregation is to average the attributions assigned to each input word across the different output words (Selvaraju et al., 2016). However, Wu et al. (2023) observe that attribution scores from different output words are not inherently comparable. For example, the attribution scores for function words (e.g., “the”, “is”, “have”) are often disproportionately larger than the scores for content words with clear semantic meaning (e.g., verbs and nouns). Therefore, it is necessary to normalize the scores prior to the aggregation, so that the scores $[a_{n,1},...,a_{n,M}]$ become comparable for $1\leq m\leq M$. Figure 2 plots the normalized scores of an example case, where each index in the Y-axis refers to an input prompt token, while that in the X-axis is an output response token. A greater normalized attribution score is brighter. In this example, the user attempts to direct the model to output information that does not exist, namely the French president in 1250. The model successfully realizes that this thing does not exist and refuses to answer. The model response can be realized as three parts, “There was no”, “president in France”, and “in 1250”. According to the figure, the first span is generated heavily because of the tokens “Who” and “president”, while the model uses both “France” and “1250” to respond to the second span “president in France”. Finally, the model emphasizes the date “1250” again by referencing the same information from the prompt. Overall, these explanations align with human understanding and highlight the usage of this method in the future. However, current research on attribution-based explaining for generative LLMs is still in its early stages, and only a limited number of methods have been proposed.
2.2 Case Studies: Usability of Attribution Methods for LLMs
Figure 3: A general pipeline of model diagnosis with attribution explanations.
The attribution map offers a partial insight into the operational mechanics of LLMs (Chandrasekaran et al., 2018; Hase & Bansal, 2020; Ye & Durrett, 2022a). Accordingly, we propose a general pipeline that leverages attribution scores to analyze LLM behaviors, as shown in Figure 3. First, given the target LLM and an input prompt, we compute attribution scores of input tokens relative to the output tokens. Second, we extract a feature vector from the attribution map, tailored to the requirements of the diagnostic task at hand. Third, we train a light-weight predictor (e.g., a classifier) to diagnose whether the model behaves appropriately based on the feature vector. In the following, we provide case studies to illustrate how attribution scores could be utilized to assess LLM response quality (Adlakha et al., 2023).
2.2.1 LLM Response Quality Evaluation with Explanations
This case study explores the use of attribution-based explanations as evidence for assessing the quality of LLM-generated responses. Here, “quality” is specifically measured by the responses’ accuracy. We hypothesize that responses generated from correct rationales are likely to be more accurate. Our method involves comparing the model’s explanations with underlying rationales as a means to evaluate response accuracy. This approach could enhance the reliability of conversational systems, which is crucial for applications such as medical question answering.
Dataset.
We employ the Multi-Sentence Reading Comprehension (MultiRC) dataset (Khashabi et al., 2018) in this case study. MultiRC presents a more significant challenge than other datasets: it requires the system to answer questions based on multiple sentences from a given paragraph. The answers may not directly correspond to specific sentence spans, reflecting a more realistic scenario of modern LLMs. Each MultiRC instance is associated with a human-annotated rationale indicating which sentences are relevant for answering the question. Our study specifically employs its test subset with 950 samples in total. We consider 80% in which are the training set and we report the results evaluated on the rest samples.
Implementation.
We apply the Importance Density Score method (Wu et al., 2023) to estimate the importance of each word in the provided paragraph. Specifically, given $N$-word paragraph $X$ and the $M$-word response $\hat{Y}$, the attribution from each input word $x_{n}$ to each output word $\hat{y}_{m}$ is defined as $a_{n,m}=\frac{\partial f(y_{m}|Z_{m})}{\partial\,\mathbf{E}_{i}[x_{n}]}\cdot\mathbf{E}_{i}[x_{n}]^{\top}$, where $Z_{m}$ indicates the formatted input prompt concatenated with the first $m-1$ response word, $f$ is the language model, and $\mathbf{E}_{i}[x_{n}]$ indicates the input static embedding of $x_{n}$. This pairwise attribution score is then normalized as $\widetilde{a}_{n,m}=\lceil L\times a_{n,m}/\text{max}_{n^{\prime}}(a_{n^{\prime},m})\rceil$. Any normalized attribution $\widetilde{a}_{n,m}$ is equal or less than $b$ will is forced to 0. The overall attribution score on word $x_{n}$ is defined as its attribution density, i.e., $a_{n}=||[a_{n,1},...,a_{n,M}||_{1}/||[a_{n,1},...,a_{n,M}]||_{p}$. In our experiments, we let $L=10$, $b=2$, $p=5$, and consider Vicuna-7B-v1.1 as our language model $f$. By averaging these word attribution scores, we reach the importance of each sentence. The top-K sentences with the highest importance scores are selected as the explanation for each instance. The explanation is then concatenated with the output response and fed to a classifier to predict whether the output is correct or not. We initialize the classifier with DistillBERT-base (Sanh et al., 2019), and train it over 3 epochs with learning rate $5e^{-5}$ and weight decay $1e^{-3}$. To compare with this, we also consider the Full Paragraph or the human-annotated rationales as the explanation part to train the classifier. Following previous studies (Khashabi et al., 2018; DeYoung et al., 2019), we evaluate the accuracy of the generated responses by exactly matching all correct answers with the generated responses. Table 2 reports the macro precision, recall, F1, and AUC scores on the test set over five random seeds.
Results.
Table 2: Utilizing attribution for response quality evaluation
Setting Precision Recall F1 AUC
Random 49.40 51.79 49.61 49.03
Human Rationale 68.73 66.88 67.57 73.11
Full Paragraph 58.02 58.47 56.89 63.44
Attribution (ours) 63.25 67.69 64.12 71.53
In Table 2, we observe that isolating the rationales from the full context could best help the classifier identify the response quality. It is evident that explanations align more closely with human-annotated rationales when associated with correct responses. In particular, the precision and recall metrics for explanations derived from correct answers surpass those associated with incorrect responses. This finding serves as empirical evidence underscoring the effectiveness of attribution-based explanations in estimating the accuracy of responses generated by LLMs.
2.2.2 Hallucination Detection with Attribution Explanations
This case study explores a different aspect of LLM generation quality, focusing on the presence of hallucinations in LLM-generated responses. We show that attribution-based explanations can serve as indicators to detect LLM hallucinations. Hallucinations are defined as responses that contain information conflicting with or unverifiable by factual knowledge (Li et al., 2023c; Ji et al., 2023). For instance, if a model is asked about a fictitious entity like “King Renoit” and responds with a narrative about “The Three Musketeers”, claiming it pertains to the nonexistent king, it illustrates a hallucination. This tendency, particularly pronounced in instruction-tuned models, stems from their earnest endeavor to fulfill user requests. The issue often arises when direct commands (“tell me a story”) significantly influence the generation process, while the instruction’s subject (“about King Renoit”) is neglected. Based on this insight, we develop a hallucination detector according to the distribution of attribution scores over different types of prompting words.
Dataset.
We use the Hallucination Evaluation Benchmark (Li et al., 2024a) in this case study. Each instance from this dataset includes an input prompt, a response generated by ChatGPT (Ouyang et al., 2022), and the knowledge related to the response. Each piece of knowledge has a human annotation about whether the knowledge is valid or not. Each instance having at least one invalid knowledge is considered a hallucination response. This study focuses on 632 less confused examples from the benchmark, each of which has almost all knowledge correct or almost all wrong. We randomly select 80% of the samples to form a training set and the rest form the testing set.
Implementation.
Given a query prompt and its ChatGPT response, we aim to build a classifier to detect if the response contains hallucination. Since the gradients of ChatGPT is inaccessible, we apply the Vicuna-7B model as the substitute to compute the attribution scores. Specifically, we adopt Importance Density Scores (Wu et al., 2023) to compute the attribution scores of Vicuna. We then use the NLTK package to identify the part-of-speech (POS) tag of each query word. Finally, each query-response pair is represented with an 82-dimensional vector, where each dimension indicates the average attribution scores over a certain type of POS tagging. We develop a Support Vector Machine classifier based on the POS tagging attribution scores on the training set and report the precision, recall, F1 scores on the hallucination samples, as well as accuracy over all samples in Table  3. To compare with this, we also leverage a fine-tuned model and a prompting-based method to serve as the hallucination detector baselines.
Results.
Table 3: Utilizing attribution for hallucination detection on ChatGPT responses.
Method Language Model Precision Recall F1 Accuracy
Random - 88.41 50.34 64.11 50.59
FacTool GPT-4 95.30 72.93 82.62 73.04
Vectara DeBERTa-base 90.29 60.54 72.40 59.45
AttrScore (ours) Vicuna-7B 90.15 74.21 81.36 70.20
Mistral-7B 88.74 75.04 81.26 69.57
In the table, we first observe that all methods have demonstrated a greater performance than the Random strategy, indicating that the proposed methods and the baselines are valid hallucination detectors. In addition, we observe that the proposed methods achieve a competitive performance with FacTool using GPT-4, establishing a new benchmark in this domain. It is worth noting that, instead of using GPT-4, the attribution score for our method is calculated using a smaller 7-billion-parameter language model. This demonstrates the practicality and efficiency of our approach in weak-to-strong generalization as we could diagnose large language models with smaller models. Future work could consider extracting more effective features and using more powerful classifiers.
2.3 Challenges
2.3.1 How to Identify and Explain the Semantics of Output?
The attribution function $a_{n}=p(\hat{y}|x)-p(\hat{y}|\widetilde{x}_{n})$ is no longer faithfully attributing the model predictions in the human-interested semantic level since the model could express the same semantic meaning with various responses. Specifically, the model could assign a lower confidence to its original response than the new one, while both responses share the same semantic meaning. This is a significant difference compared with the traditional classification problem, where target label sets are manually designed so that a lower $p(\hat{y}|\cdot)$ indicates the model is less confident in predicting a specific semantic concept. Taking the sentiment analysis task as an example, an LLM may generate two different responses sharing the same predicted concepts, such as “it is a positive review” and “the audience thinks the movie is great”. Current attribution-based explanations concentrate on the literal changes in generated responses, but they do not study how the semantic meanings of these responses change. Therefore, they do not provide sufficient explanations of model-generated responses at a semantic level. In this case, the semantic level is which words of the input review lead the model to believe it is positive. Future work may tackle this challenge by proposing metrics to evaluate the semantic differences in responses.
Figure 4: Review of interpretation methods for LLM components and their applications. We categorize methods according to the target LLM modules: self-attention layers and feed-forward layers.
2.3.2 Explaining LLM Predictions Beyond Attribution
The versatility of LLM generation will inspire a variety of explanation paradigms beyond traditional attribution methods. Attribution methods aim to explain model output with the contribution of input features. This explanation task is meaningful for conventional machine learning (ML) models whose outputs are usually individual decisions with clear formats (e.g., classification, regression, object detection). The decisions are highly dependent on the input features. However, LLMs differ from traditional ML models in two aspects. First, the generative process of LLMs is stochastic, e.g., we may obtain different outputs by running an LLM twice even with the same input. Second, LLMs encode rich knowledge within their parameters, which are independent of inputs. These unique properties give rise to novel explanation paradigms. For the first aspect, an explanation task of interest would be to understand the uncertainty of LLM generation. For example, researchers (Ahdritz et al., 2024; Varshney et al., 2023; Su et al., 2024) leverage the prediction perplexity to check whether the LLM is confident during generation, identifying potential errors in less confident predictions. Second, attributing LLM predictions to their encoded knowledge instead of input patterns could provide a new perspective. Some researchers (Yin et al., 2024a) propose the knowledge-boundary detection task to detect whether a model has specific knowledge for a given question. Nevertheless, it did not attribute the prediction to specific knowledge, so humans cannot verify the prediction process with their results yet.
3 LLM Diagnosis and Enhancement via Interpreting Model Components
This section discusses the XAI methods that interpret the internal components of large language models. Additionally, it delves into the insights these methodologies offer, which can be instrumental in refining and enhancing the design of language models. LLMs adopt transformers as the basic architecture, which typically comprises two types of major components: self-attention layers and feed-forward layers. In the following, we review the research that focuses on interpreting each of these components respectively.
3.1 Understanding the Self-Attention Module
A multi-head self-attention module contains multiple self-attention heads capturing different kinds of word-word relations, which are modeled with weights $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D_{1}\times D_{2}}$. Specifically, the relation of words $w_{i}$ and $w_{j}$ is computed as $\mathbf{A}_{i,j}\propto(\mathbf{x}_{i}\mathbf{W}_{q})\cdot(\mathbf{x}_{j}\mathbf{W}_{k})^{\top}$, where $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{1\times D_{1}}$ are contextual embeddings of the words. The most straightforward interpretation is analyzing the attention score matrix $\mathbf{A}$ given an input sequence to study the relations between words (Vig, 2019; Hoover et al., 2020). In practice, these intuitive explanations would be majorly used to present case studies via visualization. With this strategy, Wang et al. (2023b) conduct case studies on in-context sentiment analysis, where they find that the label words from the in-context examples serve as anchors of the final prediction. In specific, these anchors aggregate information from the examples to generate informative representations from the lower layers, while the deeper layers utilize these representations to make final predictions. This insight motivates them to reweight the attention scores of these anchors to achieve better inference accuracy. Some researchers (Dar et al., 2023; Wu et al., 2023) extend this framework to globally analyze the attention weights $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ by feeding the static word embeddings of words from an interested vocabulary, instead of their contextual embeddings. For example, with this approach, Wu et al. (2023) find that instruction tuning empowers LLMs to follow human intentions by encouraging them to encode more word-word relations related to instruction words. On the other hand, some mathematical models are proposed to theoretically explain the self-attention mechanism, such as Sparse Distributed Memory (Bricken & Pehlevan, 2021) and Transformer Circuits (Elhage et al., 2021). In particular, the Transformer Circuits provide mechanistic interpretability to transformer-based models, which breaks models down into human-understandable pieces. Although these theoretical analyses on self-attention solids a foundation for future research, their direct application is largely underexplored.
3.2 Understanding the Feed-Forward Module
A feed-forward network is formalized as $\mathbf{x}^{\prime}=\sigma(\mathbf{x}\mathbf{W}_{u})\mathbf{W}_{v}^{\top}$, where $\mathbf{x}\in\mathbb{R}^{1\times D_{1}}$ is the intermediate contextual representation of an input word, $\sigma$ is a non-linear operation, and $\mathbf{W}_{u},\mathbf{W}_{v}\in\mathbb{R}^{D_{1}\times D_{3}}$ are model parameters. Feed-forward networks can be understood as key-value memories (Sukhbaatar et al., 2015; Geva et al., 2021), where each key or value is defined as $\mathbf{W}_{u}[d]\in\mathbb{R}^{D_{1}}$ and $\mathbf{W}_{v}[d]\in\mathbb{R}^{D_{1}}$, respectively. That is, each feed-forward network obtains $D_{3}$ key-value pairs, called memories. One simple way to interpret the semantic meaning of memory is collecting the words that could maximally activate the key or value vector of that piece of the memory (Geva et al., 2021; Dar et al., 2023), which has demonstrated strong interpretability of the extracted word lists. However, it is critical to be aware that the key or value vectors are polysemantic (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023), indicating that this simple approach might not provide concise explanations for each key-value pair. It has been shown that the word list of each key-value pair has 3.6 human interpretable patterns on average (Geva et al., 2021). To alleviate the limited interpretability caused by the nature of polysemantic, Wu et al. (2023) propose to interpret the principal components of these key or value vectors, leading to a more concise explanation for each word list, such as “medical abbreviations” and “programming tasks and actions”. Other work examines individual memories by measuring the changes of predictions after perturbing their corresponding activations, where it reveals that some memories encode specific knowledge (Dai et al., 2022) and some others capture general concepts (Wang et al., 2022b). By leveraging the explanations of key-value memories, we could locate and update the memories associated with a specific piece of knowledge to perform model editing (Dai et al., 2022; Meng et al., 2022a; b; Hase et al., 2024), i.e., modifying outdated or incorrect knowledge. Another usage of these weight explanations is model pruning, where LLMs can be condensed up to 66.6% of their initial parameters by exclusively maintaining redundant neurons, resulting in an approximate 1.4x increase in inference speed (Dalvi et al., 2020).
Besides interpreting and analyzing model weights, some work studies model activations (e.g., $\mathbf{x}\mathbf{W}_{u}$ or $\mathbf{x}^{\prime}$) to interpret their functionality. The probing technique is the most popular way for this purpose (Belinkov et al., 2018; Tenney et al., 2018; Jawahar et al., 2019; Rogers et al., 2021), identifying whether a specific concept is encoded within the representations. The basic idea is developing an auxiliary classifier $g$ to map from the representations $\mathbf{x}^{\prime}$ to the interested concept space $\mathcal{C}$, such as syntax and part-of-speech knowledge, and the performance of $g$ interprets how much information encoded in $\mathbf{x}^{\prime}$ is related to the concepts in $\mathcal{C}$. This technique motivates developing better parameter-efficient (Chen et al., 2022), domain-specific (Das et al., 2023), and robust (Bai et al., 2021; Wang et al., 2023a) LLMs. Recent studies (Chen et al., 2023a; Ahdritz et al., 2024) also apply the probing method to detect the knowledge boundary of a LLM so that the hallucinated responses could be reduced. Some researchers (Bricken et al., 2023; Cunningham et al., 2023) point out another direction to interpret the model hidden activations, called dictionary learning, which is motivated by the assumption of superposition (Elhage et al., 2022; Sharkey et al., 2022). The superposition assumes that LLMs would learn an over-complete set of non-orthogonal features, thereby transcending the limitations imposed by the dimensionality of the representation space. Therefore, the researchers aim to reconstruct and interpret these features to understand the internals of the model. Practically, they develop a sparse auto-encoder $g$ to reconstruct the representations $\{\mathbf{x}_{n}\}$, which shows that humans could well interpret the learned sparse features of $g$ according to their most activation words. Their research shows that this method could be used for more controllable generation. Specifically, if forcing a sparse feature to be activated, then the language model $f$ would change its response to perform the particular behavior of that sparse feature. For example, given “1,2,3,4,5,6,7,8,9,10” as input, the model originally generates numbers as output. However, when they are forced to magnify the activations of a sparse feature called “DNA”, the model changes its output to “AGACCAGAGAGAAC”. In general, while the explanation techniques for feed-forward networks primarily offer insights for model development, they have also demonstrated promising applications in areas such as model editing and controllable generation.
3.3 Challenges
Interpreting the functionality of internal modules is still in its infancy, and we identify two challenges to be tackled in this direction.
3.3.1 Complexity of Individual Models and Their Interactions
The transformer-based language model contains two types of modules that collaborate based on the residual mechanism (He et al., 2016), which enables later modules to utilize, enhance, and/or discard outputs from preceding modules. Formally, the output of the $l$-th module is denoted as $x^{l}=f^{l}(x^{l-1})+x^{l-1}$, where $f^{l}$ could be a self-attention module or a feed-forward network. Research in this area aims to interpret how different modules $f^{i}$ and $f^{j}$ work together for $i\neq j$. Pilot studies (Elhage et al., 2021; Olsson et al., 2022) find that stacked self-attention modules could form Induction Heads, which demonstrate a strong correlation with the in-context learning capability. Specifically, the induction head encourages the model to predict the word “B” followed by a sequence “AB…A”. Their study finds a specific phase during pre-training LLMs where both induction heads and the in-context learning capability emerge from the model. Following this track, researchers observe diverse functional heads within LLMs for different tasks, such as “Name Mover Head” and “Duplicate Token Head” for the object identification task (Wang et al., 2022a), “Single Letter Head” and “Correct Letter Head” for the multiple-choice question answering task (Lieberum et al., 2023), and “Capitalize Head” as well as “Antonym Head” for the general purpose tasks (Todd et al., 2023). Although these studies have indeed deepened our understanding of cross-module effects, their analyses are grounded on specific tasks or scenarios, leaving it uncertain whether these findings possess broad generalizability.
3.3.2 Nature of Polysemantic and Superposition Assumption
Interpreting the functionality of a single neuron (one row/column vector of a weight matrix) fails in analyzing large language models since a single neuron could be activated by multiple and diverse meanings, called polysemantic (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023). This nature leads to poor interpretability: explaining a single neuron usually does not reflect a concise human concept. Some researchers (Elhage et al., 2022; Sharkey et al., 2022) assume that this phenomenon is caused by the superposition of an over-complete set of features learned by the models. Based on this assumption, we may reach another level of explanation by decomposing the model weights to reconstruct a large number of features. However, the two critical problems of this approach are still unclear: (1) How do we ensure our reconstructed features faithfully represent the latent features learned by the models? (2) How do we interpret our reconstructed features with human language?
4 LLM Debugging with Sample-based Explanation
In this section, we discuss sample-based explanation strategies for LLMs, which aim to trace back the answers generated by LLMs to specific training samples (i.e., documents) or document fragments in the corpora. The utility of sample-based explanations for LLMs is multifaceted. First, tracing back the predictions of LLM to the training samples can provide evidence for the generation results, which facilitates model debugging in cases of errors and increases the trustworthiness of the model from users when the outcomes are accurate. In addition, it can also help researchers understand how LLMs generalize from training samples. If the outputs of LLMs can be traced back to exact subsequences directly spliced from the training data, it might suggest that the LLM is simply memorizing the data. In contrast, if the generation results and the influencing training samples are abstractly related, it could indicate that LLMs can understand the concepts and generate responses by reasoning from input prompts.
In this section, we start by systematically reviewing traditional sample-based explanation strategies, including gradient-based methods and embedding-based methods, as well as some preliminary explorations to generalize them to LLMs. We then analyze the challenges associated with generalizing the above strategies to LLMs with unique transformer structures and unprecedented numbers of parameters. Finally, we discuss the insights to address the challenges, as well as open challenges worthy of further investigation.
4.1 Literature Review of Sample-based Explanation
In this section, we denote the input space and output space as $\mathcal{X}$ and $\mathcal{Y}$, respectively. In the context of large language models (LLMs), $\mathcal{X}$ is the space of token sequences known as the prompts, and $\mathcal{Y}$ could be the space of discrete labels in classification tasks or the space of token sequences as output in generation tasks***Here, we should note that, in the pretraining phase of LLM with language modeling, the model either predicts masked tokens in $x_{i}$ (i.e., masked language modeling), or autoregressively predicts the next token (i.e., causal language modeling) in $x_{i}$. Therefore, in some works, $y_{i}$ is omitted, and only $x_{i}$ is included for discussions.. Accordingly, we have a training dataset $\mathcal{D}_{train}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{N}$ with $N$ samples drawn from the joint space $\mathcal{X}\times\mathcal{Y}$, on which an LLM model $f_{\theta}$ is trained with pretrained parameters $\hat{\theta}\in\mathbb{R}^{P}$. We also have a test sample $z=(x,y)$ of interest, where we want to explain the generation of $y$ from $x$ based on training samples in $\mathcal{D}_{train}$ (which can be viewed as the information source). The goal of sample-based explanation is to measure the influence of a training sample $z_{i}\in\mathcal{D}_{train}$ or a certain segment within $z_{i}$, such that the generation of LLMs can be well-explained and backed up by the selected training samples.
4.1.1 Influence Function-based Methods
One strategy to quantify the influence of a training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ to a test sample $z$ is through the influence function (Koh & Liang, 2017; Han et al., 2020). It measures the change of the prediction loss $\mathcal{L}\left(z,\theta\right)$ for the test sample $z$, when the training sample $z_{i}$ undergoes a hypothetical modification in the dataset $\mathcal{D}_{train}$ during model training. This modification results in an altered set of optimal model parameters, denoted as $\hat{\theta}_{-z_{i}}$. The most common modification of a training sample is to remove it from the dataset, where the influence of the removal of a training sample $z_{i}$ on the loss at test sample $z$ can be computed as follows:
$\mathcal{I}(z_{i},z)=-\nabla_{\theta}\mathcal{L}(z,\hat{\theta})^{\top}\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}(z_{i},\hat{\theta}),$ (1)
where $\nabla_{\theta}\mathcal{L}(z,\hat{\theta})$ is the gradient of the loss function $\mathcal{L}$ on the test sample $z$ evaluated at the optimal parameters $\hat{\theta}$, and $\mathbf{H}_{\hat{\theta}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}^{2}\mathcal{L}(z_{i},\hat{\theta})$ denotes the Hessian matrix of the LLM model at parameter $\hat{\theta}$. If we denote the number of parameters in $\hat{\theta}$ as $P$, the naïve inversion of the Hessian matrix $\mathbf{H}_{\hat{\theta}}$ leads to $\mathcal{O}(NP^{2}+P^{3})$ time complexity and $\mathcal{O}(P^{2})$ space complexity (Schioppa et al., 2022), which is clearly infeasible for large models. To improve efficiency, Koh & Liang (2017) adopt an iterative approximation process, i.e., LiSSA (Linear time Stochastic Second-Order Algorithm), to calculate the Hessian-Vector Product (HVP) in Eq. (1), where the memory complexity can be reduced to $\mathcal{O}(P)$ and time complexity to $\mathcal{O}(NPr)$ ($r$ is the number of iterations). To further reduce the complexity, Pruthi et al. (2020) propose an alternative to Eq. (1), i.e., TracIn, which measures the influence of $z_{i}$ on $z$ by calculating the total reduction of the loss on $z$ whenever $z_{i}$ is included in the minibatch during model training. The TracIn measurement can be formulated as follows:
$\mathcal{I}_{\operatorname{TracIn}}\left(z_{i},z\right)=\sum_{t:z_{i}\in\mathcal{B}_{t}}\mathcal{L}\left(z,\theta_{t}\right)-\mathcal{L}\left(z,\theta_{t+1}\right)\approx\frac{1}{b}\sum_{t:z_{i}\in B_{t}}\eta_{t}\nabla_{\theta}\mathcal{L}\left(z_{i},\theta_{t}\right)\cdot\nabla_{\theta}\mathcal{L}\left(z,\theta_{t}\right),$ (2)
where $\mathcal{B}_{t}$ is the $t$-th mini-batch fed into the model during training, $\theta_{t}$ is the parameter checkpoint at the $t$-th step, $\eta_{t}$ is the step size, and $b$ is the size of the mini-batch. According to the above equation, TracIn only leverages gradient terms, where Hessian $\mathbf{H}_{\hat{\theta}}$ is removed from the influence measurement. This substantially improves the efficiency. However, such complexity is still prohibitive for large models from both the computational and memory perspectives. In addition, TracIn can only estimate the influence of adding/removing the sample to the loss, where variants of the vanilla influence function defined in Eq. (1) can measure the influence of other modifications of the training sample $z_{i}$, such as perturbation (e.g., masking out a segment of a document $x_{i}$). To adapt the vanilla influence function of Eq. (1) to explain transformers, Schioppa et al. (2022) propose to use Alnordi iteration (Arnoldi, 1951) to find the dominant eigenvalues and eigenvectors of the Hessian matrix on randomly sampled subsets $\mathcal{D}_{sub}$, with $|\mathcal{D}_{sub}|\ll|\mathcal{D}_{train}|$. In such a case, the diagonalized Hessian can be cheaply cached and inverted, where the computational and memory complexity can be substantially reduced. Previous work mainly focuses on reducing the complexity of calculating the influence of a single training sample. Observing that finding the most influential training sample on $z$ needs to iterate Eq. (1) overall $N$ training samples, Guo et al. (2021) propose to use fast KNN to pre-filter a small subset of influence-worthy data points from $\mathcal{D}_{train}$ as candidates to explain small pretrained language models, whereas Han & Tsvetkov (2022) propose to iteratively find a small subset $\mathcal{D}_{sub}\subset\mathcal{D}_{train}$ whose gradient is the most similar to that of the downstream task examples. Recently, Grosse et al. (2023) propose to use the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation to scale influence functions to LLMs up to 52B parameters. For adaptation, only influences mediated by the multi-layer perceptron (MLP) layers are considered, whereas weights from the self-attention layers are fixed, as previous work has demonstrated that knowledge is mainly encoded in the MLP layers (Wang et al., 2023d). In addition, based on the assumption that weights from different MLP layers are independent, the EK-FAC approximated influence can be formulated as the sum of influences mediated by each layer:
$\mathcal{I}_{\text{EKFAC}}(z_{i},z)=\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(z,\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta}),$ (3)
where $\theta^{(l)}$ denotes the weights of the $l$-th MLP layer, and $\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}$ is the EK-FAC approximated Gauss-Newton Hessian for $\theta^{(l)}$. Since the inversion of $L$ small ${K_{l}\times K_{l}}$ matrices (i.e., $\mathcal{O}(L\times K_{l}^{3})$) is substantially more efficient than the inversion of a large $L{K_{l}\times LK_{l}}$ matrix (i.e., $\mathcal{O}{((LK_{l}})^{3})$), $\mathcal{I}_{\text{EKFAC}}$ can be adaptable to very large models, let alone the HVP can be further simplified by utilizing the EK-FAC properties.
Recently, influence function-based explanation has been used in efficient finetuning of LLMs, where influence function has been used to select a small subset of training samples given few-shot validation samples for a specific downstream task, where the training overhead can be substantially improved (Xia et al., 2024).
4.1.2 Embedding-based Methods
Another strategy for sample-based explanation involves leveraging the hidden representations within the transformer architecture, which is recognized for encoding high-level semantics from textual data, to calculate the semantic similarity between $z$ and $z_{i}$. The similarity can also be used to measure the influence of $z_{i}$ on $z$ as explanations (Rajani et al., 2019). Specifically, Akyurek et al. (2022) propose to represent the training sample $z_{i}$ and test sample $z$ by concatenating the input and output as $z^{cat}_{i}=[x_{i}||y_{i}]$, $z^{cat}=[x||y]$. The concatenation is feasible for generation tasks where the output $y$ lies in the same token sequence space as the input prompt $x$. The similarity between $z_{i}$ and $z$ can then be calculated as follows:
$\mathcal{I}_{\text{emb}}\left(z_{i},z\right)=\frac{f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\cdot f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)}{\left\|f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\right\|\left\|f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)\right\|},$ (4)
where $f^{(l)}_{\hat{\theta}}$ is the sub-network that outputs the $l$-th layer intermediate activation of the pretrained LLM $f_{\hat{\theta}}$. The Eq. (4) has a similar form as the vanilla influence function defined in Eq. (1) as well as its TracIn alternative defined in Eq. (2), which assigns a score $\mathcal{I}$ for the explainee $z$ for each training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ as the explanation confidence of the sample $z_{i}$.
Compared with the influence function methods introduced in the previous part, embedding-based methods are computationally efficient, as for each explainee $z$, the explanation score from a training sample $z_{i}$ requires only one forward pass of the transformer network. In addition, the calculation can be easily paralleled for different training samples. However, the disadvantage is also evident: These methods lack a theoretical foundation and may fail to identify important training samples that may not be semantically similar to the test sample. Consider the following toy example: Training samples $z_{i}$ = (“1+1=”, “2”) and $z_{j}$ = (“2+2=”, “4”) make the LLM gain the ability to conduct arithmetic calculation, which explains why prompting the model with $x$ = “100+100” gives the results $y$=“200”. However, the embeddings between the test sample $z$ and the two training samples $z_{i}$ and $z_{j}$ can be very different when calculated via Eq. (4(Akyurek et al., 2022). Therefore, embedding-based methods may not be able to faithfully find the training samples where the explanations require generalization ability beyond semantic similarity.
4.2 Case Study: EK-FAC-based Influence Estimation
In this part, we implement the EK-FAC-approximated influence function proposed in Grosse et al. (2023), and verify its scalability and effectiveness on LLMs with billions of parameters, including GPT2-1.5B (Radford et al., 2019), LLaMA2-7B (Touvron et al., 2023b), Mistral-7B (Jiang et al., 2023), and LLaMA2-13B.
4.2.1 Experimental Design
We use the SciFact dataset (Wadden et al., 2020) as the corpora, which contains the abstract of 5,183 papers from the domain of basic science and medicine. The explainee LLMs are obtained by finetuning the pretrained LLMs for 20,000 iterations, where AdamW (Loshchilov & Hutter, 2018) is used as the optimizer, and the learning rate and weight decay are set to 1e-5, 1e-3, respectively. Then, we use 500 samples from the corpora to estimate the (i) uncentered covariance matrices of the activations and pre-activation pseudo-gradients $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$, and (ii) the variances of the projected pseudo-gradient $\bm{\Lambda}^{(l)}$ for each selected dense layer $l$, and cache them on the hard disk (details see Eqs. (16) and (20) in Grosse et al. (2023)). We select the c_fc layer for GPT2-1.5B, and gate_proj layer for LLaMA2-7B, Mistral-7B, and LLaMA2-13B***All the implementation and layer names are based on the huggingface transformers, where the details can be found in https://huggingface.co/docs/transformers/en/index..
For evaluation, we randomly select 200 samples from the corpora to construct the test set, which we name SciFact-Inf. Specifically, for the $j$-th selected sample $z_{j}=(x_{j},x_{j})$ (here $y_{j}=x_{j}$ as label equals the input in language modeling), we use the first three sentences in $x_{j}$, i.e., $\hat{x}_{j}$, to generate a completion $\hat{y}_{j}$ with the finetuned LLM (here, $\hat{y}_{j}$ does not equal the remaining sentences in $x_{j}$), and we aim to explain the generation of $\hat{y}_{i}$ from $\hat{x}_{j}$ with the finetuned LLM with the training samples via EK-FAC approximated influence scores defined in Eq. (3). Ideally, the $j$-th training sample $z_{j}$ itself should be the most influential sample w.r.t. the generation of $\hat{y}_{j}$ for test sample $\hat{z}_{j}$, which facilitates quantitative analysis of the effectiveness of Eq. (3).
In our implementation, for each test sample $\hat{z}_{j}$, we first calculate the EK-FAC approximated HVP part of the influence $\mathcal{I}_{\text{EKFAC}}(z_{i},\hat{z}_{j})$, i.e., $\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(\hat{z}_{j},\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}$, which is shared for all training samples $z_{i}$. Specifically, we record the layer-wise gradient $\nabla_{\theta^{(l)}}\mathcal{L}(\hat{z}_{j},\hat{\theta})$ and calculate the HVP with the cached $\mathbf{Q}_{A}^{(l)}$, $\mathbf{Q}_{S}^{(l)}$ as Eq. (21) in Grosse et al. (2023). We then go through candidate training samples (1 positive and 99 negative), calculate the gradient $\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta})$, and take inner-product with the approximate HVP as the layer-wise influence. Finally, the layer-wise influences are summed up as Eq. (3) as the total influence $\mathcal{I}_{\text{EKFAC}}(z_{i},\hat{z}_{j})$. We rank the influence and calculate the top-$K$ hit rate of the positive training sample.
4.2.2 Results and Analysis
Table 4: Effectiveness of EK-FAC approximated influence function on the established SciFact-Inf dataset. Time (Pre.) stands for the time for precomputing the $\mathbf{Q}_{A}$, $\mathbf{Q}_{S}$, and $\bm{\Lambda}$. Time (Inf.) stands for the time for calculating the influence of 100 training samples per test sample. Experiments for GPT2-1.5B, LLaMA2-7B, Mistral-7B are done on 4 A100 GPUs, whereas experiments for LLaMA2-13B are done with 8 A100 GPUs.
Strategy LLM Recall@5 Recall@10 Time (Pre.) Time (Inf.)
Random - 0.0100 $0$
Inf. Func. GPT2-1.5B 0.6368 0.7363 0h 27min 0min 28sec
Mistral-7B 0.6418 0.6866 2h 05min 1min 47sec
LLaMA2-7B 0.8063 0.8308 1h 37min 1min 34sec
LLaMA2-13B 0.7811 0.8940 3h 11min 3min 08sec
The experimental results are summarized in Table 4. From Table 4 we can find that, the EK-FAC approximated influence function achieves a good accuracy in finding the training sample that has the greatest influence on the generation of a test sample, even if only the influences mediated by a small part of dense layers are considered. In addition, we find that the main computational bottleneck in calculating the EK-FAC-based influence is to estimate the covariances $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$ and variance $\bm{\Lambda}^{(l)}$, which can take hours when 500 training samples are used for the estimation. However, after the estimation, it is relatively cheap to calculate the influence of 500 training samples for each test sample, which takes around 3 minutes for a 13B LLaMA2 model on 8 A100 GPUs. This demonstrates the scalability of EK-FAC-based influence by assuming the independence of different dense layers and using EK-FAC to simplify the computation.
4.3 Challenges
Overall, explaining the generation of LLMs by tracing back to the training samples is still an emerging area. Open questions need to be addressed to further advance the field. In this section, we identify three main challenges as follows, which can serve as directions for future explorations.
4.3.1 Strong Assumptions for Scalability
The unprecedented number of parameters in modern LLMs causes severe scalability issues for sample-based explanation strategies. This is especially evident for the gradient-based methods, as the HVP in Eq. (1) induces both high computational and space complexity. To address the bottleneck, strong assumptions are usually required to make it feasible for large models. For example, TracIn (Pruthi et al., 2020) simplifies the second-order term in Eq. (1) via first-order approximation. Schioppa et al. (2022) assume the Hessian to be low rank. Grosse et al. (2023) that assume that the weights from different layers of the LLMs are independent, as well as the tokens in different steps, such that EK-FAC can be appropriately applied to approximate the influence function. From the above analysis, we can find that while the method from Grosse et al. (2023) has the best scalability, it also has the strongest assumption, which may fail to hold in practice. While highly efficient to compute, embedding-based methods make the implicit assumption that semantics similarly implies explainability, which we have demonstrated may not always be the case. Therefore, how to improve the scalability with weak assumptions needs to be investigated in the future.
4.3.2 Explainability v.s. Understandability
Despite the advantage of influence/embedding similarity to provide a qualitative measurement of a specific training sample as the explanation for LLM generation, the understandability of the identified sample can still be weak, where the connection between the selected training samples and the generation may not be understandable to human beings. Specifically, Grosse et al. (2023) cautions that the sign of influence score of the training tokens may be difficult for humans to connect to the positive or negative influence on the generation results. This severely jeopardizes the usability of the identified training samples. In addition, Grosse et al. (2023) also found that, since LLMs are usually not trained to the minimum to avoid overfitting (and due to overparameterization, the number of local minimums may be large), the connection between influence defined in Eq. (1) with the counterfactual loss of removing the sample $z_{i}$ at $z$ is also weak. For the embedding-based methods, since most LLM models are black box transformer models, the similarity of embeddings can also be hard to interpret by human beings; therefore, it is imperative to improve the interpretability of the identified training samples, such that tracing back becomes more meaningful.
4.3.3 LLM-Oriented Sample-based Explanations
Finally, we observed that both gradient-based and embedding-based methods are loosely connected to the LLM, as well as the backbone transformer networks. For example, algorithms like TracIn (Pruthi et al., 2020) are designed to scale up influence functions to large models, which are not specific for LLMs. Similarly, the embedding-based method proposed in Akyurek et al. (2022) is applicable to most machine learning models with latent representations. Grosse et al. (2023) considers the specialty of LLMs by utilizing the knowledge neuron assumption of the backbone transformers (Wang et al., 2023d) to simplify the influence function, where the weights considered are constrained to the MLP layers, which may not fully utilize the property of transformers. Therefore, how to further utilize the property of the LLM and the backbone transformer to design LLM-tailored sample-based influence/similarity (either to reduce the computational/space overhead or to improve the explanation quality) is highly promising for future work.
5 Explainability for Trustworthy LLMs and Human Alignment
In previous sections, we explore the use of explanation techniques for assessing and improving the performance of LLMs. In this section, we shift the focus towards examining LLM trustworthiness. As LLMs are increasingly integrated into various applications of daily life, including high-stakes areas like healthcare, finance, and legal advice, it is crucial that their responses not only are accurate but also align with human ethical standards and safety protocols (Liu et al., 2023b; Li et al., 2023f). Thus, the need arises to extend the scope of explanations from merely evaluating the accuracy of LLMs to scrutinizing their trustworthiness. Herein, we delve into how explanation techniques, discussed in the previous sections, can be instrumental in assessing LLMs across key aspects of trustworthiness like security, privacy, fairness, toxicity, and honesty. It is worth noting that while explainability itself is an aspect of trustworthiness, it holds the promise of serving as a foundational tool for addressing other trustworthiness concerns.
5.1 Security
LLMs are known to be vulnerable to attacks and exploitation, such as spreading misinformation, launching phishing attacks, and poisoning training data (Derner et al., 2023). For enhanced safety, LLMs are designed to reject certain types of prompts that may result in the generation of harmful content, e.g., by precluding prompts likely to elicit unsafe outputs from the training phase. However, jailbreak techniques can circumvent these restriction measures and manipulate LLMs into producing malicious contents. Malevolent users (i.e., attackers) can craft special prompts that compel or induce LLMs to prioritize instruction following over rejections (Liu et al., 2023c; Li et al., 2023a). For example, through Prefix Injection, attackers can use out-of-distribution prompt prefixes that are less likely to be rejected (walkerspider, 2022; Wei et al., 2023). Another approach, called Refuse Suppression, involves directing or persuading models to ignore established safety protocols (Wei et al., 2023; Zeng et al., 2024), where the instruction following ability is then employed to perform the attack.
Existing methods mainly rely on prompt engineering for attacking LLMs, but they usually have low attack success rates and significant time costs (Li et al., 2024c). Thus, by understanding and engineering latent representations of LLMs, explanation methods provide a viable way to design advanced attacks and discover the potential vulnerabilities of LLMs (Liu et al., 2021). For example, a recent work extracts “safety patterns” via explaining the latent space of LLMs using representation engineering. Specifically, these patterns can be captured from the activation differences between malicious queries and benign queries. The salient portion of difference vectors’ dimensions is localized and utilized to generate features of safety patterns. The safety patterns reflect the internal protection mechanisms within LLMs. Circumventing these patterns leads to novel attacks, which helps exploring potential vulnerabilities of LLMs (Li et al., 2024c). Besides, a deeper understanding of fine-tuning can shed light on the reliability of existing safety measures. In particular, Jain et al. (2023) use networking pruning, attention map activation, and probing classifiers to track the changes of model capabilities from pre-training to fine-tuning. These tools are helpful in finding salient weights that identify crucial neurons to restore pre-training capabilities. These neurons have demonstrated that the capabilities gained during fine-tuning can be removed easily through fine-tuning on other unrelated tasks. This finding casts doubt on the robustness of current safety alignments in LLMs.
5.2 Privacy
Recent studies have revealed that LLMs such as ChatGPT can leak extensive amounts of training data through a method known as divergence attacks. These attacks utilize specially crafted prompts to lead the model away from its standard chatbot-style generation (Nasr et al., 2023). The risk of private data exposure through such means poses a serious challenge to the development of ethically responsible models. This issue is compounded by strategies similar to jailbreak attacks, where misalignment is exploited to induce LLMs into operating in an unconventional “developer mode” via out-of-distribution prompts (Li et al., 2023a). Traditional data pre-processing techniques, such as data sanitization (Lison et al., 2021), are impractical as defenses for LLMs due to the enormous scale of training data.
Enhancing LLM privacy involves two strategic approaches: (1) preventing the models from memorizing sensitive data, and (2) establishing safeguards against the release of sensitive information during content generation. The latter can employ techniques used in jailbreak defenses, treating prompts that solicit private information as potentially malicious. The former approach requires identifying whether LLMs possess specific knowledge, which is traditionally tested by crafting question-answering (QA) tasks to determine if LLMs can provide answers. However, this approach faces limitations due to LLMs’ sensitivity to the phrasing of QA prompts, while the optimal prompt is usually unknown. To tackle the challenge, explanatory techniques can serve as a tool to confirm whether LLMs have internalized certain knowledge. For instance, via explaining the relation between factual knowledge and neuron activations (Meng et al., 2022a; Dai et al., 2022; Hase et al., 2024), we may investigate whether and where a piece of factual knowledge is stored within transformers. In addition, Yin et al. (2024a) recently proposes the concept of “knowledge boundary” and develops a gradient-based method to explore whether LLMs master certain knowledge independent of the input prompt.
5.3 Fairness
Despite LLMs’ powerful generation capabilities, their widespread applications also bring concerns about exacerbating bias issues in society, as LLMs are able to learn social biases within human-generated corpus (Gallegos et al., 2023). For example, in a gender bias case, “[He] is a doctor” is much more likely than “[She] is a doctor” because of the gender stereotype. In this subsection, we focus on fairness issues that refer to biases related to race, gender, and age within human communities (Li et al., 2023g). There is a rich body of literature quantifying fairness issues within LLMs through various tests (Adebayo et al., 2023). Interpretation complements these methods by providing a unique perspective for bias mitigation, focusing on unraveling the mechanisms through which biases are embedded into LLMs. A research direction within this domain is the examination of biased attention heads. For instance, Ma et al. (2023) detect stereotype encodings through probing attention heads and evaluating their attributions via Shapley values (Lundberg & Lee, 2017). The results unveil that approximately 15% to 30% of attention heads across six transformer-based models are linked to stereotypes. These attention heads tend to specialize in maintaining various stereotypes, offering a potential pathway for developing effective de-biasing techniques. Moreover, scoring head biases with gradient-based metrics offers another way for identifying biased heads (Yang et al., 2023a). Furthermore, recent work has placed LLM representations under scrutiny (Zou et al., 2023a). Typically, specific templates relevant to a given concept or function are designed beforehand. Then, representations closely aligned with the concepts or functions are examined using principal component analysis (PCA). From this analysis, a vector is derived from the first principal component to predict a certain bias.
To achieve fair model predictions, a diverse range of mitigation techniques have been proposed to debias models. One stream of work proposes to debias LLMs at the embedding level, refining embeddings through various approaches. For example, a recent work attempts to alter biased embeddings with minimal alterations to make them orthogonal to neutral embeddings (Rakshit et al., 2024). Additionally, some studies concentrate on removing biases at the level of attention heads. Ma et al. (2023) address this by pruning attention heads that significantly contribute to certain biases. Similarly, another study adapts movement pruning (Sanh et al., 2020) to dynamically choose an optimal subset of low-bias attention heads (Joniak & Aizawa, 2022). Beyond modifying embeddings and pruning attention heads, another strategy involves targeting a specific group of neurons known to propagate biases. It unlearns the biases by retraining weight vectors for these neurons (Yu et al., 2023). Besides, bias mitigation can also be approached from a data-centric perspective using a few training samples (Thakur et al., 2023). This work uses a pre-trained model to find the most biased training examples, and then modifies these examples to fine-tune the model.
5.4 Toxicity
Toxicity is another form of harmful content that LLMs may produce. This issue arises because LLMs are often trained on extensive online text corpora that have not been thoroughly filtered, containing elements of toxicity that can hardly be fully eliminated. Toxicity can be identified by interpreting LLM components like the feed-forward layers and attention heads. For instance, recent work reveals how toxicity is represented within LLMs by identifying multiple vectors promoting toxicity within the MLP layers and further recognizes relevant dimensions through singular value decomposition (Lee et al., 2024). Furthermore, the exploration of geometric structures in per-layer representations offers another way to detect toxicity. Balestriero et al. (2023) applied spline formulations within MLPs to extract seven characteristic input features, demonstrating their utility in describing the domain of prompts and classifying toxic remarks.
The insights into how toxicity manifests within LLMs also shed light on mitigation strategies. Motivated by the finding that toxicity can be reduced by manipulating relevant vectors, Lee et al. (2024) develops a method called direct preference optimization (DPO), where paired toxic and non-toxic samples are utilized to fine-tune models so that non-toxic content is promoted. By examining the changes in the parameter matrices during the fine-tuning process, it substantiates that even minor adjustments to these critical vectors can reduce toxicity. Built on the observation that LLMs’ representations are updated by outputs from attention layers (Elhage et al., 2021), another work attempts to reduce toxicity by identifying the “toxicity direction” and then adjusting representations in the opposite direction (Leong et al., 2023).
5.5 Truthfulness
One prominent drawback of LLMs is their tendency to confidently produce false statements. These statements fall into two main categories: 1) statements that contradict learned knowledge within models, a problem often related to models’ honesty; 2) statements that are factually incorrect and appear to be fabricated by models, a phenomenon commonly referred to as hallucination. In the following, we delve into various approaches that aim to understand aforementioned two behaviors by leveraging explainability tools.
5.5.1 Honesty
Honesty of LLMs describes models’ ability to produce true statements based on their learned information, where dishonest behaviors significantly compromise the perceived trustworthiness of LLMs. Numerous studies have been undertaken to understand how and why dishonest behaviors happen by looking inside LLMs. One notable work attempts to distinguish dishonesty by training a classifier to predict the accuracy of statements (Azaria & Mitchell, 2023). The classifier is simply trained on top of activations from the hidden layers of LLMs. These activations are generated from true and false statements. The classifier reaches an accuracy range between 60% and 80%, which suggests that LLMs may be internally aware of the truthfulness of their outputs (Azaria & Mitchell, 2023). Furthermore, research by Campbell et al. (2023) localizes dishonesty behaviors at the level of attention heads. This study employs specially crafted prompts to induce dishonest responses, and then trains logistic classifiers on models’ activations over true/false statements. It also employs activation patching to substitute lying activations with honesty ones. Intriguingly, both approaches have witnessed the importance of layer 23--29 in flipping dishonesty behaviors. Besides, another popular method tries to study the geometric structure of true/false statements(Marks & Tegmark, 2023). Typically, these structures are visualized by projecting representations of statements onto two principal components. A clear linear structure and the truth directions can be derived to mitigate the dishonest behaviors.
5.5.2 Hallucinations
Hallucinations in LLMs can arise due to poor data quality, biases, outdated knowledge, and the lack of explicit knowledge (Xu et al., 2024; Zhu et al., 2023b). However, whether LLMs are aware of their hallucination behaviors remains an open question. Recent work investigates this question by examining models’ hidden representation space (Duan et al., 2024). It examines three hidden states involving a question, its correct answer and its incorrect answer, which are used to compute an “awareness” score. This metric quantifies the uncertainty of LLMs regarding their own answers, finding that adversarially induced hallucination can increase models’ awareness. Additionally, Li et al. (2024b) illustrates the major differences between models’ output and their inner activations, identifying these discrepancies as a potential source of hallucination. By training linear probing classifiers on each attention head’s activations, the most specialized attention head is identified. An orthogonal probe is subsequently trained with the identified specialized probe. The experiments reveal that “truth” might exist in a subspace instead of a single direction (Li et al., 2024b). Another work investigates the source of hallucination by analyzing patterns of source token contributions through perturbations (Xu et al., 2023c). Their findings suggest that hallucinations may stem from the models’ excessive dependence on a restricted set of source tokens. Besides, the static distribution of source token contribution, termed as “source contribution staticity”, can be used as another indicator of hallucinations.
Building on the above insights into LLM hallucinations, Duan et al. (2024) apply PCA to derive the direction of the correct answer’s final hidden state, and enhance the hidden representations with this direction to reduce hallucinations. In contrast, Li et al. (2024b) adopts a different approach, by intervening on top-$K$ specialized attention heads, while minimizing the influence of the rest attention heads within models. Different from PCA that identifies a single principle direction, this work adopts two distinct techniques to find multiple directions of intervention. First, they use orthogonal vectors of each probe’s hyperplane, which is similar to PCA. Second, they leverage vectors that connect the mean of the true and false distributions (Li et al., 2024b). The vectors derived from mean shift has been demonstrated more effective than those from probe classifiers, which presents another feasible strategy for identifying directions of truth.
5.6 Challenges
We discuss the challenges in employing explanations to improve models’ trustworthiness and enhance alignment from two aspects: 1) limitations of existing detection techniques, and 2) shortcomings of mitigation strategies based on explanations.
5.6.1 Challenges of Existing Detection Methods
Current detection methods primarily focus on the level of LLM modules, such as layers, attention heads, and representations. However, we still lack a finer-grained understanding of how knowledge is encoded within LLMs, such as at the level of individual neurons and weight parameters. Furthermore, we lack general and robust strategies to identify these knowledge reliably. For instance, to identify gender biases, attention heads might be examined and then the related heads pruned (Li et al., 2024b). This approach requires analyzing each model individually, rather than adopting a general approach. Moreover, existing localization approaches rely either on probing classifiers or casual scrubbing, which might not be reliable. For probing classifiers, the pre-designed biases used to train these classifiers are crucial to their performance. On the other hand, casual cleaning usually introduces new variables that complicate the analysis.
5.6.2 Challenges of Mitigation Strategies
Since LLMs are trained on vast corpora of data, it is impractical to mitigate the aforementioned trustworthiness issues from the data perspective. It is widely believed that XAI, i.e. understanding the inner mechanisms of LLMs, is a fundamental way to solve these issues. Thus, mitigation methods for LLMs are typically developed based on explanations. Existing explanations are implemented using techniques from mechanistic interpretability and representation engineering (Zhao et al., 2024). While both streams of methods can alleviate these issues, they fail to fully address them. For example, principal component analysis (PCA) is often used to find geometric structures related to these issues, but cannot fix these issues. Another popular mitigation method is casual scrubbing, which introduces certain "positive" activations to replace those demonstrated to be responsible for specific issues. However, the identified directions and patched activations can only mitigate issues to a certain extent. Moreover, the changes to either representations or activations could also influence other aspects of models’ capabilities, which we are yet unable to evaluate.
6 LLM Enhancement via Explainable Prompting
A key distinction between LLMs and traditional machine learning models lies in the LLMs’ ability to accept flexibly manipulated input data, namely prompts, during model inference (Liu et al., 2023a). LLMs generally give precedence to the information presented in these prompts when generating outputs. Therefore, to mitigate the opacity issue in LLM predictions, we can enhance prompts with understandable content, which is then prioritized over the LLMs’ inherent and implicit knowledge. These enriched prompts can include domain-specific insights, contextual information, or a step-by-step reasoning chain. In response, LLMs might reveal their decision-making processes during inference, which improves the explainability of their behaviors.
6.1 Chain of Thoughts (CoT) Prompting
The Chain of Thought (CoT) approach significantly enhances LLMs in tackling complex tasks (Wei et al., 2022). While LLMs are adept at generating human-like responses, they often lack transparency in their reasoning processes. This limitation makes it difficult for users to assess the credibility of the responses, especially for questions that require detailed reasoning.
To bridge this gap, recent efforts have incorporated in-context learning with human-crafted explanations directly into prompts (Wei et al., 2022; Huang et al., 2023a; Yao et al., 2023b; Besta et al., 2023). Among these approaches, the CoT method stands out by employing explicit knowledge to guide the reasoning process. Formally, we define the language model as $f_{\theta}$, and input prompt as $X=\{x_{1},y_{1},x_{2},y_{2},...,x_{n}\}$, where $x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1}$ denote the example question-response pairs for in-context learning, and $x_{n}$ is the actual question. In a standard question-answering scenario, we have the model output as $y_{n}=\operatorname*{arg\,max}_{Y}p_{\theta}(Y|x_{1},y_{1},x_{2},y_{2},...,x_{n})$. This approach, however, does not provide insights into the reasoning process behind the answer $y_{n}$. Therefore, the CoT method proposes to include human-crafted explanations $e_{i}$ for the $i$-th in-context example, resulting in a modified input format $X=\{x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}\}$. Given the input, the model will output not only $y_{n}$ but also the generated explanation $e_{n}$:
$e_{n},y_{n}=\operatorname*{arg\,max}_{Y}p_{\theta}(Y|x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}).$ (5)
Besides allowing for a more transparent and understandable interaction with LLMs, the CoT approach is also practically useful as it augments LLMs’ functionality by opening a window for users to control the models’ thought processes. Specifically, the usefulness of CoT methods lies in several key aspects:
  • Reducing Errors in Reasoning: By breaking down complex problems into a series of smaller tasks, CoT reduces errors in complex calculations and logic-oriented tasks, leading to a more precise resolution of intricate problems (Wei et al., 2022; Qin et al., 2023; Zhang et al., 2023; Wang & Zhou, 2024b).
  • Providing Adjustable Intermediate Steps: CoT enables the outlining of traceable intermediate steps within the problem-solving process. This feature enables users to trace the model’s thought process from inception to conclusion, and to adjust the prompts if undesirable model behaviors are observed (Lyu et al., 2023; Wang et al., 2023d).
  • Facilitating Knowledge Distillation: The step-by-step reasoning processes derived from larger LLMs can serve as a specialized fine-tuning dataset for smaller LLMs. It allows smaller models to learn complex problem-solving by following explanations, effectively teaching them to tackle intricate questions with enhanced reasoning capabilities (Magister et al., 2022).
6.2 Extended Methods of Explainable Prompting
Advanced techniques beyond CoT methods have been developed to broaden the range of reasoning paths available to LLMs towards enhancing the transparency and understandability of the decision-making process (Yao et al., 2023b; Besta et al., 2023; Yao et al., 2023a; Dhuliawala et al., 2023; Lyu et al., 2023). We introduce several notable examples below.
Tree-of-Thoughts (ToT). Proposed by Yao et al. (2023b), ToT advances beyond the traditional linear Chain of Thought reasoning, offering a more versatile structure that allows models to navigate through multiple reasoning paths. ToT makes the reasoning process of LLMs more interpretable by closely aligning it with human thought processes, as humans naturally consider multiple options and possible outcomes in both forward planning and retrospective analysis to reach conclusions (Sloman, 1996; Stanovich, 1999). This capability enhances the capacity of LLMs to tackle complex challenges that require the ability to consider and reevaluate different strategies, such as devising game strategies or generating creative content. By simulating the way humans think and make decisions, ToT not only makes their thought process more understandable to human users, but also improves the models’ effectiveness in handling complex tasks.
Graph of Thoughts (GoT). Proposed by Besta et al. (2023), GoT transforms the output of LLMs into a graph format. This format visualizes information pieces as nodes and their connections as edges, enabling a more intricate and connected form of reasoning compared to previous methods CoT and ToT. By organizing data into nodes (individual concepts or pieces of information) and edges (relationship between these concepts), GoT makes the logical connections within complex systems more understandable (Yao et al., 2023a). This graphical representation brings several benefits for understanding complex information. Firstly, it enables dynamic modification of relationships between concepts, offering a clear visualization of how changing one element affects the others. This is crucial in fields like legal reasoning (Cui et al., 2023; Boche et al., 2024), scientific research (Ding et al., 2023; Choudhury et al., 2023), and policy analysis (Chen et al., 2023c), where the inter-dependencies between various factors can be intricate and subtle. Secondly, GoT enables an assessment of the significance of each node within the graph, providing insights into which pieces of information are most critical to the task. This level of adaptability and clarity makes GoT exceptionally powerful for analyzing and navigating complex information networks.
6.3 Case Study: Is CoT Really Making LLM Inferences Explainable?
6.3.1 Background and Experimental Settings
Despite the apparent intuitiveness of the CoT prompt design, a critical question remains unanswered: Does CoT really make LLM inferences explainable? In other words, can the information provided through CoT faithfully reflect the underlying generation process of LLMs? We use multi-hop question-answering (QA) as the scenario to investigate this problem.
In QA systems, answering multi-hop questions remains a significant challenge. Instead of leveraging a single information source, multi-hop questions require synthesizing information from multiple pieces or sources of data into a coherent and logical sequence. While LLMs show good performance in single-hop QA tasks (Radford et al., 2019), their efficacy significantly declines in multi-hop situations (Tan et al., 2023; Kim et al., 2023a; Zhong et al., 2023). This discrepancy highlights the need for more advanced methods to effectively handle the intricacy of multi-hop reasoning.
CoT Prompts for Multi-hop QA. To address the above challenge, our case study applies the CoT technique. CoT relies on high-quality templates as the in-context learning prompt, where we give an example as below. Here, $[x]$ denotes the test question. The “Thoughts” following each “Question” are step-by-step problem-solving statements for the multi-hop questions. The thoughts in the templates align the generation process of LLMs with human cognitive problem-solving patterns.
Question: What is the capital of the country where Plainfield Town Hall is located?
Thoughts: Plainfield Town Hall is located in the country of the United States of America. The capital of the United States is Washington, D.C.
Answer: Washington, D.C.
Question: Who has ownership of the developer of the Chevrolet Corvette (C4)?
Thoughts: The developer of Chevrolet Corvette (C4) is Chevrolet. Chevrolet is owned by General Motors.
Answer: General Motors
Question: [x]
CoT Faithfulness for Explanation: To quantitatively measure the faithfulness of CoTs, we select fidelity as the corresponding metrics (Zhao et al., 2023b; Wachter et al., 2017):
$Fidelity=\frac{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)-\mathbbm{1}\left(\hat{\mathbf{y}}_{i}^{mislead}=\mathbf{y}_{i}\right)\right)}{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)\right)}\times 100\%,$ (6)
where $\mathbf{y}_{i}$ denotes the ground truth label, $\hat{\mathbf{y}}_{i}$ denotes the original model output with CoT, while $\hat{\mathbf{y}}_{i}^{mislead}$ denotes the model output with misleading information inserted in the "Thoughts" section. In the following, we give an example. Given the target question, the correct step-by-step thoughts should be: “Ellie Kemper is a citizen of the United States of America. The president of the United States of America is Joe Biden.” To mislead the model, we replace the thoughts with incorrect information (the underlined text) and ask the model to generate a new answer based on incorrect thoughts. If the model still generates the correct answer after the modification, we believe that the CoT information does not faithfully reflect the true process of the answer generation. On the other hand, if it generates an answer corresponding to the incorrect thoughts, then we claim the thoughts are faithful.
Question: What is the capital of the country where Plainfield Town Hall is located?
Thoughts: Plainfield Town Hall is located in the country of the United States of America. The capital of United States is Washington, D.C.
Answer: Washington, D.C.
Question: Who has ownership of the developer of the Chevrolet Corvette (C4)?
Thoughts: The developer of Chevrolet Corvette (C4) is Chevrolet. Chevrolet is owned by General Motors.
Answer: General Motors
Question: [Who is the head of state of the country where Ellie Kemper holds a citizenship?]
Thoughts: Ellie Kemper is a citizen of Croatia. The head of state in Croatia is Zoran Milanović.
Answer:
Experimental Settings. We evaluate the performance on the MQUAKE-CF dataset Zhong et al. (2023), which includes 1,000 cases for each $K$-hop questions, $K\in\{2,3,4\}$, which totally consists of 3,000 questions. Our evaluation applies various language models, including GPT-2 (Radford et al., 2019) with 1.5 billion parameters, GPT-J (Wang & Komatsuzaki, 2021) with 6 billion parameters, LLaMA (Touvron et al., 2023a) with 7 billion parameters, Vicuna-v1.5 (Chiang et al., 2023) with 7 billion parameters, LLaMA2-chat-hf (Touvron et al., 2023b) with 7 billion parameters, Falcon (Almazrouei et al., 2023) with 7 billion parameters, Mistral-v0.1 (Jiang et al., 2023) with 7 billion parameters, and Mistral-Instruct-v0.2 (Jiang et al., 2023) with 7 billion parameters. These models have demonstrated proficiency in both language generation and comprehension.
6.3.2 Experiment Results
Performance Improvement. The performance reported in Table 5 for multi-hop question answering highlights the effectiveness of the CoT across various models. GPT-J exhibits the most dramatic improvement, particularly in 3-hop questions with over a 200% increase in accuracy, suggesting that CoT’s coherent reasoning greatly enhances LLMs’ question-answering ability. While GPT-2 shows modest gains, the performance of GPT-J and other models, such as LLaMA, Vicuna-v1.5, LLaMA2, Falcon, Mistral-v0.1, and Mistral-v0.2, indicates that the CoT method may be more beneficial for more advanced models. LLaMA2 and Mistral-v0.2 display considerable improvements of 78.4% and 63.8%, respectively, in 3-hop questions, further supporting this observation. Overall, the results underscore the potential of coherent reasoning techniques, i.e., CoT, in advancing the question-answering capabilities of LLMs across different model architectures and sizes.
Table 5: Multi-hop question answering performance on MQUAKE-CF.
Datasets MQUAKE-CF
Question Type 2-hops 3-hops 4-hops
Edited Instances Base Enhanced Improve Base Enhanced Improve Base Enhanced Improve
GPT-2 (1.5B) 13.6 15.9 16.9% $\uparrow$ 11.6 8.9 23.3% $\downarrow$ 7.0 8.4 20.0% $\uparrow$
GPT-J (6B) 23.1 51.9 124.7% $\uparrow$ 10.1 30.5 202.0% $\uparrow$ 21.3 49.8 133.8% $\uparrow$
LLaMA (7B) 47.7 65.1 36.4% $\uparrow$ 29.6 39.3 32.8% $\uparrow$ 52.4 62.9 20.0% $\uparrow$
Vicuna-v1.5 (7B) 41.3 56.3 36.3% $\uparrow$ 22.7 29.7 30.9% $\uparrow$ 31.6 53.1 68.2% $\uparrow$
LLaMA2 (7B) 36.7 58.7 60.0% $\uparrow$ 17.0 30.3 78.4% $\uparrow$ 29.2 49.1 68.1% $\uparrow$
Falcon (7B) 42.3 61.7 45.7% $\uparrow$ 23.2 31.7 35.7% $\uparrow$ 33.3 48.6 45.7% $\uparrow$
Mistral-v0.1 (7B) 49.0 69.3 41.5% $\uparrow$ 30.0 42.3 41.1% $\uparrow$ 48.7 63.2 29.9% $\uparrow$
Mistral-v0.2 (7B) 44.0 56.3 28.0% $\uparrow$ 23.0 37.7 63.8% $\uparrow$ 32.9 56.2 70.9% $\uparrow$
Faithfulness Evaluation of CoT. Table 6 illustrates the impact of accurate versus misleading CoTs on the performance of LLMs. The Fidelity metric indicates how faithfully the model’s output reflects the reasoning process described in the CoT. Ideally, a high Fidelity score suggests that the model’s final response is directly based on the provided CoT, validating it as a faithful explanation of the model’s reasoning pathway. However, as we will discuss below, a low Fidelity may not always imply a lack of faithfulness in the model’s reasoning, which calls for developing more effective evaluation methods in future research.
GPT-J and LLaMA exhibit high fidelity scores across different question types, indicating a strong adherence to the given reasoning paths. Conversely, models such as Vicuna-v1.5, LLaMA2, Falcon, Mistral-v0.1, and Mistral-v0.2 show relatively high mislead accuracy scores with lower fidelity scores. In the experiments, we observe that these models usually rely on their own generated thoughts instead of using incorrect information provided in the CoT. Mistral-v0.2, in particular, demonstrates the lowest fidelity scores and highest misleading accuracy scores, suggesting a potential self-defense ability against false information. The lower fidelity scores of later models may be attributed to their improved training processes on more diverse and high-quality datasets, enabling them to develop a better understanding of context and reasoning. As a result, they are more likely to generate their own correct reasoning paths.
In conclusion, the generated thoughts can be generally viewed as faithful explanations of their output answer. While high fidelity scores generally indicate a model’s adherence to the provided CoT, low fidelity scores do not necessarily imply a lack of faithfulness, especially when the model demonstrates the ability to reject misleading information. Further research on CoT faithfulness and the development of more sophisticated evaluation metrics could contribute to the advancement of interpretable and reliable language models.
Table 6: CoT Faithfulness Evaluation on MQUAKE-CF.
Datasets MQUAKE-CF
Question Type 2-hops 3-hops 4-hops
Edited Instances Correct Mislead Fidelity Correct Mislead Fidelity Correct Mislead Fidelity
GPT-2 (1.5B) 15.9 5.2 67.3% 8.9 2.9 67.4% 8.4 1.3 84.5%
GPT-J (6B) 51.9 7.3 85.9% 30.5 1.8 94.1% 49.8 2.0 96.0%
LLaMA (6B) 65.1 9.9 84.8% 39.3 6.1 84.5% 62.9 6.0 90.5%
Vicuna-v1.5 (7B) 56.3 21.7 61.5% 29.7 12.7 57.3% 53.1 16.1 69.7%
LLaMA2 (7B) 58.7 17.0 71.0% 30.3 8.3 72.5% 49.1 12.0 75.6%
Falcon (7B) 61.7 24.0 61.1% 31.6 15.0 52.6% 48.6 23.1 52.4%
Mistral-v0.1 (7B) 69.3 24.0 65.4% 42.3 13.0 69.3% 63.2 18.4 70.8%
Mistral-v0.2 (7B) 56.3 47.9 14.8% 37.7 22.0 41.6% 56.2 37.3 33.6%
6.4 Challenges
Within machine learning, explanation faithfulness refers to the degree to which an explanation accurately reflects the decision-making process of the model it is intended to elucidate (Li et al., 2022c). An explanation is considered as faithful if it causes the model to make the same decision as the original input. In this context, the challenge faced by explainable prompting (e.g., CoT prompt) lies in two aspects: (1) directing language models to generate explanations that are genuinely representative of the models’ internal decision-making processes, and (2) preventing language models from depending on potentially biased CoT templates.
Regarding the first challenge, our case study has revealed that relatively small language models may generate answers that do not align with the provided CoT rationales. Therefore, these rationales do not accurately represent the decision-making process within these models. Some efforts have been made to bolster the CoT capabilities of smaller language models by implementing instruction tuning with CoT rationales (Kim et al., 2023b; Ho et al., 2022). These methods can help improve the explanation faithfulness of CoT for small language models, thereby addressing this issue to some extent. Nevertheless, it remains a challenging problem of how to ensure the generated explanations (i.e., “what the model says”) are faithful to the internal mechanism (i.e., “what the model thinks”) of language models.
Regarding the second challenge, recent research shows that explanations in the CoT can be heavily influenced by the introduction of biasing prompt templates into model input (Turpin et al., 2024). This is because existing CoT requires carefully designed templates to prompt language models to produce explanations. If incorrect or biased information is encoded in such templates, the generated explanations could be misleading. Recently, Wang & Zhou (2024a) propose a novel decoding strategy to implement CoT with prompting, which could mitigate this issue. However, how to effectively help language models get rid of the template reliance still remains to be underexplored.
7 LLM Enhancement via Knowledge-Augmented Prompting
Enhancing models with external knowledge can significantly improve the control and interpretability of decision-making processes. While LLMs acquire extensive knowledge through pre-training on web-scale data, this knowledge is embedded implicitly within the model parameters, making it challenging to explain or control how this knowledge is utilized during inference. Additionally, LLMs may not always encompass the unique knowledge specific to certain domains, nor keep pace with the constantly evolving information in the world. To address these limitations, this section discusses Retrieval-Augmented Generation (RAG) for the explicit integration of external knowledge into the mechanism of LLM-based AI systems, aiming to yield more interpretable predictions.
7.1 Preliminaries: Retrieval-Augmented Generation
By fetching relevant information from external databases or the internet, RAG ensures that LLM outputs are accurate and up-to-date. It addresses LLMs’ limitation of relying on fixed and potentially outdated knowledge bases. RAG operates in two steps: (1) Retrieval: It locates and fetches pertinent information from an external source based on the user’s query; (2) Generation: It incorporates this information into the model’s generated response. Given an input query $x$ and the desired output $y$, the objective function of RAG can be formulated as (Guu et al., 2020):
$\max_{\phi,\theta}\,\log p(y|x)=\max_{\phi,\theta}\,\log\sum_{z\in\mathcal{K}}p_{\phi}(y|x,z)\cdot p_{\theta}(z|x),$ (7)
where $z$ stands for the external knowledge retrieved from a knowledge base $\mathcal{K}$. Thus, the target distribution is jointly modeled by a knowledge retriever $p_{\theta}(z|x)$ and an answer reasoning module $p_{\phi}(y|x,z)$. The knowledge $z$ serves as a latent variable. An RAG model is trained to optimize the parameters, so that it learns to retrieve relevant knowledge $z$ and to produce correct answers $y$ based on $z$ and $x$. As LLMs possess stronger text comprehension and reasoning abilities, they can directly serve as the reasoning module $p_{\phi}$ without further training. In this case, RAG can be treated as a data-centric problem:
$\max_{z\in\mathcal{K}}\,\log p(y|z,x)=\max_{z\in\mathcal{K}}\,\frac{p(z|x,y)}{p(z|x)}p(y|x),$ (8)
where the goal is to find appropriate knowledge that supports the desired output. The interpretability of RAG-based models comes from the information in $z$: (1) $z$ usually elucidates or supplements the task-specific information in $x$; (2) $z$ could explain the generation of output $y$. Unlike other deep models that directly estimate $p(y|x)$ in an end-to-end manner, where the decision process is not comprehensible, the RAG process provides justification or rationale $z$ that supports the result.
Existing Retrieval-Augmented Generation (RAG) approaches can be categorized based on when they integrate external knowledge into the model’s workflow. The first category incorporates external knowledge at the inference stage. For instance, Karpukhin et al. (2020) employ dense vectors to identify related documents or text passages, enhancing the data retrieval step of RAG. Similarly, Lewis et al. (2020) refine the data retrieval process to ensure only the most pertinent information influences the model’s output. The second category integrates external knowledge during the model tuning stage. Some representative approaches include Guu et al. (2020); Borgeaud et al. (2022); Nakano et al. (2021). Generally, these methods embed a retrieval mechanism into the model’s training phase, enabling the model to utilize external data more efficiently from the outset.
7.2 Enhancing Decision-Making Control with Explicit Knowledge
The incorporation of explicit external knowledge through RAG enhances the precision and controllability of decision-making in LLMs. This method leverages real-time information from external databases to produce responses that are not only accurate but also tailored to the specific requirements of each query. Below, we explore the mechanisms by which RAG achieves a more controllable and directed content generation process, with references to key papers that have contributed to these advancements.
7.2.1 Reducing Hallucinations in Response
“Hallucination" in the context of LLMs refers to instances where these models generate information that, while coherent and contextually appropriate, is not based on factual accuracy or real-world evidence (Huang et al., 2023c). This issue can lead to the production of misleading or entirely fabricated content, posing a significant challenge to the reliability and trustworthiness of LLMs’ outputs. RAG offers a powerful solution to mitigate the problem of hallucinations in LLMs. By actively incorporating up-to-date, verified external knowledge at the point of generating responses, RAG ensures that the information produced by the model is anchored in reality. This process significantly enhances the factual basis of the model’s outputs, thereby reducing the occurrence of hallucinations. Shuster et al. (2021) applies neural-retrieval-in-the-loop architectures to knowledge-grounded dialogue, which significantly reduces factual inaccuracies in chatbots, as confirmed by human evaluations. Siriwardhana et al. (2023) introduces RAG-end2end, which joint trains retriever and generator components together. Their method demonstrates notable performance improvements across specialized domains like healthcare and news while reducing knowledge hallucination.
7.2.2 Dynamic Responses to Knowledge Updating
RAG empowers LLMs with the ability to incorporate the most current information, keeping their decision-making processes aligned with the latest developments. This feature is especially vital in fast-evolving fields such as medicine and technology, where the need for timely and accurate information is paramount (Meng et al., 2022b). For example, research by (Izacard & Grave, 2020) demonstrates significant enhancements in output relevance and accuracy through real-time information retrieval. Similarly, Han et al. (2023) suggest using retrieved factual data to correct and update the knowledge within pre-trained LLMs efficiently. Additionally, Wang et al. (2023e) introduce a method for integrating newly retrieved knowledge from a multilingual database directly into the model prompts, facilitating updates in a multilingual context.
7.2.3 Domain-specific Customization
RAG enhances LLMs by incorporating knowledge from specialized sources, enabling the creation of models tailored to specific domains. Research by Guu et al. (2020) illustrates how integrating databases specific to certain fields into the retrieval process can empower models to deliver expert-level responses, boosting their effectiveness in both professional and academic contexts. Shi et al. (2023) have applied this concept in the medical domain with MedEdit, utilizing an in-context learning strategy to merge relevant medical knowledge into query prompts for more accurate medical advice. Moreover, recent research finds that LLMs struggle to capture specific knowledge that is not widely discussed in the pre-training data. Specifically, Mallen et al. (2023) observe that LLMs often fail to learn long-tail factual knowledge with relatively low popularity, finding that simply increasing model size does not significantly enhance the recall of such information. However, they note that retrieval-augmented LLMs surpass much larger models in accuracy, particularly for questions on well-known subjects, suggesting that this method can effectively bridge knowledge gaps. Similarly, Kandpal et al. (2023) highlights LLMs’ challenges with acquiring rare knowledge and proposes that retrieval augmentation offers a viable solution, minimizing reliance on extensive pre-training for capturing nuanced, less common information.
7.3 Challenges
We discuss the challenges in RAG that are relevant to its explainability aspects: (1) In the retrieval stage $p_{\theta}(z|x)$, does the retrieved information $z$ always elucidate the task-specific information contained in the input $x$? (2) In the generation stage $p_{\phi}(y|x,z)$, does $z$ effectively serve as an explanation for the generation of output $y$? Please note that our goal is not to exhaustively discuss all the limitations of RAG in this paper as RAG itself is a broad topic in NLP research. For a more detailed examination of the broader limitations of RAG, we direct readers to other reviews (Gao et al., 2023).
7.3.1 Retrieval Accuracy Bottlenecks
Existent RAG methods typically rely on similarity search to pinpoint relevant information (Lewis et al., 2020; Gao et al., 2023), which represents a substantial improvement over basic keyword searches (Robertson et al., 2009). However, these methods may struggle with complex queries that demand deeper comprehension and nuanced reasoning. The recent “lost-in-the-middle” phenomenon (Liu et al., 2024) has revealed that an ineffective retrieval can result in the accumulation of extraneous or conflicting information, negatively affecting the generation quality. To address this challenge, recent RAG approaches have integrated adaptive learning processes (Asai et al., 2023). This advancement enables the retrieval system to refine their performance over time through feedback, adapting to evolving language use and information updates, ensuring their responses remain relevant and accurate. Nonetheless, efficiently handling intricate and multi-hop questions remains a significant challenge, highlighting the need for ongoing research to enhance the capabilities of RAG systems.
7.3.2 Controllable Generation Bottlenecks
In-context learning stands out as the premier method for incorporating external knowledge to boost the capabilities of LLMs such as GPT-4 (Asai et al., 2023; Gao et al., 2023). Despite its effectiveness, there’s no surefire way to ensure that these models consistently leverage the provided external knowledge within the prompts for their decision-making processes. In practice, to achieve thorough coverage, commonly used dense retrieval usually returns a large volume of content, including both relevant and redundant information to the input question. Unfortunately, redundant information in the model prompt raises the computational cost and can mislead LLMs to generate incorrect answers. Recent research shows the retrieved information can degrade the question-answering task performance (Yoran et al., 2023; Petroni et al., 2020; Li et al., 2022a). Some recent work proposes to fine-tune the LLM to improve resilience to noise and reduce hallucinations. However, such approach still cannot prevent oversized retrieval information decrease the system interpretability (Yoran et al., 2023; Xu et al., 2023b). The challenge of optimizing the use of external explanations to achieve more precise and controlled decision-making in LLMs is an ongoing issue that has yet to be fully addressed.
8 Training Data Augmentation with Explanation
This section explores the generation of synthetic data from explanations using large language models, a technique poised to enhance various machine learning tasks. In machine learning, limited data availability often constrains model performance, presenting a significant challenge across many domains. A viable solution is data augmentation, where LLMs, with their robust generative capabilities, can be utilized for text data augmentation (Whitehouse et al., 2023), such as transforming existing text samples into new variants (Dai et al., 2023). Nevertheless, there are several challenges to be tackled for effective text augmentation. First, for utility, the generated samples need to exhibit diversity compared to the original data. Second, these samples should be exhibit useful patterns relevant to the downstream tasks. To address these challenges, explanation methods offer a valuable tool to guide data augmentation by providing supplemental contexts and useful rationales (Carton et al., 2021). Using LLMs for explanation-guided data augmentation is a nascent but promising field. In this section, we aim to outline feasible frameworks and discuss potential applications, offering directions for future research in this field.
Explanations can be particularly beneficial in data augmentation within two scenarios. In the first scenario, explanations are used to delineate desired model behaviors or to identify existing deficiencies, which effectively guides the data augmentation process of LLMs. The second scenario involves employing LLMs to directly produce explanatory texts, which serve as supplementary information to enrich the dataset.
8.1 Explanation-guided Data Augmentation for Mitigating Shortcuts
Machine learning models are prone to make predictions with spurious correlations, also known as shortcuts (Geirhos et al., 2020), which are misaligned with human reasoning processes. This dependency on shortcuts underlies various challenges in machine learning, notably diminishing a model’s ability to generalize to out-of-distribution samples and weakening its resilience against adversarial attacks (Xu et al., 2020). The extent to which a model depends on shortcut features is intrinsically linked to its interpretability. Essentially, if a model’s predictions are predominantly based on such unreliable features, it indicates that the underlying mechanisms of the model are not fully comprehensible or interpretable from a human perspective. Therefore, post-hoc explanation techniques usually play a crucial role in detecting undesirable correlations between input and predictions within deep models (Liu et al., 2018; 2021). For example, Du et al. (2021) adopt Integrated Gradient (IG) to attribute a model’s predictions to its input features, showing that the model tends to treat functional words, numbers, and negation words as shortcuts and strongly rely on these shortcut features for prediction in natural language understanding tasks.
Data augmentation can be applied to training better models that are less sensitive to shortcut features. Explanatory information such as counterfactuals (Wang & Culotta, 2021) has been incorporated in data augmentation to improve model robustness. It generates counterfactual samples by first identifying critical features (e.g., word tokens) and then replacing these features with their antonyms, along with reversing their associated labels. Subsequently, the generated samples are combined with the original ones to train downstream models. Furthermore, these techniques can be extended to enhance the out-of-distribution performance of smaller models (Sachdeva et al., 2023; Wen et al., 2022). Namely, large language models could serve as an effective tool to augment data. For example, LLMs are able to synthesize examples that represent outlier cases or rare occurrences, which helps smaller models generalize better on unseen data (Xu et al., 2023a). This could be helpful in building robust models in scenarios where data are scarce or confidential (Tang et al., 2023a). Besides, LLMs are promising in improving models’ safety by generating adversarial examples that are more valid and natural compared to conventional approaches (Wang et al., 2023f). First, the most vulnerable words are identified with attribution-based methods. Then, these words are replaced by LLMs in a manner that maintains the integrity of the original text. The quality of these examples can be examined with an external classifier. Subsequently, these adversarial examples are employed to train downstream models, effectively fortifying them against potential attacks and boosting their security. Similarly, LLMs are also helpful in mitigating biases such as fairness issues in models (He et al., 2023). This work claims that it automatically identifies underrepresented subgroups, and chooses an optimal augmentation strategy that avoids hurting other groups. New group examples are generated using LLMs with human-providing labels. The experiments observe improvements on both underrepresented groups and overall model performance. This approach prevents skewed datasets from leading to unfair outcomes for a certain demographic, thereby potentially promoting fairness in society.
8.2 Explanation-enhanced Data Enrichment
As a powerful generative tool, LLMs have been leveraged to directly generate natural language explanations as augmented data. This strategy relies on LLMs’ understanding abilities to assist smaller models in their learning tasks. One objective in such work is to add natural language explanation generated by LLMs to training data, so as to enhance the performance of small models. Li et al. (2022b) introduce explanations from LLMs that facilitate the training of smaller models to enhance their reasoning capabilities and acquire explanation generation abilities. Specifically, three explanation generation approaches from LLM are utilized to train smaller models, including (1) explanations generated through chain of thought prompting, (2) explanations by rationalization prompting conditioned on golden labels and (3) hybrid approach that combines the first two approaches. With the best case scenario in a reasoning case, a 9.5% improvement on accuracy compared to 60x larger GPT-3 has been seen (Li et al., 2022b). It is worth noting that LLMs including ChatGPT and GPT-4 are competitive in generating coherent and logical explanations, except for limited languages (Whitehouse et al., 2023). Explanations from LLMs have also been utilized to mitigate spurious correlations in aspect-based sentiment analysis tasks (Wang et al., 2023c). This study proposes using LLMs to generate explanations for the sentiment labels of aspects in sentences (Wang et al., 2023c). These explanations provide reasoning grounded in contextual semantics rather than relying on spurious correlations between words and labels. The explanations are integrated into the training of aspect-based sentiment analysis models through two methods: augmenting the training data with the explanations or distilling knowledge from the explanations through mimicking behaviors. By focusing on the explanations instead of superficial textual cues, the models can better learn the true associations between text and sentiment and become more robust, improving both in-domain performance and generalization ability (Wang et al., 2023c). Another line of work involves integrating LLM rationales as additional supervision to guide the training of smaller models. Experiments have shown that this approach not only requires fewer training data but also outperforms traditional fine-tuning and distillation methods (Hsieh et al., 2023).
In addition to the existing application of augmentation techniques summarized above, we envision that incorporating additional text information can also be practical and efficient in enhancing the performance of various models. For example, one promising application lies in the realm of guiding the parameter learning process of small models by using automatically generated explanations. Previous research has investigated this avenue by directing the attention of natural language inference models towards human-crafted explanations (Stacey et al., 2022). As human-crafted explanations are both arduous and non-transferable, utilizing LLMs as generators presents a more economic and versatile alternation. Another potential application is to enhance model performance on complex tasks using natural language explanations from LLMs. For instance, code translation generation tasks incorporate explanations as an intermediate step, improving model performance by 12% on average (Tang et al., 2023b). The result shows that explanations are particularly useful in zero-shot settings. Apart from assisting in smaller models, LLMs have demonstrated their ability to improve their own reasoning abilities by generating reliable rationales (Huang et al., 2022). Further, Krishna et al. (2023) embed post-hoc explanations, attributing scores to all input features, into natural language rationales. This approach improves model accuracy by 10-25% across multiple tasks via in-context learning. Another study explicitly investigates LLMs’ ability to generate post-hoc explanations in natural language. The experiments show that LLMs spot key features with 75% accuracy (Kroeger et al., 2023). These studies present a novel strategy for leveraging post-hoc explanations, beyond traditional natural language explanations, to enrich training data so as to bolster model performance.
8.3 Challenges
8.3.1 Computational Overhead
Conventional post-hoc explanations, built on well-trained models, are often resource-intensive tasks. The first scenario mentioned above leverages interpretability techniques to accurately diagnose dataset issues. This process typically requires multiple rounds of model training and applying interpretability methods to develop fair and robust models. Consequently, the crafting process can be both time and energy-consuming. Given these challenges, exploring the development of data-centric evaluation metrics is crucial. These metrics can offer a more efficient way to assess data issues, bypassing traditional, cumbersome explanation methods. By focusing on these data-centric measurements, data issues can be diagnosed and fixed before training. The number of training rounds needed is then significantly reduced. This shift not only streamlines model development but also helps reduce computational overhead, making the whole process more practical and efficient.
8.3.2 Data Quality and Volume
Despite their advanced capabilities, LLMs still have limitations when dealing with highly specialized or niche contexts. For example, one of the most prominent issues is “hallucination”, where models generate plausible but incorrect or misleading responses. This could adversely affect the quality of augmented data, potentially introducing more biases to which LLMs are also vulnerable. Another challenge is controlling the relevance of LLM-generated content. That is, the explanations or data points may seem reasonable but often lack factual accuracy or nuances specific to a domain. Currently, we lack robust metrics to effectively measure the quality and relevance of these generated data relative to the original tasks. Determining the precise amount of data required is also challenging, often leading to new dataset imbalances. Managing the quality of LLM-generated data is an immense challenge, as augmented data can introduce other biases (Zhao & Vydiswaran, 2021). This stems from LLMs’ limited ability to accurately control the quantity and distribution of generated data. Moreover, crafting effective prompts is more of an art than a science, adding uncertainty around generated data quality. Together, these factors underscore the complexities and challenges in fully harnessing LLMs’ potential for data augmentation and related tasks.
9 Generating User-Friendly Explanation for XAI
Previous sections mainly focused on quantitative explanations with LLM via numerical values. For example, sample-based explanation discussed in Section 4 aims to assign each training sample an influence score (see Eqs.1-4) that measures the confidence that we can use that training sample to explain the prediction of a test sample. However, using numerical values for explanations is not intuitive, which can be difficult to understand by practitioners with little domain knowledge (Latif & Zhai, 2024; Lee et al., 2023; Li et al., 2020). User-friendly explanations, on the contrary, aim to generate human-understandable explanations, e.g., natural language-based descriptions, regarding certain data, the reason why a model makes certain predictions, or what role a neuron plays in the network, such that the generated explanations can be well-understood by both researchers and practitioners.
Given an explainee $e$, which can be a data sample $(x_{i},y_{i})$, a neuron $\theta_{i}$ from a pretrained model $f_{\theta}$, or a prediction result $\hat{y}$ based on the input $x$, generating user-friendly explanation aims to map the explainee $e$ to a sequence of natural language tokens as the explanation for the explainee $e$, such that the generated explanations can be easily comprehended by human beings.
9.1 User-friendly Data Explanation with LLMs
Data explanation refers to the process of translating difficult materials (e.g., program codes, long documents) into concise and straightforward language so that they are easy to understand by humans. Language models have long been used to generate explanations for textual data (Dai & Callan, 2019). Since modern LLMs are trained on corpora composed of codes, math, and papers, they can be leveraged to explain data beyond pure textual content. For example, Chen et al. (2021) have demonstrated that pretrained GPT models possess the ability to understand and generate codes, where explanatory comments are generated simultaneously that facilitate the understanding of programmers. In addition, Welleck et al. (2022) propose to explain math theorems by providing detailed derivations, so that the theorems are easier to understand. Recently, LLMs have also been used to elucidate academic papers (Castillo-González et al., 2022), making difficult content to be easily understood by individuals with little domain knowledge.
9.2 Explaining Small Models with LLMs
Recently, there has been growing interest in leveraging LLMs to generate free-text explanations for small models. For example, to explain black-box text classifiers, Bhattacharjee et al. (2023) propose a prompting-based strategy to identify keywords $K=\{k_{1},k_{2},...,k_{n}\}$ in the input texts $x$ with pretrained LLMs that are informative for the label $y$, and ask LLMs to substitute them with another set of keywords $K^{\prime}=\{k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{n}\}$, such that changed text $x^{\prime}$ changes the label prediction to $y^{\prime}$. They view the textual mapping rule “if we change $K$ into $K^{\prime}$ in $x$, then $y$ will be classified as $y^{\prime}$” as the counterfactual explanation for the model. In addition, to explain the neuron of a pretrained language model (e.g., GPT2), Bills et al. (2023) propose to summarize the neuron activation patterns into textual phrases with a larger language model (e.g., GPT4), where the neuron activation patterns are expressed as a sequence of (token, attribution score) pairs. To verify the identified patterns, they generate activation patterns according to the phrases via the same LLM and compare their similarity with the true activation patterns of the neuron, where the phrases with high scores are considered more confident to serve as the explanation for the neuron.
The explaining ability of LLMs is not necessarily limited to text models. For example, Zhao et al. (2023a) propose using pretrained vision-language models to generate explanations for a neuron $\theta_{i}$ of an image classification model. Specifically, for each class $y=y_{c}$, they first find regions in images with label $y_{c}$ that have maximum activation of the neuron $\theta_{i}$ as the surrogate explainees for $\theta_{i}$, and prompt LLMs such as ChatGPT to generate candidate explanations (words, short phrases) for the class label $y_{c}$. Then, they use the pretrained vision-language model CLIP (Radford et al., 2021) to match the candidate explanations with the surrogate explainees as the explanations for the neuron $\theta_{i}$. Recently, LLMs have also found applications in explaining recommender systems Zhu et al. (2023a). Specifically, Yang et al. (2023c) found that LLMs can well interpret the latent space of sequential recommendation model after alignment, whereas Lei et al. (2023) propose to align user tokens of LLMs with the learned user embeddings of small recommendation model to generate explanations of user preferences encoded in the embeddings. Recently,  Schwettmann et al. (2024) propose a unified framework to explain all models where inputs and outputs can be converted to textual strings. Specifically, the explainer LLM is used as an agent to interact with the explainee model by iteratively creating inputs and observing outputs from the model, where the textual explanations are generated by viewing all the interactions as the context.
9.3 Self-Explanation of LLMs
Due to the black-box nature of LLMs, it is promising to generate user-friendly explanations for the LLMs themselves, such that the operational mechanics and the predictions of LLMs can be well-understood by human experts. Based on whether the LLM needs to be retrained to generate explanations for themselves, the self-explanation of LLM can be categorized into two classes: fine-tuning based approach and in-context based approach, which will be introduced in the following parts.
Fine-tuning based approaches. Given sufficient exemplar explanations on the labels of the training data (e.g., in recommendation datasets such as the Amazon Review datasets (He et al., 2017) or the Yelp dataset (Zhou et al., 2020), users have provided explanations on why they have purchased certain items, which can be viewed as explanations for the ratings), LLMs can learn to generate explanations for their predictions as an auxiliary task through supervised learning. One exemplar method is P5 (Geng et al., 2022), which fine-tunes the pre-trained language model T5 (Raffel et al., 2020) on both the rating and explanation data to generate an explanation alongside the recommendations. Recently, several works have improved upon P5 (Cui et al., 2022; Zhu et al., 2024), which fine-tunes different LLMs such as GPT2, LLaMA, Vicuna, etc., and propose different prompt learning strategies (Li et al., 2023d) with generating explanation as the auxiliary task. With explanations introduced as additional supervision signals to fine-tune pretrained LLMs for recommendations, the performance can be improved with good explainability.
In-context based approaches. In many applications, there is often a lack of sufficient exemplar explanations. However, the unique capability of modern LLMs to reason and provide answers through human-like prompts introduces the potential for in-context based explanations. Here, explanations for predictions are crafted solely based on the information within the prompt. A leading approach in this domain is the Chain-of-Thoughts (CoT) prompting (Wei et al., 2022), which provides few-shot examples (with or without explanations) in the prompt and asks the LLM to generate answers after reasoning step-by-step, where the intermediate reasoning steps that provide more context for generating the final answer can be viewed as explanations. However, CoT generates reasoning first and then based on which generates predictions, where the reasoning steps can influence prediction results (Lyu et al., 2023). If explanations are generated after the prediction, since the explanation is conditioned on the predicted label, it can provide a more faithful post-hoc explanation of why the model makes certain decisions (Lanham et al., 2023). The application of in-context based self-explanation of LLMs is broad. For example, Huang et al. (2023d) explore generating zero-shot self-explanation of sentiment analysis with LLMs by directly asking them to generate explanations alongside the predictions. In addition, Huang et al. (2023a) propose a chain-of-explanation strategy that aims to explain how LLMs can detect hate speech from the textual input. Lu et al. (2022) find that CoT can generate well-supported explanations for question answering with scientific knowledge.
9.4 Challenges
9.4.1 Usability v.s. Reliability
Many existing methods rely on prompts to generate user-friendly explanations, which are not as reliable as numerical methods with good theoretical foundations. Ye & Durrett (2022b) find that the explanations by CoT may not be factually grounded in the inputs. Therefore, they believe that these explanations are more suitable as post-hoc explanations regarding why the LLM makes certain predictions (no matter whether the predictions are right or wrong). However, the validity of viewing CoT explanations as post-hoc justifications has been questioned by recent findings from Turpin et al. (2024), which uses biased datasets (e.g., the few-shot examples in the prompt always answer “A” for multiple choice questions) to show that the generated explanations may be plausible, but systematically unfaithful to represent the true reasoning process of the LLMs. This issue arises from biased reasoning steps which inadvertently skew the predictions. Thus, there’s a growing need for more theoretical scrutiny of user-friendly explanations to ensure faithfulness and credibility.
9.4.2 Constrained Application Scenarios
Currently, the utilization of LLMs to explain smaller black-box models is mainly limited to those that deal with data with rich textual information (Bhattacharjee et al., 2023; Lei et al., 2023). Although Zhao et al. (2023b) propose a strategy to explain image classifiers, the ability to match candidate textual explanations with image patterns still relies on the pretrained vision-language model CLIP. This method may not be applicable to other domains such as graph machine learning (e.g., graph neural networks), or time-series analysis (recurrent neural networks), where large pretrained models have demonstrated little progress compared to natural language processing and computer vision. Therefore, there is a compelling need to devise more versatile strategies for explaining models across a wider range of fields. This endeavor could depend on the fundamental research on combining LLM with other domain-specific tasks, such as the development of Graph-Language Models that are applicable to unseen graphs in a zero-shot manner.
10 LLMs for Interpretable AI System Design
An intriguing but challenging problem in XAI is creating model architectures or even AI systems that are inherently interpretable (Rudin, 2019), where different model components represent clear and comprehensible concepts or functionalities that are easily distinguishable from one another. Machine learning models such as support vector machines (Hearst et al., 1998) and tree-based models (Song & Ying, 2015) were classical techniques for achieving model interpretability. In the deep learning era, typical research areas in this context include concept-bottleneck models (Koh et al., 2020; Yuksekgonul et al., 2022), disentangled representation learning (Denton et al., 2017; Higgins et al., 2016), and network dissection (Bau et al., 2017; 2018). Nevertheless, under the traditional deep learning setting, the usability of these techniques remains limited because of two major challenges. First, it is difficult to define the spectrum of concepts or functionalities the model is expected to capture. Second, the efficacy of interpretable models often falls short compared to black-box models, thereby constraining their practical utility.
Large foundation models, such as large language models (LLMs) and vision language models (VLMs), provide opportunities to bridge the gap. By leveraging the common-sense knowledge embedded within them, foundation models can design interpretable architectures by providing cues that encourage creating and using the features or procedures within AI workflows. This is different from traditional deep learning pipelines where the deep models automatically discover the features during the training process, which may not end up with model components with clear meanings. Furthermore, LLMs can decompose complex tasks into simpler and collaborative sub-tasks, enhancing both the system’s interpretability and its overall performance.
10.1 Designing Interpretable Network Architectures with LLMs
Representative methods for developing interpretable deep architectures include Generalized Additive Models (GAMs) (Zhuang et al., 2021; Lou et al., 2012) and Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022). These models map inputs into a human-understandable latent space, and then apply a linear transformation from this space to the target label. For example, to build a classifier that diagnoses arthritis, we can let the model identify features such as “bone spurs” and “sclerosis”, and then use these interpretable features for the final decision. However, these approaches often require the involvement of experts to define the latent space, which can limit the learning capabilities of deep models. Some work tries to automate the discovery of semantic concepts during model training, such as by requiring independence between concepts (Higgins et al., 2016; Yu et al., 2020) or clustering data (Ghorbani et al., 2019), but they lack direct control over the outcomes and does not ensure the clarity of the concepts. One promising strategy is to utilize LLMs to provide comprehensible concept candidates. Menon & Vondrick (2022) use human language as an internal representation for visual recognition, and create an interpretable concept bottleneck for downstream tasks. By basing the decision on those comprehensible concepts, the model architecture itself is provided with better transparency. Similarly, a recent approach Labo (Yang et al., 2023b) constructs high-performance CBMs without manual concept annotations. This method controls the concept selection in bottlenecks by generating candidates from the LLMs, which contain significant world knowledge (Petroni et al., 2019) that can be explored by prompting a string prefix. Human studies further indicate that those LLM-sourced bottlenecks are much factual and groundable, maintaining great inherent interpretability for model designs. Besides the concept-based models, another promising strategy is to employ LLMs to enhance the conventional architectures that are inherently interpretable, such as GAMs and Decision Trees (DTs). Singh et al. (2023) leverages the knowledge captured in LLMs to enhance GAMs and DTs, where LLMs are only involved during the augmented model training instead of the inference process. For GAMs training, LLMs can provide decoupled embeddings for enhancement. For DTs training, LLMs are able to help generate improved features for splitting. The LLM-augmented GAMs and DTs enable full transparency, where only the summing coefficients and input key phrases are required for interpretation. With the extra information from LLMs, augmented GAMs and DTs are capable of achieving better generalization performance compared with non-augmented ones.
10.2 Designing Interpretable AI Workflows with LLM Agents
Traditional deep models are usually designed in an end-to-end manner. The internal workflows are not quite understandable to general users. By utilizing common-sense world knowledge, LLMs can break down complex problems into smaller ones and organize the workflows among them, leading to more interpretable design of AI systems (Feng et al., 2023). A recent example on interpretable AI workflow design comes from  Shen et al. (2024), where an LLM-powered agent leverages ChatGPT to integrate various off-the-shelf AI models (e.g., from Hugging Face (Jain, 2022)) to handle different downstream application tasks. In order to handle the complicated tasks in a transparent workflow, LLMs serve as a pivotal role in coordinating with external models with language mediums to harness their powers. By planning the target task, selecting candidate models, executing decomposed subtasks and summarizing responses, LLMs can help disassemble tasks based on user requests, and assign appropriate models to the tasks based on the model descriptions. Similarly, to transparentize the workflow, Liu et al. (2023d) introduces a task decomposer to analyze the user prompts and break it down into a number of subtasks for solving using LLMs. Each subtask is well managed and attributed with description, domain, inputs, and outputs. In this way, the AI systems are then capable of handling intricate user prompts with a step-by-step understandable workflow. Under the prompting paradigm, Khot et al. (2022) also employs LLMs to solve complex tasks by decomposition. Drawing inspiration from software libraries where the workflows are trackable, the decomposer and shared subtasks are designed in a modular manner. One step further, Wang et al. (2024) introduces an interactive planning approach for complex tasks, which enhances the error correction on initial LLM-generated plans by integrating plan execution descriptions and providing self-explanation of the feedback. Such interactive nature enables better workflow transparency in long-term planning and multi-step reasoning task scenarios.
10.3 Challenges
10.3.1 Planning Feasibility in Complicated Scenarios
Despite the task planning capability of LLMs, it is still challenging to be applied to certain scenarios in real-world applications due to the feasibility issues. One typical scenario is the few-shot planning cases (Guo et al., 2023), where acquiring large datasets for training is either impractical or cost-prohibitive thus making feasible planning about unseen cases from sparse exemplars extremely challenging. To better assist the interpretable designs, LLM planning needs to generalize well without extensive supervision and is expected to have the ability to integrate information from prior experiences as well as knowledge. Besides, another important scenario lies in the dynamic planning settings (Dagan et al., 2023), in which LLMs integrate feedback from the environment iteratively, letting the agent take thinking steps or augment its context with a reasoning trace. Dynamic scenarios urgently and frequently involve high computational costs resulting from the iterated invocations of LLMs, and still face challenges in dealing with the limits of the context window and recovering from hallucinations on planning.
10.3.2 Assistance Reliability with Knowledge Gaps
LLMs exhibit remarkable proficiency in encapsulating real-world knowledge within their parameters, but they resort to hallucinations and biases with high confidence when certain knowledge is missing or unreliable. Although a growing number of techniques has been proposed, such as retrieval augmentation (Guu et al., 2020), searching integration (Nakano et al., 2021) and multi-LLM collaboration (Feng et al., 2023), to expand LLM knowledge, such discrepancy in knowledge may perpetually exist owing to the continuously evolving character of human understanding (Ji et al., 2023). As a result, a crucial research challenge keeps rising, i.e., how to effectively detect and mitigate the LLM knowledge gaps from humans when employing LLMs for designs. We will need further research on evaluating and developing robust LLM mechanisms to address the knowledge-gapping problems, with the goal of helping improve LLM reliability, reducing hallucinations and mitigating biases. Furthermore, the intersections between the knowledge gaps and the safety aspects are also of great challenges to be solved, which may pose some security concerns especially when using LLMs for downstream models or workflow designs.
11 Emulating Humans with LLMs for XAI
This section discusses how LLMs can be leveraged to serve XAI by playing the role of humans. Building explainable models requires two main steps where humans are in the loop: (1) collecting a dataset with human-annotated rationales to train the models; (2) collecting human feedback on the quality of explanations produced by the models for evaluation. The significant cost and time required for human involvement raise the main challenge in scaling up this procedure. LLMs emerge as a promising solution to this challenge, thanks to their capability to emulate human reasoning and produce responses that closely resemble human-generated content. In the following, we introduce the methods that demonstrate LLMs’ ability to generate human-like annotations and feedback, contributing to the creation of explainable models.
11.1 Emulating Human Annotators for Training Explainable Models
Incorporating human-understandable rationales into model development has shown its effectiveness in enhancing both the transparency and performance of the system for various NLP tasks, such as question answering (Li et al., 2018; Wu et al., 2020a), sentiment analysis (Du et al., 2019b; Antognini & Faltings, 2021), and common sense reasoning (Rajani et al., 2019; Camburu et al., 2021). We use the term rationales to describe supportive evidence that justifies the connection between inputs and outputs (Gurrapu et al., 2023). Traditionally, the rationales are collected by leveraging human annotations (Camburu et al., 2018; Wang et al., 2019) or applying expert-designed rules (Alhindi et al., 2018; Li et al., 2018), resulting in expensive costs or limited quality. Recently, researchers in automatic annotation (Ding et al., 2022; Belal et al., 2023; Gilardi et al., 2023) have begun to explore the potential of leveraging advanced LLMs to emulate human annotators in annotating the target labels of task-specific examples. These studies found that advanced LLMs show comparable annotation qualities against average crowd human annotators on most tasks with a lower cost, pointing out the scalability of using machine-emulated annotators. Inspired by these works, some studies (Huang et al., 2023b; a) attempt to leverage advanced LLMs to collect rationales by applying the chain-of-thought technique. Specifically, researchers provide several input-rationale-output demonstrations within the input text to prompt the LLMs to generate rationale and output for an unlabeled input instance. The quality of such annotated rationales largely relies on the in-context learning capabilities of LLMs, leading to uncontrollable annotation quality on uncommon tasks. Other scholars (Yao et al., 2023a; Chen et al., 2023b; Luo et al., 2023) propose a human-in-the-loop LLM-based annotation framework based on the active-learning architecture. This framework initially collects a small seed dataset with human-annotated rationales and labels. This seed dataset is used to train an explainable classifier for this downstream task. Then, each unlabeled sample is passed through the trained explainable classifier. This is followed by a selection strategy that chooses representative samples according to metrics such as explanation plausibility, prediction uncertainty, and sample diversity. Finally, LLMs are leveraged to annotate the rationales and labels of these selected unlabeled samples. This procedure could be repeated multiple times, and the trained explainable classifier from the latest time is the final output of this framework. Compared with other methods, this approach balances the annotation quality and the cost budget in developing explainable models by using LLM-emulated annotators.
11.2 Emulating Human Feedback for Evaluating Explainable Models
The explanations generated by the explainable models could be classified into two categories: extractive and abstractive (Gurrapu et al., 2023). Extractive explanations derive directly from the input data, exemplified by attribution-based methods that emphasize specific segments of the input text. In contrast, abstractive explanations are generated in a free-form text manner, such as chain-of-thought (CoT) responses (Wei et al., 2022), offering a more nuanced interpretation. The quality of extractive explanations is typically assessed through their agreement with annotated rationales (DeYoung et al., 2020), such as accuracy, recall, and precision. However, evaluating abstractive explanations presents a significant challenge, as it is impractical to exhaustive all reasonable abstractive results comprehensively. To automatically assess abstractive explanations, early studies first collect some free-text rationales, and then apply LLMs to estimate the similarity between the explanation and the rationales (Cheng et al., 2023; Li et al., 2023b). A higher similarity between the abstraction explanation and the annotated rationales indicates a more transparent model. Recently, some researchers directly use LLMs to check the rationality of the model explanations without referring to human-annotated rationales (Miao et al., 2023; Bills et al., 2023), emphasizing the potential of emulating human feedback with advanced LLMs.
11.3 Challenges
11.3.1 Uncontrollable Credibility of Emulation
While LLMs can assist in rationale collection and explanation evaluation, their behaviors of collected results may not always match human annotators, primarily due to hallucinated responses in their unfamiliar domains (Ji et al., 2023). This issue leads to unreliable annotations or feedback, as LLMs confidently generate factually incorrect conclusions. The quality of data gathered from this process is compromised, impacting the development of XAI systems. To improve the quality of annotations and feedback, future research could focus on incorporating hallucination detection (Dhuliawala et al., 2023) and retrieval augmented generation (Ren et al., 2023) techniques. These methods could enhance the reliability of LLM outputs, making them more comparable to human-generated content in the context of XAI development.
11.3.2 Ethical Considerations in LLM Annotation
When LLM annotators keep human annotators away from subjective scenarios, such as hate speech detection (Huang et al., 2023b), LLMs also have a chance to inject unethical opinions into their annotated datasets. Although most advanced LLMs are fine-tuned to align with human values (Ouyang et al., 2022), such as being helpful, honest, and harmless, many studies have shown that this protection mechanism can be jailbroken (Wei et al., 2023; Zou et al., 2023b), causing the model to produce values-violating answers. Ensuring LLM annotators follow ethical guidelines is worth further exploration.
12 Discussion and Conclusion
XAI research is undergoing a significant transformation and experiencing rapid expansion in the era of large models. In previous sections, we have introduced XAI methodologies with an emphasis on their usability. In this final section, we provide a high-level overview of the overarching challenges that persist in the field and suggest directions for future endeavors.
  • Circumvent the interpretability-accuracy tradeoff. The advent of modern LLMs (e.g., ChatGPT) has a significant impact on this tradeoff. Traditionally, in many applications, people are willing to sacrifice a certain degree of performance for better transparency. A corresponding XAI strategy is to train and deploy an inherently interpretable model that mimics the black-box model (Che et al., 2016). However, applying this strategy to LLMs presents a challenge due to the difficulty in identifying an interpretable model that can match the performance levels of LLMs. This requires the creation of XAI strategies that can circumvent this tradeoff, where enhanced interpretability can contribute to improved accuracy. This is consistent with the goal of Usable XAI discussed in this paper.
  • Data-driven AI vs. XAI. Data-driven AI refers to developing AI models that operate based on large volumes of training data. This approach often leads to “black-box” models, as it emphasizes results over the clarity of decision-making pathways. Currently, the development of XAI techniques lags behind the advancement of LLMs because the latter easily scale up with data-driven methods – they ingest gigantic amounts of texts from the Internet to train. However, we believe that XAI might still catch up because of several opportunities. (1) We may run out of data. It was predicted that “we will have exhausted the stock of high-quality language data before 2026” ***https://www.livemint.com/mint-top-newsletter/techtalk20102023.html. Should the accumulation of more data cease to yield substantial improvements, the focus might shift towards enhancing model interpretability to leverage existing data more effectively. (2) The model is relatively stable. As the Transformer architecture of LLMs is pretty mature and stable, it will attract more attention to interpret their inner workings. (3) Leveraging LLMs for XAI. The advancement of XAI research can be accelerated if it can properly use the knowledge and human-like capabilities of LLMs.
  • The objective matters for explanation. During the transition from the era of classical machine learning (when SVMs and decision trees dominate) to the deep learning era (when convolutional and recurrent neural networks became popular), XAI techniques put a strong emphasis on achieving complete transparency within models, as if “anything less than fully transparent is not transparent”. However, as LLMs begin to match or even exceed human capabilities across various tasks, the importance of certain XAI problems shifts. For example, when Recurrent Neural Networks (RNNs) were widely adopted for text generation, we are interested in how the output is linguistically derived because RNNs often produce nonsensical sentences. Nowadays we are less interested in this for LLMs, as they are proficient in generating coherent text. Nevertheless, our focus may shift to explaining how LLMs construct output with factual information, as LLMs are prone to producing hallucinations. Similar observations exist in human cognition, which can be categorized into system-1 and system-2 styles: system-1 handles intuitive and unconscious tasks that are less explainable, while system-2 encompasses logical thinking, planning, and reasoning (Goyal & Bengio, 2022). Given the vast scale and complexity of LLMs, achieving absolute transparency across all aspects of these models appears increasingly unfeasible in the immediate future. Thus, prioritizing meaningful and feasible objectives of explanation, customized for specific tasks, becomes essential in enhancing the utility of AI systems in practical applications.
  • Evaluation remains challenging for XAI in LLMs. Traditional XAI has developed a comprehensive taxonomy of explanation problems and formats, accompanied by clear definitions for each category (Han et al., 2022; Doshi-Velez & Kim, 2017; Rudin et al., 2022). However, the established taxonomy cannot be simply grafted into the study of LLM because of two reasons. First, certain XAI challenges lose their prominence in the context of LLMs, while some approaches become too complex for practical application. Second, while XAI is becoming a common pathway of solving problems for LLMs, the exploration of LLMs’ inner mechanics has branched into various directions. For instance, there has been a notable trend towards leveraging insights from human behavior and limitations to interpret LLMs, such as whether LLMs can lie (Azaria & Mitchell, 2023), can LLMs keep secrets (Mireshghallah et al., 2023), the impact of politeness in prompts on LLMs (Yin et al., 2024b), and even how they can be "hypnotized" (Li et al., 2023e). These diverse approaches have not converged to unified methodologies in interpreting LLM behaviors, which makes it challenging for evaluation. A potential risk is the resultant explanations might give users a false sense that they accurately understand the model, especially when users attempt to shoehorn certain human knowledge or concepts to explain LLMs (Schut et al., 2023).
Conclusion.
In this paper, we hope to guide readers through a crucial yet frequently underappreciated aspect of Explainable AI (XAI) – usability. To this end, we present 10 strategies for advancing Usable XAI within the LLM paradigm, including (1) leveraging explanations to reciprocally enhance LLMs and general AI systems, and (2) enriching XAI approaches by integrating LLM capabilities. Unlocking the potential of XAI’s usability can help address various challenges in LLM such as human alignment. We also provide case studies to several critical topics, aiming to provide resources for interested developers. We further discuss open challenges at the end of each strategy, suggesting directions for future work in this evolving area.
Acknowledgement
The work is, in part, supported by NSF (#IIS-2223768, #IIS-2223769, #IIS-2310261, #DRL-2101104). The views and conclusions in this paper are those of the authors and should not be interpreted as representing any funding agencies.

### v2

@@
 Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era
-Abstract
+Abstract.
 Explainable AI (XAI) refers to techniques that provide human-understandable insights into the workings of AI models.
-Recently, the focus of XAI is being extended towards Large Language Models (LLMs) which are often criticized for their lack of transparency.
-This extension calls for a significant transformation in XAI methodologies because of two reasons.
+Recently, the focus of XAI is being extended toward explaining Large Language Models (LLMs).
+This extension calls for a significant transformation in the XAI methodologies for two reasons.
 First, many existing XAI methods cannot be directly applied to LLMs due to their complexity and advanced capabilities.
-Second, as LLMs are increasingly deployed across diverse industry applications, the role of XAI shifts from merely opening the “black box” to actively enhancing the productivity and applicability of LLMs in real-world settings.
-Meanwhile, unlike traditional machine learning models that are passive recipients of XAI insights, the distinct abilities of LLMs can reciprocally enhance XAI.
-Therefore, in this paper, we introduce Usable XAI in the context of LLMs by analyzing (1) how XAI can benefit LLMs and AI systems, and (2) how LLMs can contribute to the advancement of XAI.
+Second, as LLMs are increasingly deployed in diverse applications, the role of XAI shifts from merely opening the “black box” to actively enhancing the productivity and applicability of LLMs in real-world settings.
+Meanwhile, the conversation and generation abilities of LLMs can reciprocally enhance XAI.
+Therefore, in this paper, we introduce Usable XAI in the context of LLMs by analyzing (1) how XAI can explain and improve LLM-based AI systems and (2) how XAI techniques can be improved by using LLMs.
 We introduce 10 strategies, introducing the key techniques for each and discussing their associated challenges.
@@
 The code used in this paper can be found at: https://github.com/JacksonWuxs/UsableXAI_LLM.
-**footnotetext: Equal contribution
-1 Introduction
-Explainability holds great promise in understanding machine learning models and providing directions for improvement. In practice, users have high expectations for model explainability:
+Large Language Models, Explainable AI (XAI), Usability
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Machine learning
+1. Introduction
+Explainability holds great promise for understanding machine learning models and providing directions for improvement. In practice, users have high expectations for model explainability:
 1. Through explanation, can we know if a model works properly?
-2. Does explainability help developing better models?
-First, explanations are expected to illuminate whether a model operates in accordance with human expectations.
-For example, does the model leverage reliable evidence and domain knowledge in its decision-making?
+2. Does explainability tell us how to develop better models?
+First, explanations are expected to illuminate whether a model operates as humans expect.
+For example, does the model leverage reliable evidence and domain knowledge in its decision making?
 Does the model contain bias and discrimination?
@@
 Second, in recognition of model imperfections, we aspire for explainability to inform the development of better models.
-For example, how to adjust the behaviors of a model if we find it is using unreliable or unreasonable features in making predictions?
+For example, how to adjust the behaviors of a model if we find that it uses unreliable or unreasonable features in making predictions?
 Can we improve the performance of a model by aligning its behavior with human preferences?
 Therefore, the question arises: Have these expectations been met?
-In recent years, the body of literature on Explainable AI (XAI) has expanded rapidly to improve model transparency (Du et al., 2019a; Murdoch et al., 2019; Tjoa & Guan, 2020; Došilović et al., 2018; Rudin et al., 2022), encompassing a wide array of methods customized for different data modalities, including visual (Zhang & Zhu, 2018), textual (Danilevsky et al., 2020), graph (Yuan et al., 2022), and time-series data (Zhao et al., 2023c).
-Some literature delves into specific techniques, such as attention methods, generalized additive models, and causal models.
-Additionally, some offer reviews on general principles and categorizations or initiate discussions on evaluating the faithfulness of explanations (Yang et al., 2019).
-Despite the progress, the last mile of XAI – making use of explanations – has not received enough attention.
-In many cases, we seem to be satisfied with just acquiring explanations and their associated visualizations, sometimes followed by qualitative analysis of the model’s strengths and weaknesses.
-While these explanations can reveal a model’s imperfections, quantifying model properties (e.g., fairness, security, rationality) or taking the next concrete steps toward model improvement remains a difficult task.
-The challenges in achieving usable explainability are twofold.
-First, there is an inherent conflict between AI automation and human engagement in XAI.
-On one hand, humans need to define explainability that the model should follow, or scrutinize explanation to identify if any vulnerabilities exist in the model.
-On the other hand, the requirement for human oversight introduces substantial costs, posing challenges to the scalability and practical implementation of model debugging and improvement in AI workflows.
-Second, many of the current approaches view explainability as a purely technical matter, ignoring the needs of practitioners and non-technical stakeholders.
-Existing XAI methods are mainly developed as statistical and mathematical tools.
-However, there exists a noticeable disparity between the objectives of these tools and the expectations of practitioners across various application domains (Malizia & Paternò, 2023).
-An explanation that satisfies a technical audience might offer little value to a non-technical audience.
-While the opacity issues have not yet been fully resolved for traditional deep models (e.g., multi-layer perceptrons, convolutional and recurrent neural network), the recent advancements of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023b; Chiang et al., 2023) appear to have exacerbated the challenge we are facing.
+Since the rise of deep learning, the body of literature on Explainable AI (XAI) has expanded rapidly to improve model transparency (Du et al., 2019a; Murdoch et al., 2019; Tjoa and Guan, 2020; Došilović et al., 2018; Rudin et al., 2022), encompassing a wide array of methods customized for different data modalities, including visual (Zhang and Zhu, 2018), textual (Danilevsky et al., 2020), graph (Yuan et al., 2022), and time-series data (Zhao et al., 2023c).
+In addition, some offer reviews of general principles and initiate discussions on evaluating the faithfulness of explanations (Yang et al., 2019; Doshi-Velez and Kim, 2017).
+Despite the progress, the last mile of XAI – making use of explanations – has not received the attention it merits.
+In many cases, we seem to be satisfied with just acquiring explanations and their associated visualizations, sometimes followed by qualitative discussions of the model’s strengths and failure cases.
+However, how to quantify model properties (e.g., fairness, security, rationality) and improve models by utilizing explanations remains a difficult task.
+While the opacity issues have not yet been fully resolved for traditional deep models (e.g., multi-layer perceptrons, convolutional and recurrent neural networks), the recent advancements of Large Language Models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023b; Chiang et al., 2023) appear to have exacerbated the challenge we are facing.
 Firstly, LLMs typically possess a significantly larger model size and a greater number of parameters.
 This increased model complexity intensifies the difficulty of explaining their inner workings.
-Second, different from traditional ML models that primarily focus on low-level pattern recognition tasks such as classification and parsing, LLMs can handle more complex tasks such as generation, reasoning and question answering.
+Second, unlike traditional ML models that primarily focus on low-level pattern recognition tasks such as classification and parsing, LLMs can handle more complex tasks (Zhao et al., 2023d) such as generation, reasoning, and question answering.
 Understanding the exclusive abilities of LLMs presents novel challenges for XAI techniques.
 Considering the transformative impact of LLMs across various applications, ensuring the explainability and ethical use of LLMs has become an imminent and pressing need.
-Meanwhile, the emergent capabilities of LLMs also present new opportunities for XAI research.
+Meanwhile, the emerging capabilities of LLMs also present new opportunities for XAI research.
 Their human-like communication and commonsense reasoning skills offer prospects for achieving explainability in ways that could potentially augment or replace human involvement.
-Figure 1: The contributions and outline of this paper. We define Usable XAI in the context of LLMs with seven strategies of enhancing LLMs with XAI, and three strategies of enhancing XAI with LLMs.
+Figure 1. The contributions and outline of this paper. We define Usable XAI in the context of LLMs with seven strategies of enhancing LLMs with XAI, and three strategies of enhancing XAI with LLMs.
 Defining “Usable XAI”.
-In light of the above considerations, in the context of LLMs, we define Usable XAI which includes two aspects as follows.
-(1) Utilizing Explainability to Enhance LLM and AI Systems.
-Beyond just producing explanations or enhancing the transparency of LLMs, we explore whether these explanations can pinpoint issues for model debugging or improve the overall performance of LLMs or AI models at large, such as accuracy, controllability, fairness, and truthfulness.
-(2) Utilizing LLMs to Enhance XAI Frameworks.
+In light of the above considerations, in the context of LLMs, we define Usable XAI which includes two aspects.
+(1) Leveraging XAI to improve LLMs and AI Systems.
+Beyond just producing explanations of LLMs, we explore whether these explanations can pinpoint issues for model debugging or improve the overall performance of LLMs or AI models, such as accuracy, controllability, and trustworthiness.
+(2) Enhancing the usability of XAI through LLMs.
 The human-like communication ability of LLMs can enhance model explanations in terms of user-friendliness, by converting the numerical values into understandable language.
@@
 Contribution of this paper.
-In this paper, we investigate 10 strategies towards usable XAI techniques in the context of LLMs.
-These strategies are organized into two major categories: (1) Usable XAI for LLMs; (2) LLM for Usable XAI, as shown in Figure 1.
-Additionally, we conduct case studies to substantiate the discussion on selected techniques.
-For each strategy, we also explore the open challenges and areas that require further investigation in future work.
-  • Usable XAI for LLMs.
-We introduce how interpretation can be utilized to enhance AI pipelines, including LLMs and small models.
-First, we investigate how explanations could be utilized to diagnose and enhance LLMs in terms of utility.
-We study three types of post-hoc explanation methods, targeting LLM predictions (Section 2), LLM components (Section 3), and training samples (Section 4), respectively.
-Second, we focus on how explanations could be leveraged to scrutinize and boost model trustworthiness (Section 5), including security, fairness, toxicity, and truthfulness, which is crucial to achieving human alignment.
-Third, we discuss how explainability could guide the augmentation of data, including both inference data (i.e., prompts) and training data.
-Specifically, we discuss two strategies of crafting explainable prompts for LLMs: Chain-of-Thought prompts (Section 6) and knowledge-enhanced prompts (Section 7).
-Furthermore, we introduce leveraging LLM explanations to augment training data for improving small models (Section 8).
-  • LLM for Usable XAI.
-In this part, we investigate strategies for leveraging the advanced capabilities of LLMs to address the challenges in traditional XAI domains, thus enhancing the usability of XAI in practice.
+In this paper, we investigate 10 strategies towards usable XAI techniques in the context of LLMs, as shown in Figure 1.
+For each strategy, we also explore the open challenges that require further investigation in future work.
+  • Leveraging XAI to improve LLMs and AI Systems We introduce how interpretation can be utilized to enhance LLMs and AI models.
+We follow the LLM development pipeline in reverse order—spanning post-hoc analysis, prompt improvement, and data enhancement.
+First, we investigate how post-hoc explanations could be utilized to diagnose and enhance LLMs.
+We study three types of post-hoc explanation methods, targeting LLM predictions (Section 2), LLM components (Section 3), and training samples (Section 4), respectively, followed by how explanations could be leveraged to scrutinize and boost LLM trustworthiness (Section 5), including security, fairness, toxicity, and truthfulness.
+Second, we discuss two strategies for designing explainable prompts to guide LLM reasoning: Chain-of-Thought prompts (Section 6) and knowledge-enhanced prompts (Section 7).
+Third, we introduce the use of LLM explanations to augment training data toward improving AI models (Section 8).
+  • Enhancing XAI Usability through LLMs In this part, we investigate strategies for leveraging the advanced capabilities of LLMs to address the challenges in traditional XAI domains, thus enhancing the usability of XAI in practice.
 First, we examine ways to enhance the user-friendliness of explanations through the generative capabilities of LLMs (Section 9).
@@
 Differences between this paper and existing surveys.
-Many surveys have been conducted to examine Explainable AI (Du et al., 2019a; Tjoa & Guan, 2020; Došilović et al., 2018) or Interpretable Machine Learning (Murdoch et al., 2019).
+Many surveys have been conducted to examine Explainable AI (Du et al., 2019a; Tjoa and Guan, 2020; Došilović et al., 2018; Murdoch et al., 2019).
 This paper differs from existing work as we focus on explanation methods for large language models.
-Meanwhile, different from the existing survey (Zhao et al., 2023b) that mainly reviews explanation methods for LLMs, our paper puts an emphasis on the XAI usability in LLM studies.
-To the best of our knowledge, the most related paper to our survey is (Luo & Specia, 2024), which also discusses several aspects where explanations can improve LLM performance.
+Meanwhile, different from the existing survey (Zhao et al., 2023a; Ferrando et al., 2024) that reviews explanation methods for LLMs, our paper places an emphasis on the XAI usability in LLM studies.
+To the best of our knowledge, the most related paper to our survey is (Luo and Specia, 2024), which also discusses several aspects where explanations can improve LLM performance.
 Nevertheless, this light-weight investigation lacks a thorough examination of XAI methods (e.g., sample-based explanation, interpretable workflows, explainable prompts) and how LLMs can benefit existing XAI frameworks (e.g., data augmentation, improving user-friendliness, XAI evaluation).
 Finally, our paper contributes further by providing detailed case studies and open-sourced codes, fostering future research in applying explanations effectively within the LLM context.
-2 LLM Diagnosis via Attribution Methods
+2. LLM Diagnosis and Enhancement via Attribution Methods
 This section introduces attribution methods as post-hoc explanations for LLMs, and how we can discover model defects with attribution scores.
-We start with revisiting existing attribution methods, and then discuss which methods are still suitable for explaining LLMs.
-Since LLMs widely serve both classification and generation tasks, our discussion categorizes the attribution methods accordingly.
-After that, we explore case studies of applying attribution methods to assess LLM-generated output quality.
-Finally, we discuss future work of designing novel post-hoc explanation methods for LLMs.
-2.1 Literature Review of Attribution Methods
-The attribution-based explanation quantifies the importance of each input feature that contributes to making predictions.
-Given a language model $f$ with a prediction $\hat{y}=f(x)$ according to the $N$-words input prompt $x$, the explainer $g$ assesses the influence of input words in $x$ as $\mathbf{a}=g(x,\hat{y},f)\in\mathbb{R}^{N}$.
-Typically, the sign of $a_{n}\in\mathbf{a}$ indicates word $x_{n}$ positively or negatively influences $\hat{y}$, and a greater value of $|a_{n}|$ indicates a stronger impact.
+We start by reviewing existing attribution techniques and discussing which methods are still suitable for explaining LLMs.
+After that, we explore cases that apply attribution methods to assess LLM-generated output quality.
+Finally, we discuss future work of designing novel post-hoc explanations for LLMs.
+2.1. Attribution Methods for LLMs
+The attribution-based explanation quantifies the importance of each input feature that contributes to the prediction.
+Given a language model $f$ with a prediction $\hat{y}=f(x)$ according to the $N$-words input prompt $x$, the explainer $g$ assesses the influence of input words in $x$ as: $a=g(x,\hat{y},f)$, where $a$ stores the importance scores of words.
 In text classification, $\hat{y}$ denotes a specific class label.
 In text generation, $\hat{y}$ represents a varying length of generated text.
-Many existing attribution-based explanation methods focus on classification tasks and cannot be directly applied to the generation task.
 The primary distinction between them is that: classification is limited to a specific set of predictions, while generation encompasses an endless array of possibilities.
-For instance, in sentiment analysis, a language model can be instructed to output a number between 0 and 1 that indicates the positivity of input text by adding a linear layer and a sigmoid function on top of the language model.
-However, in the generative setting, the model can express this positivity in numerous expressions, such as “the reviewer definitely loves this movie” and “it is a strong positive movie review”.
-This distinction poses a unique challenge in adapting explanation methods from classification to generation tasks.
-In the following, we review related works based on the scenarios they are applicable to.
-2.1.1 Attributing Inputs for Label Classification
-Common attribution methods (Du et al., 2019a; Murdoch et al., 2019) developed for traditional deep models include gradient-based methods, perturbation-based methods, surrogate methods, and decomposition methods.
-We introduce the general idea and representative examples for each category, followed by the analysis of their suitability for explaining large language models.
-Perturbation-based Explanation.
-Perturbation-based methods assess the importance of input features by perturbing them and monitoring changes in prediction confidence, i.e., $a_{n}=p(\hat{y}|x)-p(\hat{y}|\widetilde{x}_{n})$, where $\widetilde{x}_{n}$ refers to the input sequence with the $n$-th feature being perturbed.
-Each feature could refer to a word (Li et al., 2016a), a phrase (Wu et al., 2020b), or a word embedding (Li et al., 2016b).
-The underlying principle is that perturbing a more important feature should result in a more pronounced alteration in the model’s prediction confidence.
-However, this method has limitations, particularly in its assumption that features are independent, which is not always the case with textual data due to word inter-dependencies.
+In the following, we review related works based on the tasks to which they are applicable.
+2.1.1. Attribution Methods for Classification
+Common attribution methods (Du et al., 2019a; Murdoch et al., 2019) for deep models include gradient-based methods, perturbation-based methods, surrogate methods, and decomposition methods.
+  • Perturbation-based explanation assesses the importance of input features by perturbing them and monitoring changes in prediction confidence, i.e., $a_{n}=p(\hat{y}|x)-p(\hat{y}|x_{/n})$, where $x_{/n}$ refers to the input sequence with the $n$-th feature being perturbed or deleted (Li et al., 2016a, b; Wu et al., 2020b).
+This approach has limitations, particularly in its assumption that features are independent, which is not always the case with textual data.
 Additionally, it is computationally intensive for explaining LLMs, requiring $N$ inferences for an input of $N$ words.
-Gradient-based Explanation.
-Gradient-based methods offer a computationally efficient approach for estimating model sensitivity to input features based on gradients $\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}$, where $\mathbf{x}_{n}$ refers to the embedding of word $x_{n}$.
-Some methods employ the $L_{2}$-norm of gradients to assess word importance (Li et al., 2016a), i.e., $a_{n}=\|\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}\|_{2}$.
-This approach only requires a single inference and one backpropagation pass.
-Some extended methods multiply the gradient with the word embedding (Kindermans et al., ; Ebrahimi et al., 2018; Mohebbi et al., 2021), i.e., $a_{n}=\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}\cdot\mathbf{x}_{n}$.
-These methods may yield explanations with limited faithfulness for deep models (Shrikumar et al., 2017), as gradients only reflect the local relationship between input variation and output variation.
-To address this, Integrated Gradients (IG) has been proposed (Sundararajan et al., 2017; Sikdar et al., 2021; Sanyal & Ren, 2021; Enguehard, 2023), which accumulates gradients as input transitions from a reference point to the actual data point.
-Nevertheless, IG entails multiple rounds of inference and backpropagation, thus significantly increasing computational demands.
-Surrogate-based Explanation.
-Surrogate-based explanation methods understand complex models by constructing a simpler model $g$ trained on $\mathcal{D}(x,\hat{y})=\{(\widetilde{x}_{k},\widetilde{y}_{k})\}_{k=1}^{K}$, where $\mathcal{D}(x,\hat{y})$ denotes a dataset constructed for the target instance $(x,\hat{y})$; $\widetilde{x}_{k}$ is usually obtained by perturbing $x$, and $\widetilde{y}_{k}=f(\widetilde{x}_{k})$.
-The surrogate model $g$, ranging from basic linear models to sophisticated decision trees, serves as a proxy to approximate the decision boundary of the target model $f$ for a specific instance $(x,\hat{y})$.
-Notable examples include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017), and TransSHAP (Kokalj et al., 2021), where the first two are designed for general deep neural networks and the last one is tailored for Transformer-based language models.
-Nevertheless, a significant limitation of them is their intensive reliance on repeated interactions with the target model, a process that is impractical for LLMs.
-Decomposition-based Explanation.
-Decomposition-based methods assign linearly additive relevance scores to inputs, effectively breaking down the model’s prediction.
-Layer-wise Relevance Propagation (Montavon et al., 2019) and Taylor-type Decomposition (Montavon et al., 2017) are well-known techniques for computing these relevance scores.
-These methods have been adapted for Transformer-based language models in various research (Voita et al., 2019; 2020; Wu & Ong, 2021).
-However, a primary challenge in implementing decomposition-based explanations is the need for tailored decomposition strategies to accommodate different model architectures.
-Although many large language models are based on the Transformer framework, there are key variations between them, such as LLaMA (Touvron et al., 2023a) and GPT (OpenAI, 2023), particularly in aspects like positional encoding strategy and feed-forward network design.
+  • Gradient-based explanation offers a computationally efficient approach for estimating model sensitivity to input features based on gradients $\frac{\partial p(\hat{y}|x)}{\partial\mathbf{x}_{n}}$, where $\mathbf{x}_{n}$ refers to the embedding of word $x_{n}$ (Li et al., 2016a; Kindermans et al., [n.
+d.]; Shrikumar et al., 2017; Ebrahimi et al., 2018; Mohebbi et al., 2021).
+This approach only requires a single inference and one backpropagation pass, achieving an efficient way to estimate the input word sensitivities.
+However, it is still expensive to compute a backward pass on such giant LLMs.
+  • Surrogate-based explanation understands the target model $f$ by constructing a simpler model $g$ trained on $\mathcal{D}(x,\hat{y})=\{(\widetilde{x}_{k},\widetilde{y}_{k})\}_{k=1}^{K}$, where $\widetilde{x}_{k}$ is usually obtained by perturbing $x$, and $\widetilde{y}_{k}=f(\widetilde{x}_{k})$ (Ribeiro et al., 2016; Lundberg and Lee, 2017; Kokalj et al., 2021).
+The surrogate model $g$, ranging from basic linear models to sophisticated decision trees, serves as a proxy to approximate the decision boundary of $f$ for the target instance $(x,\hat{y})$.
+Nevertheless, a significant limitation of this approach is its repeated interactions with the target model, which leads to huge computing requirements.
+In addition, it is impractical to find a simple and explainable model $g$ that can estimate the advanced LLMs well.
+  • Decomposition-based explanation assigns linearly additive relevance scores to inputs, effectively breaking down the model’s prediction (Montavon et al., 2019, 2017; Voita et al., 2019, 2020; Wu and Ong, 2021).
+However, a primary challenge in implementing such an approach is the need for tailored decomposition strategies to accommodate different model architectures.
 This challenge poses a limitation on the universal applicability of decomposition methods for general-purpose interpretation.
-Table 1: Time complexity analysis on different attribution methods for the generative task.
+2.1.2. Attribution Methods for Generation
+With advancements in generative AI, interpreting why an output is generated has become increasingly important.
+The explanation of LLM generation can be defined as attributing the overall confidence $p(\hat{y}|x)$ to the input $x$, where $\hat{y}$ denotes the generated response $\hat{y}=[\hat{y}_{1},...,\hat{y}_{M}]$ with $M$ words.
+An approach is to treat the text generation process as a sequence of word-level classification tasks.
+This perspective allows for the application of existing classification-based explanation techniques to assess the influence of each input word $x_{n}$ in relation to each output word $\hat{y}_{m}$.
+In particular, Wu et al.
+(2023b) proposes a gradient-based method for input-output attribution.
+Specifically, the importance score $a_{n,m}$ defined between input token $x_{n}$ to output token $\hat{y}_{m}$ can be estimated as:
+(1) $a_{n,m}=p(\hat{y}_{m}|x,\hat{y}_{1:(m-1)})-p(\hat{y}_{m}|x_{/n},\hat{y}_{1:(m-1)})\approx\frac{\partial f(\hat{y}_{m}|x,\hat{y}_{1:(m-1)})}{\partial\mathbf{E}[x_{n}]}\cdot\mathbf{E}[x_{n}],$
+where $\hat{y}_{1:(m-1)}$ denotes the first $m-1$ tokens of response, $x_{/n}$ means removing the $n$-th input token from $x$, and $\mathbf{E}[x_{n}]$ is the embedding of token $x_{n}$.
+The total contribution of each input word $x_{n}$ can be estimated by averaging the attributions $a_{n,m}$ for $m=1,...,M$ (Selvaraju et al., 2016).
+Figure 2 (left) plots an example, where each index in the Y-axis refers to an input token, while that in the X-axis is an output token.
+A higher attribution score is brighter.
+Given that not all output tokens warrant interpretation, Qi et al.
+(2024) outlines a two-step procedure where informative tokens are first identified, and then attribution is made to these tokens using gradient-based or other attribution techniques (Yin and Neubig, 2022).
+It also develops a contrastive explanation method, which aims to explain why the model predicted one target token $\hat{y}_{m}$ over an alternative token $\hat{y}^{\prime}_{m}$.
+Furthermore, a Python library has been developed to incorporate existing attribution methods for sequential generation of language models (Sarti et al., 2023).
+Time complexity is a critical consideration when applying attribution methods to language generation models like LLMs.
+In particular, the perturbation-based and gradient-based approaches may be limited by their huge demand for computing resources, while the surrogate-based and decomposition-based strategies are restricted by their limited generalizability across different model architectures.
+To emphasize the computing costs of applying attribution-based approaches for generative LLMs, we present the time complexity of several representative methods in Table 1, by assuming the model takes $N$ words as input and predicts $M$ words as output.
+It shows that existing attribution methods require a large number of forward or backward operations, emphasizing the exploration of efficient methods.
+Table 1. Time complexity analysis on different attribution methods for the generative task.
 Method Forward Backward Notes
@@
 SHAP $\mathcal{O}(2^{N})$ 0 -
-To summarize, the traditional explanation methods are not always suitable for LLMs.
-In particular, the perturbation-based and gradient-based explanations are relatively easy to extend for attributing LLM responses to the input prompts, while the surrogate-based and decomposition-based methods become significantly challenging to do so.
-Specifically, surrogate-based methods suppose that an explainable small model could approximate the decision boundary of the target model around a local example, but there are limited explainable models for the text generation task.
-Meanwhile, decomposition-based methods require designing decomposition strategies for different layers, which is challenging for big LLM architectures.
-Another primary concern is their significant demand for computing resources.
-Given an $N$-words input prompt and an $M$-words output response, we present the time complexity of several representative explanation methods in Table 1.
-It demonstrates that existing methods either require a large number of forward operations or backward operations.
-Therefore, improving the efficiency of the attribution-based explanation is an important direction for future research and development.
-2.1.2 Attributing Inputs for Text Generation
-Figure 2: An example of attribution saliency map between input and output tokens.
-The explanation of generative models can be defined as attributing the overall confidence $p(\hat{y}|x)$ to the input $x$, where $\hat{y}$ denotes the generated response $\hat{y}=[\hat{y}_{1},...,\hat{y}_{M}]$ with $M$ words.
-One method to achieve this is by treating the text generation process as a sequence of word-level classification tasks.
-This perspective allows for the application of existing classification-based explanation techniques to assess the influence of each input word $x_{n}$ in relation to each output word $\hat{y}_{m}$, resulting in a corresponding attribution score $a_{n,m}$.
-After gathering the attributions $a_{n,m}$ for $m=1,...,M$, we perform an aggregation to determine the overall contribution of each input word $x_{n}$.
-This is accomplished by aggregating the individual attributions for all output words corresponding to the input word, denoted as $a_{n}=\mathrm{Aggregate}([a_{n,1},...,a_{n,M}])$.
-The simplest approach for this aggregation is to average the attributions assigned to each input word across the different output words (Selvaraju et al., 2016).
-However, Wu et al.
-(2023) observe that attribution scores from different output words are not inherently comparable.
-For example, the attribution scores for function words (e.g., “the”, “is”, “have”) are often disproportionately larger than the scores for content words with clear semantic meaning (e.g., verbs and nouns).
-Therefore, it is necessary to normalize the scores prior to the aggregation, so that the scores $[a_{n,1},...,a_{n,M}]$ become comparable for $1\leq m\leq M$.
-Figure 2 plots the normalized scores of an example case, where each index in the Y-axis refers to an input prompt token, while that in the X-axis is an output response token.
-A greater normalized attribution score is brighter.
-In this example, the user attempts to direct the model to output information that does not exist, namely the French president in 1250.
-The model successfully realizes that this thing does not exist and refuses to answer.
-The model response can be realized as three parts, “There was no”, “president in France”, and “in 1250”.
-According to the figure, the first span is generated heavily because of the tokens “Who” and “president”, while the model uses both “France” and “1250” to respond to the second span “president in France”.
-Finally, the model emphasizes the date “1250” again by referencing the same information from the prompt.
-Overall, these explanations align with human understanding and highlight the usage of this method in the future.
-However, current research on attribution-based explaining for generative LLMs is still in its early stages, and only a limited number of methods have been proposed.
-2.2 Case Studies: Usability of Attribution Methods for LLMs
-Figure 3: A general pipeline of model diagnosis with attribution explanations.
-The attribution map offers a partial insight into the operational mechanics of LLMs (Chandrasekaran et al., 2018; Hase & Bansal, 2020; Ye & Durrett, 2022a).
-Accordingly, we propose a general pipeline that leverages attribution scores to analyze LLM behaviors, as shown in Figure 3.
-First, given the target LLM and an input prompt, we compute attribution scores of input tokens relative to the output tokens.
-Second, we extract a feature vector from the attribution map, tailored to the requirements of the diagnostic task at hand.
-Third, we train a light-weight predictor (e.g., a classifier) to diagnose whether the model behaves appropriately based on the feature vector.
-In the following, we provide case studies to illustrate how attribution scores could be utilized to assess LLM response quality (Adlakha et al., 2023).
-2.2.1 LLM Response Quality Evaluation with Explanations
-This case study explores the use of attribution-based explanations as evidence for assessing the quality of LLM-generated responses.
-Here, “quality” is specifically measured by the responses’ accuracy.
-We hypothesize that responses generated from correct rationales are likely to be more accurate.
-Our method involves comparing the model’s explanations with underlying rationales as a means to evaluate response accuracy.
-This approach could enhance the reliability of conversational systems, which is crucial for applications such as medical question answering.
-Dataset.
-We employ the Multi-Sentence Reading Comprehension (MultiRC) dataset (Khashabi et al., 2018) in this case study.
-MultiRC presents a more significant challenge than other datasets: it requires the system to answer questions based on multiple sentences from a given paragraph.
-The answers may not directly correspond to specific sentence spans, reflecting a more realistic scenario of modern LLMs.
-Each MultiRC instance is associated with a human-annotated rationale indicating which sentences are relevant for answering the question.
-Our study specifically employs its test subset with 950 samples in total.
-We consider 80% in which are the training set and we report the results evaluated on the rest samples.
-Implementation.
-We apply the Importance Density Score method (Wu et al., 2023) to estimate the importance of each word in the provided paragraph.
-Specifically, given $N$-word paragraph $X$ and the $M$-word response $\hat{Y}$, the attribution from each input word $x_{n}$ to each output word $\hat{y}_{m}$ is defined as $a_{n,m}=\frac{\partial f(y_{m}|Z_{m})}{\partial\,\mathbf{E}_{i}[x_{n}]}\cdot\mathbf{E}_{i}[x_{n}]^{\top}$, where $Z_{m}$ indicates the formatted input prompt concatenated with the first $m-1$ response word, $f$ is the language model, and $\mathbf{E}_{i}[x_{n}]$ indicates the input static embedding of $x_{n}$.
-This pairwise attribution score is then normalized as $\widetilde{a}_{n,m}=\lceil L\times a_{n,m}/\text{max}_{n^{\prime}}(a_{n^{\prime},m})\rceil$.
-Any normalized attribution $\widetilde{a}_{n,m}$ is equal or less than $b$ will is forced to 0.
-The overall attribution score on word $x_{n}$ is defined as its attribution density, i.e., $a_{n}=||[a_{n,1},...,a_{n,M}||_{1}/||[a_{n,1},...,a_{n,M}]||_{p}$.
-In our experiments, we let $L=10$, $b=2$, $p=5$, and consider Vicuna-7B-v1.1 as our language model $f$.
-By averaging these word attribution scores, we reach the importance of each sentence.
-The top-K sentences with the highest importance scores are selected as the explanation for each instance.
-The explanation is then concatenated with the output response and fed to a classifier to predict whether the output is correct or not.
-We initialize the classifier with DistillBERT-base (Sanh et al., 2019), and train it over 3 epochs with learning rate $5e^{-5}$ and weight decay $1e^{-3}$.
-To compare with this, we also consider the Full Paragraph or the human-annotated rationales as the explanation part to train the classifier.
-Following previous studies (Khashabi et al., 2018; DeYoung et al., 2019), we evaluate the accuracy of the generated responses by exactly matching all correct answers with the generated responses.
-Table 2 reports the macro precision, recall, F1, and AUC scores on the test set over five random seeds.
-Results.
-Table 2: Utilizing attribution for response quality evaluation
-Setting Precision Recall F1 AUC
-Random 49.40 51.79 49.61 49.03
-Human Rationale 68.73 66.88 67.57 73.11
-Full Paragraph 58.02 58.47 56.89 63.44
-Attribution (ours) 63.25 67.69 64.12 71.53
-In Table 2, we observe that isolating the rationales from the full context could best help the classifier identify the response quality.
-It is evident that explanations align more closely with human-annotated rationales when associated with correct responses.
-In particular, the precision and recall metrics for explanations derived from correct answers surpass those associated with incorrect responses.
-This finding serves as empirical evidence underscoring the effectiveness of attribution-based explanations in estimating the accuracy of responses generated by LLMs.
-2.2.2 Hallucination Detection with Attribution Explanations
-This case study explores a different aspect of LLM generation quality, focusing on the presence of hallucinations in LLM-generated responses.
-We show that attribution-based explanations can serve as indicators to detect LLM hallucinations.
-Hallucinations are defined as responses that contain information conflicting with or unverifiable by factual knowledge (Li et al., 2023c; Ji et al., 2023).
-For instance, if a model is asked about a fictitious entity like “King Renoit” and responds with a narrative about “The Three Musketeers”, claiming it pertains to the nonexistent king, it illustrates a hallucination.
-This tendency, particularly pronounced in instruction-tuned models, stems from their earnest endeavor to fulfill user requests.
-The issue often arises when direct commands (“tell me a story”) significantly influence the generation process, while the instruction’s subject (“about King Renoit”) is neglected.
-Based on this insight, we develop a hallucination detector according to the distribution of attribution scores over different types of prompting words.
-Dataset.
-We use the Hallucination Evaluation Benchmark (Li et al., 2024a) in this case study.
-Each instance from this dataset includes an input prompt, a response generated by ChatGPT (Ouyang et al., 2022), and the knowledge related to the response.
-Each piece of knowledge has a human annotation about whether the knowledge is valid or not.
-Each instance having at least one invalid knowledge is considered a hallucination response.
-This study focuses on 632 less confused examples from the benchmark, each of which has almost all knowledge correct or almost all wrong.
-We randomly select 80% of the samples to form a training set and the rest form the testing set.
-Implementation.
-Given a query prompt and its ChatGPT response, we aim to build a classifier to detect if the response contains hallucination.
-Since the gradients of ChatGPT is inaccessible, we apply the Vicuna-7B model as the substitute to compute the attribution scores.
-Specifically, we adopt Importance Density Scores (Wu et al., 2023) to compute the attribution scores of Vicuna.
-We then use the NLTK package to identify the part-of-speech (POS) tag of each query word.
-Finally, each query-response pair is represented with an 82-dimensional vector, where each dimension indicates the average attribution scores over a certain type of POS tagging.
-We develop a Support Vector Machine classifier based on the POS tagging attribution scores on the training set and report the precision, recall, F1 scores on the hallucination samples, as well as accuracy over all samples in Table  3.
-To compare with this, we also leverage a fine-tuned model and a prompting-based method to serve as the hallucination detector baselines.
-Results.
-Table 3: Utilizing attribution for hallucination detection on ChatGPT responses.
-Method Language Model Precision Recall F1 Accuracy
-Random - 88.41 50.34 64.11 50.59
-FacTool GPT-4 95.30 72.93 82.62 73.04
-Vectara DeBERTa-base 90.29 60.54 72.40 59.45
-AttrScore (ours) Vicuna-7B 90.15 74.21 81.36 70.20
-Mistral-7B 88.74 75.04 81.26 69.57
-In the table, we first observe that all methods have demonstrated a greater performance than the Random strategy, indicating that the proposed methods and the baselines are valid hallucination detectors.
-In addition, we observe that the proposed methods achieve a competitive performance with FacTool using GPT-4, establishing a new benchmark in this domain.
-It is worth noting that, instead of using GPT-4, the attribution score for our method is calculated using a smaller 7-billion-parameter language model.
-This demonstrates the practicality and efficiency of our approach in weak-to-strong generalization as we could diagnose large language models with smaller models.
-Future work could consider extracting more effective features and using more powerful classifiers.
-2.3 Challenges
-2.3.1 How to Identify and Explain the Semantics of Output?
-The attribution function $a_{n}=p(\hat{y}|x)-p(\hat{y}|\widetilde{x}_{n})$ is no longer faithfully attributing the model predictions in the human-interested semantic level since the model could express the same semantic meaning with various responses.
-Specifically, the model could assign a lower confidence to its original response than the new one, while both responses share the same semantic meaning.
-This is a significant difference compared with the traditional classification problem, where target label sets are manually designed so that a lower $p(\hat{y}|\cdot)$ indicates the model is less confident in predicting a specific semantic concept.
-Taking the sentiment analysis task as an example, an LLM may generate two different responses sharing the same predicted concepts, such as “it is a positive review” and “the audience thinks the movie is great”.
-Current attribution-based explanations concentrate on the literal changes in generated responses, but they do not study how the semantic meanings of these responses change.
-Therefore, they do not provide sufficient explanations of model-generated responses at a semantic level.
-In this case, the semantic level is which words of the input review lead the model to believe it is positive.
-Future work may tackle this challenge by proposing metrics to evaluate the semantic differences in responses.
-Figure 4: Review of interpretation methods for LLM components and their applications. We categorize methods according to the target LLM modules: self-attention layers and feed-forward layers.
-2.3.2 Explaining LLM Predictions Beyond Attribution
-The versatility of LLM generation will inspire a variety of explanation paradigms beyond traditional attribution methods.
-Attribution methods aim to explain model output with the contribution of input features.
-This explanation task is meaningful for conventional machine learning (ML) models whose outputs are usually individual decisions with clear formats (e.g., classification, regression, object detection).
-The decisions are highly dependent on the input features.
-However, LLMs differ from traditional ML models in two aspects.
-First, the generative process of LLMs is stochastic, e.g., we may obtain different outputs by running an LLM twice even with the same input.
-Second, LLMs encode rich knowledge within their parameters, which are independent of inputs.
-These unique properties give rise to novel explanation paradigms.
-For the first aspect, an explanation task of interest would be to understand the uncertainty of LLM generation.
-For example, researchers (Ahdritz et al., 2024; Varshney et al., 2023; Su et al., 2024) leverage the prediction perplexity to check whether the LLM is confident during generation, identifying potential errors in less confident predictions.
-Second, attributing LLM predictions to their encoded knowledge instead of input patterns could provide a new perspective.
-Some researchers (Yin et al., 2024a) propose the knowledge-boundary detection task to detect whether a model has specific knowledge for a given question.
-Nevertheless, it did not attribute the prediction to specific knowledge, so humans cannot verify the prediction process with their results yet.
-3 LLM Diagnosis and Enhancement via Interpreting Model Components
-This section discusses the XAI methods that interpret the internal components of large language models.
-Additionally, it delves into the insights these methodologies offer, which can be instrumental in refining and enhancing the design of language models.
-LLMs adopt transformers as the basic architecture, which typically comprises two types of major components: self-attention layers and feed-forward layers.
-In the following, we review the research that focuses on interpreting each of these components respectively.
-3.1 Understanding the Self-Attention Module
-A multi-head self-attention module contains multiple self-attention heads capturing different kinds of word-word relations, which are modeled with weights $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D_{1}\times D_{2}}$.
-Specifically, the relation of words $w_{i}$ and $w_{j}$ is computed as $\mathbf{A}_{i,j}\propto(\mathbf{x}_{i}\mathbf{W}_{q})\cdot(\mathbf{x}_{j}\mathbf{W}_{k})^{\top}$, where $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathbb{R}^{1\times D_{1}}$ are contextual embeddings of the words.
-The most straightforward interpretation is analyzing the attention score matrix $\mathbf{A}$ given an input sequence to study the relations between words (Vig, 2019; Hoover et al., 2020).
-In practice, these intuitive explanations would be majorly used to present case studies via visualization.
-With this strategy, Wang et al.
-(2023b) conduct case studies on in-context sentiment analysis, where they find that the label words from the in-context examples serve as anchors of the final prediction.
-In specific, these anchors aggregate information from the examples to generate informative representations from the lower layers, while the deeper layers utilize these representations to make final predictions.
-This insight motivates them to reweight the attention scores of these anchors to achieve better inference accuracy.
-Some researchers (Dar et al., 2023; Wu et al., 2023) extend this framework to globally analyze the attention weights $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ by feeding the static word embeddings of words from an interested vocabulary, instead of their contextual embeddings.
-For example, with this approach, Wu et al.
-(2023) find that instruction tuning empowers LLMs to follow human intentions by encouraging them to encode more word-word relations related to instruction words.
-On the other hand, some mathematical models are proposed to theoretically explain the self-attention mechanism, such as Sparse Distributed Memory (Bricken & Pehlevan, 2021) and Transformer Circuits (Elhage et al., 2021).
-In particular, the Transformer Circuits provide mechanistic interpretability to transformer-based models, which breaks models down into human-understandable pieces.
-Although these theoretical analyses on self-attention solids a foundation for future research, their direct application is largely underexplored.
-3.2 Understanding the Feed-Forward Module
-A feed-forward network is formalized as $\mathbf{x}^{\prime}=\sigma(\mathbf{x}\mathbf{W}_{u})\mathbf{W}_{v}^{\top}$, where $\mathbf{x}\in\mathbb{R}^{1\times D_{1}}$ is the intermediate contextual representation of an input word, $\sigma$ is a non-linear operation, and $\mathbf{W}_{u},\mathbf{W}_{v}\in\mathbb{R}^{D_{1}\times D_{3}}$ are model parameters.
-Feed-forward networks can be understood as key-value memories (Sukhbaatar et al., 2015; Geva et al., 2021), where each key or value is defined as $\mathbf{W}_{u}[d]\in\mathbb{R}^{D_{1}}$ and $\mathbf{W}_{v}[d]\in\mathbb{R}^{D_{1}}$, respectively.
-That is, each feed-forward network obtains $D_{3}$ key-value pairs, called memories.
-One simple way to interpret the semantic meaning of memory is collecting the words that could maximally activate the key or value vector of that piece of the memory (Geva et al., 2021; Dar et al., 2023), which has demonstrated strong interpretability of the extracted word lists.
-However, it is critical to be aware that the key or value vectors are polysemantic (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023), indicating that this simple approach might not provide concise explanations for each key-value pair.
-It has been shown that the word list of each key-value pair has 3.6 human interpretable patterns on average (Geva et al., 2021).
-To alleviate the limited interpretability caused by the nature of polysemantic, Wu et al.
-(2023) propose to interpret the principal components of these key or value vectors, leading to a more concise explanation for each word list, such as “medical abbreviations” and “programming tasks and actions”.
-Other work examines individual memories by measuring the changes of predictions after perturbing their corresponding activations, where it reveals that some memories encode specific knowledge (Dai et al., 2022) and some others capture general concepts (Wang et al., 2022b).
-By leveraging the explanations of key-value memories, we could locate and update the memories associated with a specific piece of knowledge to perform model editing (Dai et al., 2022; Meng et al., 2022a; b; Hase et al., 2024), i.e., modifying outdated or incorrect knowledge.
-Another usage of these weight explanations is model pruning, where LLMs can be condensed up to 66.6% of their initial parameters by exclusively maintaining redundant neurons, resulting in an approximate 1.4x increase in inference speed (Dalvi et al., 2020).
-Besides interpreting and analyzing model weights, some work studies model activations (e.g., $\mathbf{x}\mathbf{W}_{u}$ or $\mathbf{x}^{\prime}$) to interpret their functionality.
-The probing technique is the most popular way for this purpose (Belinkov et al., 2018; Tenney et al., 2018; Jawahar et al., 2019; Rogers et al., 2021), identifying whether a specific concept is encoded within the representations.
-The basic idea is developing an auxiliary classifier $g$ to map from the representations $\mathbf{x}^{\prime}$ to the interested concept space $\mathcal{C}$, such as syntax and part-of-speech knowledge, and the performance of $g$ interprets how much information encoded in $\mathbf{x}^{\prime}$ is related to the concepts in $\mathcal{C}$.
-This technique motivates developing better parameter-efficient (Chen et al., 2022), domain-specific (Das et al., 2023), and robust (Bai et al., 2021; Wang et al., 2023a) LLMs.
-Recent studies (Chen et al., 2023a; Ahdritz et al., 2024) also apply the probing method to detect the knowledge boundary of a LLM so that the hallucinated responses could be reduced.
-Some researchers (Bricken et al., 2023; Cunningham et al., 2023) point out another direction to interpret the model hidden activations, called dictionary learning, which is motivated by the assumption of superposition (Elhage et al., 2022; Sharkey et al., 2022).
-The superposition assumes that LLMs would learn an over-complete set of non-orthogonal features, thereby transcending the limitations imposed by the dimensionality of the representation space.
-Therefore, the researchers aim to reconstruct and interpret these features to understand the internals of the model.
-Practically, they develop a sparse auto-encoder $g$ to reconstruct the representations $\{\mathbf{x}_{n}\}$, which shows that humans could well interpret the learned sparse features of $g$ according to their most activation words.
-Their research shows that this method could be used for more controllable generation.
-Specifically, if forcing a sparse feature to be activated, then the language model $f$ would change its response to perform the particular behavior of that sparse feature.
-For example, given “1,2,3,4,5,6,7,8,9,10” as input, the model originally generates numbers as output.
-However, when they are forced to magnify the activations of a sparse feature called “DNA”, the model changes its output to “AGACCAGAGAGAAC”.
-In general, while the explanation techniques for feed-forward networks primarily offer insights for model development, they have also demonstrated promising applications in areas such as model editing and controllable generation.
-3.3 Challenges
-Interpreting the functionality of internal modules is still in its infancy, and we identify two challenges to be tackled in this direction.
-3.3.1 Complexity of Individual Models and Their Interactions
-The transformer-based language model contains two types of modules that collaborate based on the residual mechanism (He et al., 2016), which enables later modules to utilize, enhance, and/or discard outputs from preceding modules.
-Formally, the output of the $l$-th module is denoted as $x^{l}=f^{l}(x^{l-1})+x^{l-1}$, where $f^{l}$ could be a self-attention module or a feed-forward network.
-Research in this area aims to interpret how different modules $f^{i}$ and $f^{j}$ work together for $i\neq j$.
-Pilot studies (Elhage et al., 2021; Olsson et al., 2022) find that stacked self-attention modules could form Induction Heads, which demonstrate a strong correlation with the in-context learning capability.
-Specifically, the induction head encourages the model to predict the word “B” followed by a sequence “AB…A”.
-Their study finds a specific phase during pre-training LLMs where both induction heads and the in-context learning capability emerge from the model.
-Following this track, researchers observe diverse functional heads within LLMs for different tasks, such as “Name Mover Head” and “Duplicate Token Head” for the object identification task (Wang et al., 2022a), “Single Letter Head” and “Correct Letter Head” for the multiple-choice question answering task (Lieberum et al., 2023), and “Capitalize Head” as well as “Antonym Head” for the general purpose tasks (Todd et al., 2023).
-Although these studies have indeed deepened our understanding of cross-module effects, their analyses are grounded on specific tasks or scenarios, leaving it uncertain whether these findings possess broad generalizability.
-3.3.2 Nature of Polysemantic and Superposition Assumption
-Interpreting the functionality of a single neuron (one row/column vector of a weight matrix) fails in analyzing large language models since a single neuron could be activated by multiple and diverse meanings, called polysemantic (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023).
-This nature leads to poor interpretability: explaining a single neuron usually does not reflect a concise human concept.
-Some researchers (Elhage et al., 2022; Sharkey et al., 2022) assume that this phenomenon is caused by the superposition of an over-complete set of features learned by the models.
-Based on this assumption, we may reach another level of explanation by decomposing the model weights to reconstruct a large number of features.
-However, the two critical problems of this approach are still unclear: (1) How do we ensure our reconstructed features faithfully represent the latent features learned by the models?
-(2) How do we interpret our reconstructed features with human language?
-4 LLM Debugging with Sample-based Explanation
+Another category of attribution approaches leverages the generative capabilities of LLMs to provide citations to one or more text passages that support the output they generate.
+The attribution can be obtained by explicitly prompting LLMs to quote the curated sources of information (Weller et al., 2024).
+The cited contents can originate either from external knowledge sources or from the LLMs’ own memory.
+For instance, Sun et al.
+(2023) introduces RECITE, which first retrieves one or more relevant passages from the LLMs’ memory via sampling and then generates the final response based on these passages.
+Conversely, many systems assume the presence of an external retrieval corpus.
+In these cases, the LLM output includes both an answer string and a reference to a short segment of text from the corpus that substantiates the answer (Bohnet et al., 2022; Gao et al., 2023c; Asai et al., 2023a).
+Nevertheless, LLMs may hallucinate and produce fabricated information.
+To address this, various benchmarks have been developed to assess the faithfulness of LLM-generated attributions, i.e., whether the generated statements are fully supported by the cited references.
+For details on evaluation methods, we refer readers to Section 11.
+2.2. Usability of Attribution Methods for LLMs
+Explanations provide a lens to evaluate whether a model’s predictions are grounded in reasonable rationale and to identify opportunities for improving prompt instructions.
+Based on this, we discuss two scenarios where LLMs can benefit from attributions: model diagnosis and model enhancement.
+  • Model Diagnosis.
+By comparing the LLM’s output with attribution results, it is possible to detect ungrounded or unsupported information.
+Formally, let $x$ denote the input, $\hat{y}=f(x)$ be the target model’s output, and $a$ be the attribution.
+The goal is to construct a detector $g_{\mathrm{detect}}(a,x,\hat{y})\in\{-1,+1\}$ where $+1$ indicates that the prediction can be trusted, while $-1$ means the opposite.
+Here $g_{\mathrm{detect}}$ can be implemented as an LLM.
+For instance, Gao et al.
+(2023a) proposes a framework called RARR, which leverages attribution to identify unsupported or misleading content generated by LLMs.
+In particular, the attribution step is automated by using a query generator to produce comprehensive questions about different aspects of the text that need verification.
+In addition, Agrawal et al.
+(2024) proposes ”consistency checks”, which asks the LLM to perform attribution multiple times for the same output.
+Inconsistencies in the attribution results serve as indicators of potential issues with the factuality of the LLM’s output.
+  • Model Enhancement.
+Attribution results can serve as additional information to refine prompts, thereby improving LLM performance.
+For instance, Sun et al.
+(2023) proposes ”recitation” as a new two-step approach where an LLM $f$ first explicitly recalls relevant factual knowledge $a$ from its own memory before generating an answer for input $x$.
+The LLM inference is augmented as $\hat{y}=f(x,a)$ instead of the standard $f(x)$.
+This recitation step allows the model to better access and verify its stored knowledge, similar to how a student might first recite relevant facts they’ve learned before answering an exam question, leading to more accurate responses.
+In addition, Krishna et al.
+(2024) proposes a framework called AMPLIFY, which uses attribution results to improve LLM performance by crafting effective prompts.
+It selects samples in the validation set that were misclassified by the LLM, and uses another model (e.g., GPT-2 or BERT) to compute their attribution scores with respect to the ground truth labels.
+These samples and their corresponding explanations are used to construct prompts for future test samples.
+Figure 2. Left: An example of attribution saliency maps. Right: The pipeline of making use of attribution results to improve LLMs.
+2.3. Challenges
+We discuss two major challenges regarding attribution methods for explaining LLMs, including efficiency and faithfulness.
+First, according to our analysis in Table 1, many traditional attribution methods are computationally expensive when adapted to LLMs.
+Furthermore, these methods were primarily developed for tasks in the context of classification or regression, while the more urgent need lies in designing attribution methods tailored for LLM generation.
+The diverse and dynamic nature of LLM generation demands the development of novel explanation paradigms that go beyond existing frameworks.
+Second, while some approaches utilize the generative capabilities of LLMs to produce free-text attributions, the faithfulness of these attributions remains underexplored.
+Ensuring that the generated explanations truly reflect the model’s reasoning process is a challenge to be tackled for building trust and reliability in LLM explanations.
+{forest}
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, rectangle, draw, font=, rounded corners, align=center, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, text centered, , [Interpreting Model Internals,ver,fill=gray!20,text centered,text width=10em [Explanation of
+Model Components,ver,fill=orange!40,text centered,text width=7em [Importance Analysis on
+Single Component,fill=orange!10,text centered,text width=6em [Contextual Impacts: (Geva et al., 2022), (Ferrando et al., 2023), (Wang et al., [n.
+d.]), (Elhage et al., 2021), (Meng et al., 2022a), (Wang et al., 2022a), (Mohebbi et al., 2023)., leaf2, align=left] [Global Impacts: (Geva et al., 2021), (Dar et al., 2023), (Wu et al., 2023b)., leaf2, align=left] ] [Circuit Analysis on
+Multiple Components,fill=orange!10,text centered,text width=6em [(Cammarata et al., 2020), (Wang et al., 2022a), (Hanna et al., 2023), (Geiger et al., 2021), (Conmy et al., 2023) (Sarti et al., 2023), (Hanna et al., [n.
+d.]), (Marks et al., 2024)., leaf2, align=left] ] [Usability,fill=orange!10,text centered,text width=6em [Knowledge Editing: (Dai et al., 2022), (Wang et al., 2022b), (Elhage et al., 2021), (Meng et al., 2022a), (Meng et al., 2022b), (Hase et al., 2024), (Li et al., 2024a),
+(Gupta et al., 2024), (Wu et al., 2023a), (Wang et al., 2024c)., leaf2, align=left] [Model Pruning: (Dalvi et al., 2020), (Men et al., 2024), (Gromov et al., 2025), (Dalvi et al., 2020), (Fan et al., 2024)., leaf2, align=left] ] ] [Explanation of
+Latent Representations,ver,fill=orange!40,text centered,text width=7em [Non-Training
+based Methods,fill=orange!10,text centered,text width=6em [Projecting $\mathbf{x}^{l}$ to Output Embeddings: (nostalgebraist, 2020), (Jastrzebski et al., 2018), (Belrose et al., 2023), (Din et al., 2024)., leaf2, align=left] [Feeding $\mathbf{x}^{l}$ as Input Embeddings: (Chen et al., 2024b), (Dumas et al., 2024), (Pan et al., 2024)., leaf2, align=left] ] [Supervised Training
+based Methods,fill=orange!10,text centered,text width=6em [Identifying Concept Existence: (Hewitt and Manning, 2019),(Chen et al., 2021), (Liu et al., 2019), (Hewitt and Manning, 2019), (Burns et al., 2022),
+(Zou et al., 2023a), (MacDiarmid et al., 2024),(Zheng et al., 2024), (Hewitt and Liang, 2019), (Pimentel et al., 2020), (Voita and Titov, 2020)., leaf2, align=left] [Measuring Concept Importance: (Kim et al., 2018), (Wei et al., 2021), (Bai et al., 2022), (Nicolson et al., 2024)., leaf2, align=left] ] [Unsupervised Training
+based Methods,fill=orange!10,text centered,text width=6em [Orthogonal Assumption on Concepts: (Millidge and Black, 2022), (Wu et al., 2023b)., leaf2, align=left] [Sparse Assumption on Concepts: (Olshausen and Field, 1997), (Cunningham et al., 2023), (Bricken et al., 2023)., leaf2, align=left] ] [Usability,fill=orange!10,text centered,text width=6em [Data Pre-processing: (Ge et al., 2024), (Xu et al., 2023c), (Li et al., 2024c), (Yang et al., 2025b)., leaf2, align=left] [Training-Time Regularization: (Yin et al., 2024a), (Wu et al., 2025b)., leaf2, align=left] [Inference-Time Control: (Wu et al., 2025c), (Huang, 2024), (Stolfo et al., 2024), (Rimsky et al., 2024), (Arditi et al., 2025), (Bhattacharjee et al., 2024),
+(Zou et al., 2023a), (Rimsky et al., 2024), (Chen et al., 2023a), (Ahdritz et al., 2024), (Singh et al., 2024), (Zou et al., 2023a), (Scialanga et al., 2025)., leaf2, align=left] ] ] ]
+Figure 3. Taxonomy of explanation methods for LLM internals, including explaining model components and hidden representations.
+3. LLM Diagnosis and Enhancement via Interpreting Model Internals
+This section discusses XAI methods that interpret LLM internals.
+We delve into the insights these methodologies offer, which can be instrumental in refining and enhancing the design of language models.
+Specifically, we review research that focuses on interpreting model components and latent representations, respectively, as shown in Figure 3.
+3.1. Explanation of Model Components
+LLMs adopt transformers as their basic architectures, which typically comprise two types of major components: Self-attention Mechanism (ATT) and Feed-forward Network (FFN).
+Researchers have focused on analyzing individual components as well as their interactions to gain insights into how LLMs make their predictions.
+We start by reviewing the connection between these components, focusing on how information is transformed through the residual stream.
+We then discuss methods to analyze the contribution of single or multiple components, respectively.
+We conclude by summarizing how these methods can be applied to improve model performance.
+3.1.1. Linear Relation between Components
+We consider a decoder-only LLM $f$ with $L$ internal layers, each of which is either an ATT or FNN layer, denoted as $f^{l}$.
+There is an input embedding matrix $\mathbf{W}_{\text{e}}$ that encodes each word of input text $x$ into a $D$-dimensional vector, i.e., $\mathbf{x}^{0}=\mathbf{W}_{\text{e}}[x]\in\mathbb{R}^{|x|\times D}$.
+The internal representation at each layer is computed by $\mathbf{x}^{l}=f^{l}(\mathbf{x}^{l-1})+\mathbf{x}^{l-1}$.
+Finally, there is an output embedding matrix $\mathbf{W}_{\text{u}}$ that decodes the hidden representations for next word prediction, i.e., $f(x)=\mathbf{x}^{L}\cdot\mathbf{W}_{\text{u}}$.
+By decomposing the final prediction (Elhage et al., 2021), we have
+(2) $f(x)~{}=~{}\underbrace{\mathbf{x}^{0}\,\mathbf{W}_{\text{u}}}_{\text{Residual Stream}}~{}+~{}\sum_{l=1}^{L}\underbrace{f^{l}\Bigl{(}\mathbf{x}^{l-1}\Bigr{)}\,\mathbf{W}_{\text{u}}}_{\text{Layer Update}}.$
+Equation (2) demonstrates two critical insights in understanding LLM mechanism: (1) Model prediction is a sum of layer outputs (Mickus et al., 2022).
+That is to say, the final prediction is an additive combination of the initial input embedding and the contribution from each layer.
+This linear transformation allows us to directly trace how each component shapes the overall prediction.
+Please note that the residual mechanism does not necessarily lead to a linear structure of model hidden spaces.
+For example, ResNet (He et al., 2016) does not show such a linear structure because there are non-linear operations applied to its residual stream.
+(2) The residual stream serves as a communication channel across layers (Elhage et al., 2021).
+In particular, each layer first extracts information from the residual stream, and then injects its processed output back into it.
+This shared residual stream enables multiple layers to work collaboratively, indirectly contributing to specific tasks by continuously updating the shared residual stream.
+3.1.2. Importance Analysis on Single Component
+Inspired by the first insight of Equation (2), early works (Geva et al., 2022; Ferrando et al., 2023; Wang et al., [n.
+d.]) quantify the contextual impact of a single layer $f^{l}$ in predicting a certain word $w$ under the context $x$ by computing $f^{l}(\mathbf{x}^{l-1})\,\mathbf{W}_{\text{u}}[:,w]$, where a greater value indicates a stronger impact.
+Following this trend, some researchers (Elhage et al., 2021; Meng et al., 2022a) propose to more precisely quantify the contextual impact by conducting interventions on layer outputs, i.e., $\text{diff}(\mathbf{h}\,\mathbf{W}_{\text{u}}[:,w],\widetilde{\mathbf{h}}\,\mathbf{W}_{\text{u}}[:,w])$, where $\text{diff}(\cdot,\cdot)$ measures the difference of two inputs, $\mathbf{h}=f^{l}(\mathbf{x}^{l-1})$, and $\widetilde{\mathbf{h}}$ is a intervention of $\mathbf{h}$.
+In practice, the choices of $\widetilde{\mathbf{h}}$ can vary, such as a constant vector (Wang et al., 2022a; Mohebbi et al., 2023), a noise vector (Meng et al., 2022a), and a hidden representation of other context (Hanna et al., 2023).
+On the other hand, researchers (Geva et al., 2021; Dar et al., 2023; Wu et al., 2023b) extend this framework to analyze the global impact of a single layer $f^{l}$ in predicting word $w$ by computing $\mathbf{W}^{l}_{\text{o}}\cdot\mathbf{W}_{u}[:,w]$, where $\mathbf{W}^{l}_{\text{o}}$ is the output weight matrix of layer $f^{l}$ that maps the processed results back to the residual stream.
+3.1.3. Circuit Analysis on Multiple Components
+Motivated by many emerging abilities of LLMs, it is interesting to explore how different layers work collaboratively to perform such complex tasks, called circuit analysis (Cammarata et al., 2020).
+Specifically, the goal of circuit analysis is to identify a subset of layers, whose combination performs a particular function, such as “indirect objects identification” (Wang et al., 2022a) and “number comparison” (Hanna et al., 2023).
+Recall that the residual stream enables the communication across layers, researchers (Geiger et al., 2021) formalize the forward pass of an LLM as a directed acyclic graph, where each node is one particular layer, and an edge is a representation flow from one layer to another via the residual stream.
+Under this setting, a circuit is a sub-graph with a distinct function.
+Identifying a specific circuit consists of three steps (Conmy et al., 2023): (1) selecting an interested task and constructing a dataset, (2) building the model’s graph, and (3) measuring the importance of sub-graphs with interventions (Sec.
+3.1.2).
+This circuit identification process is time-consuming as it needs a large number of forward passes for interventions.
+To overcome this challenge, researchers (Sarti et al., 2023; Hanna et al., [n.
+d.]; Marks et al., 2024) propose a gradient-based method to approximate the intervention results with one more backward pass for each sample.
+3.1.4. Usability of Component Analysis
+The success of component analysis enables researchers to isolate the entire LLM into different parts for specific tasks.
+Therefore, the direct application of this technique is either modifying (enhancing or weakening) or removing certain parts that are essential or redundant to downstream tasks, resulting in two common applications: Knowledge Editing and Model Pruning.
+  • Knowledge Editing.
+Component analysis can be used to locate the exact position storing a certain piece of knowledge (Dai et al., 2022; Wang et al., 2022b), and then update it accordingly to control the minimum effect on other knowledge.
+Specifically, we aim to update an outdated knowledge $t=(s,r,o)$ to a new one $t^{\prime}=(s,r,o^{\prime})$.
+Given the prefix $[s,r]$ of $t$, ROME (Meng et al., 2022a) first identifies a specific layer $f^{l*}\in f$ whose activation maximally contributes to predicting the original object $o$.
+This is done by maximizing the difference $f(o|[s,r])-f(o|[s,r],do(\mathbf{h}^{l}=\mathbf{h}^{l}+\epsilon))$, where $\mathbf{h}^{l}$ denotes the output activation for input $[s,r]$ at layer $f^{l}$, and $\epsilon$ denotes a random noise to simulate a perturbation.
+ROME then modifies only the parameters of this identified layer $f^{l*}$ to update the prediction to the new object $o^{\prime}$, without affecting unrelated knowledge stored elsewhere.
+The key limitation of ROME is that it can update only one piece of knowledge at a time, leading to an unacceptable time complexity to update a batch of knowledge.
+To overcome this challenge, follow-up works (Meng et al., 2022b; Hase et al., 2024; Li et al., 2024a; Gupta et al., 2024) scale up this process by skipping the knowledge locating step in ROME, and instead, they focus on learning the representation of updated knowledge at each layer.
+This line of work has been further applied to edit user personal information for privacy (Wu et al., 2023a), and edit out harmful concepts for safety (Wang et al., 2024c).
+  • Model Pruning.
+Component analysis allows researchers to identify the specific roles of each model component for a given task, and thus, redundant or irrelevant components can be pruned to speed up the inference.
+Dalvi et al.
+(2020) directly measure the functionality of $f^{l}\in f$ by monitoring the expected difference $\mathbb{E}_{x,y\in\mathcal{D}}[p(y|x;f)-p(y|x,do(f^{l}(l)=\mathbf{0});f)]$, where $\mathcal{D}$ is a collected dataset for the downstream task.
+Recent studies (Men et al., 2024; Gromov et al., 2025) directly measure the redundancy by quantifying the similarities of hidden representations from different layers.
+Empirical studies (Dalvi et al., 2020; Fan et al., 2024) found that even over half of the model parameters are redundant for a certain task, and dropping them off can significantly reduce the inference time without sacrificing performance.
+3.2. Explanation of Latent Representations
+Many researchers have studied latent representations of LLMs (i.e., $\mathbf{x}^{l}$) to interpret model behaviors.
+It is critical to recognize that the representations are not naturally interpretable because of their polysemantic nature (Arora et al., 2018; Scherlis et al., 2022; Bricken et al., 2023), indicating that each dimension of the latent space represents multiple semantic meanings, called concepts.
+Thus, we cannot clearly understand a latent representation by reviewing the values in each dimension.
+There are various lines of work to tackle this challenge, and we categorize them according to whether they require additional training.
+3.2.1. Non-Training based Methods
+Both the input and output layer of transformer-based LLM $f$ are directly interpretable as they operate on human-readable text.
+Hidden representations across different layers share the same residual stream (see Section 3.1.1), allowing information to flow seamlessly between layers.
+Consequently, non-training based interpretation methods leverage this insight by transferring the hidden representation $\mathbf{x}^{l}$ to either the first or the last layer to facilitate interpretation.
+  • Projecting $\mathbf{x}^{l}$ to Output Embeddings.
+Given a hidden representation $\mathbf{x}^{l}$ at layer $l$, Logit Lens (nostalgebraist, 2020) understands its semantics by projecting it to the output word embedding matrix $\mathbf{W}_{\text{u}}$.
+Specifically, we interpret $\mathbf{x}^{l}$ by collecting the $K$ words whose output embeddings could maximally activate $\mathbf{x}^{l}$, i.e.,
+    (3) $\mathcal{I}=\underset{\mathcal{V}^{\prime}\subset\mathcal{V},|\mathcal{V}^{\prime}|=K}{\arg\max}\sum_{w\in\mathcal{V}^{\prime}}\mathbf{x}^{l}\cdot\mathbf{W}_{\text{u}}[:,w],$
+    where $\mathcal{V}$ is a pre-defined vocabulary set.
+The Logit Lens can be viewed as an instance of skipping all later layers after $\mathbf{x}^{l}$ (Jastrzebski et al., 2018).
+However, researchers (Belrose et al., 2023; Din et al., 2024) find that there is a distribution shift between the intermediate layer $l$ and the last layer $L$, and thus, they propose to train a translator to learn the distribution shift to improve its interpretability for representations from shallow layers.
+  • Feeding $\mathbf{x}^{l}$ as Input Embeddings.
+In contrast with moving hidden representation $\mathbf{x}^{l}$ to the output layers, Self-IE (Chen et al., 2024b) proposes feeding $\mathbf{x}^{l}$ as an input word embedding in another round of model inference, where we prompt LLM to explain the meaning of the fed representation.
+For example, we may instruct LLM to interpret $\mathbf{x}^{l}$ with the prompt “Please summarize [X] in one sentence.”, where “[X]” is a placeholder, whose embedding will be replaced with our interested hidden representation $\mathbf{x}^{l}$.
+Since LLMs are trained to follow human instructions, they are expected to generate a sentence to describe the information encoded in $\mathbf{x}^{l}$.
+3.2.2. Supervised Training based Methods
+The supervised training-based interpretation methods break the polysemantic nature of LLM latent space by leveraging some annotated datasets.
+They focus on exploring two main research questions: (1) whether the LLM latent space encodes a specific concept; and if so, (2) whether this concept impacts LLM’s predictions and how significant.
+  • Identifying Existence of Certain Concepts.
+Probing (Gupta et al., 2015) is the most traditional technique in analyzing the knowledge encoded within hidden representations of LLMs.
+In specific, the Probing technique requires an annotated dataset to train a probe $p$ with parameter $\theta$ to predict the occurrence of interested knowledge, i.e., $p:\mathbb{X}\rightarrow[0,1]$, where $\mathbb{X}$ is the latent space of LLMs at a certain layer.
+Here, the dataset consists of positive samples (with label “1”) that clearly demonstrate our interest concept and negative samples that clearly exclude such concept.
+For example, early works (Belinkov et al., 2017) use this technique to identify some morphology concepts (e.g., part-of-speech tags and morph tags) in the LLM’s latent space by constructing a dataset where each example has human annotations on these tags.
+If the probe $p$ achieves a significantly high accuracy on a certain dataset, then we could conclude that the latent representation encodes that concept.
+This technique has been used to study diverse concepts, ranging from low-level linguistic patterns (Belinkov et al., 2017; Hewitt and Manning, 2019; Chen et al., 2021; Liu et al., 2019; Hewitt and Manning, 2019) to high-level semantic meanings (Burns et al., 2022; Zou et al., 2023a; MacDiarmid et al., 2024; Zheng et al., 2024).
+The improvements in the probing technique (Hewitt and Liang, 2019; Pimentel et al., 2020; Voita and Titov, 2020) mainly focus on designing baselines to confirm that the high accuracy achieved by probe $p$ is led by the knowledge provided by the latent space $\mathbf{x}$.
+  • Measuring Importance of Certain Concepts.
+To quantify how much a specific concept impacts the model’s predictions, researchers often use TCAV (Kim et al., 2018).
+We first train a linear probe $p$ with parameter $\theta\in\mathbb{R}^{D}$ to distinguish latent representations containing concept $c$ or not, where $\theta$ is named as the concept activation vector (CAV).
+TCAV measures how sensitive the model’s output is to perturbations along this concept direction (e.g., via directional derivatives of the model’s logits).
+Intuitively, if the model’s predictions change substantially when the representation is shifted along the learned concept direction, we can conclude that the concept has a high importance for the model’s decision.
+Mathematically, the perturbation is realized by computing the gradient of $f(x)$ with respect to $\mathbf{x}^{l}$ and projecting it onto $\theta$ as $\Delta=\nabla_{\mathbf{x}^{l}}f(x)\cdot\theta$, so that a larger $|\Delta|$ indicates that small shifts in the latent space along $\theta$ lead to substantial changes in $f(x)$, underscoring the concept’s influence.
+3.2.3. Unsupervised Training based Methods
+Breaking the polysemantic constraint for interpretation can be formalized as an unsupervised learning problem on the basis vectors of the LLM latent space (Fel et al., 2023), where each basis vector refers to a clear and concise semantic concept, indicating a monosemantic nature.
+Mathematically, given a total of $N$ hidden representations $\mathbf{X}\in\mathbb{R}^{N\times D}$, we aim to learn two matrices $\mathbf{A}\in\mathbb{R}^{N\times C}$ and $\mathbf{C}\in\mathbb{R}^{C\times D}$ by minimizing the loss
+(4) $\underset{\mathbf{A},\mathbf{C}}{\arg\min}\,\,\|\mathbf{X}-\mathbf{A}\mathbf{C}\|^{2},$
+where $\mathbf{C}$ serve as the learned concepts, and $\mathbf{A}$ is the coefficient of the instances on each concept.
+There are some additional conditions to ensure that the learned concept vectors satisfy our expectations on monosemantic.
+  • Orthogonal Assumption on Concepts.
+Early works (Millidge and Black, 2022; Wu et al., 2023b) applies the Singular Vector Decomposition to learn a set of concept vectors $\mathbf{C}$ that are perfectly orthogonal, i.e.,
+    (5) $\mathbf{C}\mathbf{C}^{\top}=\mathbf{I},$
+    where $\mathbf{I}$ is an identical matrix.
+However, satisfying this hard request is computing costly in practice, leading to limited scalability in interpreting hidden representations from giant LLMs with both large $N$ and $D$.
+  • Sparse Assumption on Concepts.
+Sparse Autoencoder (Olshausen and Field, 1997) (SAE) is another alternative that relaxes the orthogonal condition to be a smoother one in which the sample coefficients matrix is asked to be sparse, i.e.,
+    (6) $\mathbf{A}=\text{Top-K}(\mathbf{X}\cdot\mathbf{C}^{\top}),$
+    where Top-K activation enforces all values to be zero unless the $K$ largest ones from the inputs.
+Specifically, researchers (Cunningham et al., 2023; Bricken et al., 2023) train a SAE $g(\mathbf{x})=\text{Top-K}(\mathbf{x}\cdot\mathbf{W}^{\top})\cdot\mathbf{W}$ to reconstruct hidden representations $\mathbf{x}$ by minimizing $\|\mathbf{x}-g(\mathbf{x})\|^{2}$, where $\mathbf{W}\in\mathbf{R}^{C\times D}$ and Top-K activation enforces all values to be zero unless the $K$ largest ones from the input.
+This approach has shown effectiveness in decomposing the latent representations from giant LLMs with a hundred billion parameters, such as GPT-4 (Gao et al., 2024) and Claude 3 (Templeton et al., 2024).
+Once these concept vectors are obtained, their representing meanings can be further described with natural language by selecting $M$ input texts that could maximally activate the related concept vectors (Bricken et al., 2023), i.e.,
+(7) $\mathcal{I}_{c}=\arg\max_{\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|=M}\sum_{x\in\mathcal{X}^{\prime}}f^{l}(x)\mathbf{C}[c].$
+However, the latest research (Wu et al., 2025c; Gur-Arieh et al., 2025) has shown that such an input-based explanation cannot well interpret their impacts on controlling LLM responses, and thus, they propose to interpret these learned concept vectors by building the connection to the output texts.
+We refer audiences to check (Shu et al., 2025) for a more detailed review of designing and interpreting SAEs for LLMs.
+3.2.4. Usability of Interpreting Latent Representations
+Latent space interpretations improve LLMs further at each stage of their entire life-cycles.
+In particular, we consider three stages of model development, including the “pre-processing” stage for data preparation, the “training” stage that trains LLM on prepared data, and the “inference” stage that uses a trained model to operate user requests.
+  • Data Pre-processing.
+Existing studies have empirically found that the same model but trained on different datasets with different diversity (Ge et al., 2024), complexity (Xu et al., 2023c), and quality (Li et al., 2024c) levels may lead different downstream task performance.
+The latent space interpretation techniques enable researchers to understand the geometric structure of a dataset and, further, to select a subset that preserves a certain geometric structure.
+For example, to obtain a subset of the dataset preserving as diverse topics as possible, SAE-GreedSelect (Yang et al., 2025b) greedily select examples to maximize the total number of activated SAE features.
+Results show that a selected subset with only 5% of total samples can train a model with comparable performance to a full-trained one, significantly reducing the training cost.
+  • Training-Time Regularization.
+Either the trained parameter $\theta$ of probe $p$ or each column vector $\mathbf{C}[c]$ learned by SAE or SVD indicates a feature in the LLM latent space.
+Therefore, these learned feature vectors can be used to perform feature engineering in the LLM latent space for model training.
+For example, Yin et al.
+(2024a) propose a feature-level constrained fine-tuning approach.
+This method effectively prevents the fine-tuned model from changing significantly from its pretrained counterpart without explicitly involving the pretrained model during fine-tuning, thereby substantially reducing computational costs.
+In contrast, Wu et al.
+(2025b) introduce a more semantically meaningful regularization approach for model training.
+They first determine whether each learned feature vector $\mathbf{C}[c]$ is causally relevant to the downstream task by analyzing its natural language explanation $\mathcal{I}_{c}$.
+Subsequently, they impose a penalty during fine-tuning to discourage reliance on features identified as task-irrelevant (i.e., shortcuts).
+This approach effectively enhances the generalizability of the fine-tuned model.
+  • Inference-Time Control.
+As the learned feature vectors by probing or SAEs break the polysemantic nature and achieve monosemantic, it is possible to directly steer LLM latent space toward a specific direction without sacrificing other abilities.
+Formally, let $\mathbf{z}$ denote the learned feature vector to an interested concept, we can erase it from the hidden representation of input $x$ by $\mathbf{x}^{\prime}=\mathbf{x}-\text{ReLU}(\mathbf{x}\cdot\mathbf{z}^{\top})\cdot\mathbf{z}$, or constantly amplify it to a certain level $\alpha$ by $\mathbf{x}^{\prime}=\mathbf{x}+\alpha\cdot\mathbf{z}$.
+The combination of these two techniques on different categories of feature vectors leverages model steering for different purposes.
+For example, Wu et al.
+(2025c) observe that certain feature vectors semantically correlate with specific safety-related concepts, such as “First Aid,” based on summarizing their natural language explanations $\mathcal{I}_{c}$.
+By enforcing a constant activation of these safety-related feature vectors during model generation, they demonstrate that the model reliably switches from generating harmful responses to explicitly rejecting inappropriate or malicious user inputs.
+Generally, this technique has shown success in steering LLMs during inference to improve their helpfulness (Huang, 2024; Stolfo et al., 2024; Rimsky et al., 2024; He et al., 2025), safety (Arditi et al., 2025; Bhattacharjee et al., 2024; Zou et al., 2023a; Rimsky et al., 2024), faithfulness (Chen et al., 2023a; Ahdritz et al., 2024), fairness (Singh et al., 2024; Zou et al., 2023a), and currentness (Scialanga et al., 2025).
+3.3. Challenges
+3.3.1. Explanation Efficiency
+The most significant challenge in analyzing the inner workings of LLMs is their demand for a large amount of computing resources.
+Circuit analysis (Section 3.1.3) is one of the examples that is bounded by this bottleneck.
+Specifically, performing circuit analysis on one sample still requires a large amount of forward inference of LLMs under different perturbed internals of that sample, leading to limited scalability to analyze over a large dataset.
+Therefore, existing works on circuit analysis focus on studying particular behaviors of LLMs, such as “Name Mover Head” and “Duplicate Token Head” for object identifications (Wang et al., 2022a), “Single Letter Head” and “Correct Letter Head” for multiple-choice question answering (Lieberum et al., 2023), and “Capitalize Head” as well as “Antonym Head” for writing (Todd et al., 2023).
+On the other hand, although SAEs (Section 3.2.3) have shown strong promise in understanding LLM latent representations, training such proxy models requires collecting LLM representations over a large corpus.
+For example, researchers (Gao et al., 2024) found that training an SAE with only 1 million features requires over 10 billion training tokens.
+Future works may explore computing-efficient methods to understand LLMs’ diverse behaviors comprehensively.
+3.3.2. Explanation Faithfulness
+Although many proposed methods have shown strong promise with certain usability cases (Section 3.1.4 and Section 3.2.4), some researchers still hold concerns about whether they faithfully interpret the inner workings of LLMs since they may not practically outperform some trivial baselines.
+For example, although many researchers (Wu et al., 2025c; Bricken et al., 2023) can apply SAEs (Section 3.2.3) to steer LLM responses for specific purposes, some research (Wu et al., 2025a; Durmus et al., 2024; Bricken et al., 2024) has observed that SAE-based methods cannot significantly outperform prompting baselines, posing a serious concern about the effectiveness of SAEs.
+Furthermore, other researchers (Heap et al., 2025; Paulo and Belrose, 2025) suspect that SAEs cannot faithfully explain LLMs’ internals, as they observe that SAEs can even learn features from randomized LLMs.
+These concerns suggest two critical questions for future works: (1) How do we ensure our learned features faithfully represent LLM internals?
+(2) In which cases are SAE-based methods more effective than prompting strategies?
+4. LLM Debugging with Sample-based Explanation
 In this section, we discuss sample-based explanation strategies for LLMs, which aim to trace back the answers generated by LLMs to specific training samples (i.e., documents) or document fragments in the corpora.
@@
 First, tracing back the predictions of LLM to the training samples can provide evidence for the generation results, which facilitates model debugging in cases of errors and increases the trustworthiness of the model from users when the outcomes are accurate.
-In addition, it can also help researchers understand how LLMs generalize from training samples.
+Second, it can also help researchers understand how LLMs generalize from training samples.
 If the outputs of LLMs can be traced back to exact subsequences directly spliced from the training data, it might suggest that the LLM is simply memorizing the data.
-In contrast, if the generation results and the influencing training samples are abstractly related, it could indicate that LLMs can understand the concepts and generate responses by reasoning from input prompts.
-In this section, we start by systematically reviewing traditional sample-based explanation strategies, including gradient-based methods and embedding-based methods, as well as some preliminary explorations to generalize them to LLMs.
-We then analyze the challenges associated with generalizing the above strategies to LLMs with unique transformer structures and unprecedented numbers of parameters.
-Finally, we discuss the insights to address the challenges, as well as open challenges worthy of further investigation.
-4.1 Literature Review of Sample-based Explanation
-In this section, we denote the input space and output space as $\mathcal{X}$ and $\mathcal{Y}$, respectively.
-In the context of large language models (LLMs), $\mathcal{X}$ is the space of token sequences known as the prompts, and $\mathcal{Y}$ could be the space of discrete labels in classification tasks or the space of token sequences as output in generation tasks***Here, we should note that, in the pretraining phase of LLM with language modeling, the model either predicts masked tokens in $x_{i}$ (i.e., masked language modeling), or autoregressively predicts the next token (i.e., causal language modeling) in $x_{i}$.
-Therefore, in some works, $y_{i}$ is omitted, and only $x_{i}$ is included for discussions..
-Accordingly, we have a training dataset $\mathcal{D}_{train}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{N}$ with $N$ samples drawn from the joint space $\mathcal{X}\times\mathcal{Y}$, on which an LLM model $f_{\theta}$ is trained with pretrained parameters $\hat{\theta}\in\mathbb{R}^{P}$.
-We also have a test sample $z=(x,y)$ of interest, where we want to explain the generation of $y$ from $x$ based on training samples in $\mathcal{D}_{train}$ (which can be viewed as the information source).
-The goal of sample-based explanation is to measure the influence of a training sample $z_{i}\in\mathcal{D}_{train}$ or a certain segment within $z_{i}$, such that the generation of LLMs can be well-explained and backed up by the selected training samples.
-4.1.1 Influence Function-based Methods
-One strategy to quantify the influence of a training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ to a test sample $z$ is through the influence function (Koh & Liang, 2017; Han et al., 2020).
-It measures the change of the prediction loss $\mathcal{L}\left(z,\theta\right)$ for the test sample $z$, when the training sample $z_{i}$ undergoes a hypothetical modification in the dataset $\mathcal{D}_{train}$ during model training.
-This modification results in an altered set of optimal model parameters, denoted as $\hat{\theta}_{-z_{i}}$.
-The most common modification of a training sample is to remove it from the dataset, where the influence of the removal of a training sample $z_{i}$ on the loss at test sample $z$ can be computed as follows:
-$\mathcal{I}(z_{i},z)=-\nabla_{\theta}\mathcal{L}(z,\hat{\theta})^{\top}\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}(z_{i},\hat{\theta}),$ (1)
-where $\nabla_{\theta}\mathcal{L}(z,\hat{\theta})$ is the gradient of the loss function $\mathcal{L}$ on the test sample $z$ evaluated at the optimal parameters $\hat{\theta}$, and $\mathbf{H}_{\hat{\theta}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}^{2}\mathcal{L}(z_{i},\hat{\theta})$ denotes the Hessian matrix of the LLM model at parameter $\hat{\theta}$.
-If we denote the number of parameters in $\hat{\theta}$ as $P$, the naïve inversion of the Hessian matrix $\mathbf{H}_{\hat{\theta}}$ leads to $\mathcal{O}(NP^{2}+P^{3})$ time complexity and $\mathcal{O}(P^{2})$ space complexity (Schioppa et al., 2022), which is clearly infeasible for large models.
-To improve efficiency, Koh & Liang (2017) adopt an iterative approximation process, i.e., LiSSA (Linear time Stochastic Second-Order Algorithm), to calculate the Hessian-Vector Product (HVP) in Eq.
+4.1. Literature Review of Sample-based Explanation
+4.1.1. Influence Function-based Methods
+One strategy to quantify the influence of a training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ to a test sample $z$ is through the influence function (Koh and Liang, 2017; Han et al., 2020), which measures the change of the prediction loss $\mathcal{L}\left(z,\theta\right)$ for $z$, when the training sample $z_{i}$ is removed from the dataset:
+(8) $\mathcal{I}(z_{i},z)=-\nabla_{\theta}\mathcal{L}(z,\hat{\theta})^{\top}\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}(z_{i},\hat{\theta}),$
+where $\nabla_{\theta}\mathcal{L}(z,\hat{\theta})$ is the gradient of loss $\mathcal{L}$ on $z$, and $\mathbf{H}_{\hat{\theta}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}^{2}\mathcal{L}(z_{i},\hat{\theta})$ is the Hessian matrix.
+To improve efficiency, Koh and Liang (2017) adopt an iterative approximation process to calculate the Hessian-Vector Product (HVP) in Eq.
 (8), where the memory complexity can be reduced to $\mathcal{O}(P)$ and time complexity to $\mathcal{O}(NPr)$ ($r$ is the number of iterations).
-To further reduce the complexity, Pruthi et al.
-(2020) propose an alternative to Eq.
-(1), i.e., TracIn, which measures the influence of $z_{i}$ on $z$ by calculating the total reduction of the loss on $z$ whenever $z_{i}$ is included in the minibatch during model training.
-The TracIn measurement can be formulated as follows:
-$\mathcal{I}_{\operatorname{TracIn}}\left(z_{i},z\right)=\sum_{t:z_{i}\in\mathcal{B}_{t}}\mathcal{L}\left(z,\theta_{t}\right)-\mathcal{L}\left(z,\theta_{t+1}\right)\approx\frac{1}{b}\sum_{t:z_{i}\in B_{t}}\eta_{t}\nabla_{\theta}\mathcal{L}\left(z_{i},\theta_{t}\right)\cdot\nabla_{\theta}\mathcal{L}\left(z,\theta_{t}\right),$ (2)
-where $\mathcal{B}_{t}$ is the $t$-th mini-batch fed into the model during training, $\theta_{t}$ is the parameter checkpoint at the $t$-th step, $\eta_{t}$ is the step size, and $b$ is the size of the mini-batch.
-According to the above equation, TracIn only leverages gradient terms, where Hessian $\mathbf{H}_{\hat{\theta}}$ is removed from the influence measurement.
-This substantially improves the efficiency.
-However, such complexity is still prohibitive for large models from both the computational and memory perspectives.
-In addition, TracIn can only estimate the influence of adding/removing the sample to the loss, where variants of the vanilla influence function defined in Eq.
-(1) can measure the influence of other modifications of the training sample $z_{i}$, such as perturbation (e.g., masking out a segment of a document $x_{i}$).
-To adapt the vanilla influence function of Eq.
-(1) to explain transformers, Schioppa et al.
-(2022) propose to use Alnordi iteration (Arnoldi, 1951) to find the dominant eigenvalues and eigenvectors of the Hessian matrix on randomly sampled subsets $\mathcal{D}_{sub}$, with $|\mathcal{D}_{sub}|\ll|\mathcal{D}_{train}|$.
-In such a case, the diagonalized Hessian can be cheaply cached and inverted, where the computational and memory complexity can be substantially reduced.
-Previous work mainly focuses on reducing the complexity of calculating the influence of a single training sample.
-Observing that finding the most influential training sample on $z$ needs to iterate Eq.
-(1) overall $N$ training samples, Guo et al.
-(2021) propose to use fast KNN to pre-filter a small subset of influence-worthy data points from $\mathcal{D}_{train}$ as candidates to explain small pretrained language models, whereas Han & Tsvetkov (2022) propose to iteratively find a small subset $\mathcal{D}_{sub}\subset\mathcal{D}_{train}$ whose gradient is the most similar to that of the downstream task examples.
-Recently, Grosse et al.
-(2023) propose to use the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation to scale influence functions to LLMs up to 52B parameters.
-For adaptation, only influences mediated by the multi-layer perceptron (MLP) layers are considered, whereas weights from the self-attention layers are fixed, as previous work has demonstrated that knowledge is mainly encoded in the MLP layers (Wang et al., 2023d).
-In addition, based on the assumption that weights from different MLP layers are independent, the EK-FAC approximated influence can be formulated as the sum of influences mediated by each layer:
-$\mathcal{I}_{\text{EKFAC}}(z_{i},z)=\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(z,\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta}),$ (3)
+To further reduce the complexity, (Pruthi et al., 2020) propose TracIn, which measures the influence by calculating the total reduction of loss whenever $z_{i}$ is included in the minibatch during model training as follows:
+(9) $\mathcal{I}_{\operatorname{TracIn}}\left(z_{i},z\right)=\sum_{t:z_{i}\in\mathcal{B}_{t}}\mathcal{L}\left(z,\theta_{t}\right)-\mathcal{L}\left(z,\theta_{t+1}\right)\approx\frac{1}{|B_{t}|}\sum_{t:z_{i}\in B_{t}}\eta_{t}\nabla_{\theta}\mathcal{L}\left(z_{i},\theta_{t}\right)\cdot\nabla_{\theta}\mathcal{L}\left(z,\theta_{t}\right),$
+where $\mathcal{B}_{t}$ is the $t$-th mini-batch, $\theta_{t}$ is the parameter at the $t$-th step, $\eta_{t}$ is the step size.
+TracIn only leverages gradients, which substantially improves the efficiency.
+In addition, Schioppa et al.
+(2022) propose to use Alnordi iteration (Arnoldi, 1951) to find the dominant eigenvalues and eigenvectors of $\mathbf{H}_{\hat{\theta}}$ on randomly sampled subsets $\mathcal{D}_{sub}$, with $|\mathcal{D}_{sub}|\ll|\mathcal{D}_{train}|$, where the diagonalized Hessian can be cheaply cached and inverted.
+Observing that finding the most influential training sample on $z$ needs to iterate overall $N$ training samples, (Guo et al., 2021) propose to use fast KNN to pre-filter a small subset of influence-worthy data points from $\mathcal{D}_{train}$ as candidates to explain small pretrained language models, whereas (Han and Tsvetkov, 2022) propose to find a small subset $\mathcal{D}_{sub}\subset\mathcal{D}_{train}$ whose gradient is the most similar to the downstream task examples.
+(Grosse et al., 2023) propose to use the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation to scale influence functions to 52B LLMs.
+For adaptation, only influences mediated by the feed-forward networks (FFNs) are considered.
+Based on the assumption that weights from different FFN layers are independent, the EK-FAC approximated influence is finally formulated as the sum of influences mediated by each layer as follows:
+(10) $\mathcal{I}_{\text{EKFAC}}(z_{i},z)=\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(z,\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta}),$
 where $\theta^{(l)}$ denotes the weights of the $l$-th MLP layer, and $\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}$ is the EK-FAC approximated Gauss-Newton Hessian for $\theta^{(l)}$.
-Since the inversion of $L$ small ${K_{l}\times K_{l}}$ matrices (i.e., $\mathcal{O}(L\times K_{l}^{3})$) is substantially more efficient than the inversion of a large $L{K_{l}\times LK_{l}}$ matrix (i.e., $\mathcal{O}{((LK_{l}})^{3})$), $\mathcal{I}_{\text{EKFAC}}$ can be adaptable to very large models, let alone the HVP can be further simplified by utilizing the EK-FAC properties.
-Recently, influence function-based explanation has been used in efficient finetuning of LLMs, where influence function has been used to select a small subset of training samples given few-shot validation samples for a specific downstream task, where the training overhead can be substantially improved (Xia et al., 2024).
-4.1.2 Embedding-based Methods
-Another strategy for sample-based explanation involves leveraging the hidden representations within the transformer architecture, which is recognized for encoding high-level semantics from textual data, to calculate the semantic similarity between $z$ and $z_{i}$.
+Since the inversion of $L$ small ${K_{l}\times K_{l}}$ matrices (i.e., $\mathcal{O}(L\times K_{l}^{3})$) is substantially more efficient than the inversion of a large $L{K_{l}\times LK_{l}}$ matrix (i.e., $\mathcal{O}{((LK_{l}})^{3})$), $\mathcal{I}_{\text{EKFAC}}$ can be adaptable to very large models.
+Recently, influence functions have also been used to explain and improve the in-context learning ability of LLMs.
+Based on the finding  (Von Oswald et al., 2023) that during in-context learning, LLMs can be viewed as implicitly ”learning” an internal kernelized least square surrogate objective on the few-shot examples included in the prompt, Zhou et al.
+(2024) propose to use this hypothesized objective as the loss function to calculate the influence of each in-context sample on the LLM generation, and based on which select and re-rank the samples.
+They show that the in-context learning ability of LLMs can be improved after the sample selection and demonstration order re-ranking.
+4.1.2. Embedding-based Methods
+Another strategy for sample-based explanation leverages hidden representations within the transformer architecture, which is recognized for encoding high-level semantics from textual data, to calculate the semantic similarity between $z$ and $z_{i}$.
 The similarity can also be used to measure the influence of $z_{i}$ on $z$ as explanations (Rajani et al., 2019).
-Specifically, Akyurek et al.
-(2022) propose to represent the training sample $z_{i}$ and test sample $z$ by concatenating the input and output as $z^{cat}_{i}=[x_{i}||y_{i}]$, $z^{cat}=[x||y]$.
+Specifically, (Akyurek et al., 2022) propose to represent the training sample $z_{i}$ and test sample $z$ by concatenating the input and output as $z^{cat}_{i}=[x_{i}||y_{i}]$, $z^{cat}=[x||y]$.
 The concatenation is feasible for generation tasks where the output $y$ lies in the same token sequence space as the input prompt $x$.
-The similarity between $z_{i}$ and $z$ can then be calculated as follows:
-$\mathcal{I}_{\text{emb}}\left(z_{i},z\right)=\frac{f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\cdot f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)}{\left\|f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\right\|\left\|f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)\right\|},$ (4)
-where $f^{(l)}_{\hat{\theta}}$ is the sub-network that outputs the $l$-th layer intermediate activation of the pretrained LLM $f_{\hat{\theta}}$.
+The similarity between $z_{i}$ and $z$ can then be calculated as:
+(11) $\mathcal{I}_{\text{emb}}\left(z_{i},z\right)=\frac{f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\cdot f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)}{\left\|f^{(l)}_{\hat{\theta}}(z^{cat}_{i})^{\top}\right\|\left\|f^{(l)}_{\hat{\theta}}\left(z^{cat}\right)\right\|},$
+where $f^{(l)}_{\hat{\theta}}$ is the sub-network that outputs the $l$-th layer intermediate activation of $f_{\hat{\theta}}$.
 The Eq.
@@
 (9), which assigns a score $\mathcal{I}$ for the explainee $z$ for each training sample $z_{i}$ in the dataset $\mathcal{D}_{train}$ as the explanation confidence of the sample $z_{i}$.
-Compared with the influence function methods introduced in the previous part, embedding-based methods are computationally efficient, as for each explainee $z$, the explanation score from a training sample $z_{i}$ requires only one forward pass of the transformer network.
-In addition, the calculation can be easily paralleled for different training samples.
-However, the disadvantage is also evident: These methods lack a theoretical foundation and may fail to identify important training samples that may not be semantically similar to the test sample.
-Consider the following toy example: Training samples $z_{i}$ = (“1+1=”, “2”) and $z_{j}$ = (“2+2=”, “4”) make the LLM gain the ability to conduct arithmetic calculation, which explains why prompting the model with $x$ = “100+100” gives the results $y$=“200”.
-However, the embeddings between the test sample $z$ and the two training samples $z_{i}$ and $z_{j}$ can be very different when calculated via Eq.
-(4(Akyurek et al., 2022).
-Therefore, embedding-based methods may not be able to faithfully find the training samples where the explanations require generalization ability beyond semantic similarity.
-4.2 Case Study: EK-FAC-based Influence Estimation
-In this part, we implement the EK-FAC-approximated influence function proposed in Grosse et al.
-(2023), and verify its scalability and effectiveness on LLMs with billions of parameters, including GPT2-1.5B (Radford et al., 2019), LLaMA2-7B (Touvron et al., 2023b), Mistral-7B (Jiang et al., 2023), and LLaMA2-13B.
-4.2.1 Experimental Design
-We use the SciFact dataset (Wadden et al., 2020) as the corpora, which contains the abstract of 5,183 papers from the domain of basic science and medicine.
-The explainee LLMs are obtained by finetuning the pretrained LLMs for 20,000 iterations, where AdamW (Loshchilov & Hutter, 2018) is used as the optimizer, and the learning rate and weight decay are set to 1e-5, 1e-3, respectively.
-Then, we use 500 samples from the corpora to estimate the (i) uncentered covariance matrices of the activations and pre-activation pseudo-gradients $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$, and (ii) the variances of the projected pseudo-gradient $\bm{\Lambda}^{(l)}$ for each selected dense layer $l$, and cache them on the hard disk (details see Eqs.
-(16) and (20) in Grosse et al.
-(2023)).
-We select the c_fc layer for GPT2-1.5B, and gate_proj layer for LLaMA2-7B, Mistral-7B, and LLaMA2-13B***All the implementation and layer names are based on the huggingface transformers, where the details can be found in https://huggingface.co/docs/transformers/en/index..
-For evaluation, we randomly select 200 samples from the corpora to construct the test set, which we name SciFact-Inf.
-Specifically, for the $j$-th selected sample $z_{j}=(x_{j},x_{j})$ (here $y_{j}=x_{j}$ as label equals the input in language modeling), we use the first three sentences in $x_{j}$, i.e., $\hat{x}_{j}$, to generate a completion $\hat{y}_{j}$ with the finetuned LLM (here, $\hat{y}_{j}$ does not equal the remaining sentences in $x_{j}$), and we aim to explain the generation of $\hat{y}_{i}$ from $\hat{x}_{j}$ with the finetuned LLM with the training samples via EK-FAC approximated influence scores defined in Eq.
-(3).
+4.2. Case Study: EK-FAC-based Influence Estimation
+This case study implements the EK-FAC-based influence function (Grosse et al., 2023) and verifies its scalability and effectiveness on LLMs with billions of parameters, including GPT2-1.5B (Radford et al., 2019), LLaMA2-7B (Touvron et al., 2023b), Mistral-7B (Jiang et al., 2023), and LLaMA2-13B.
+4.2.1. Experimental Design
+We use the SciFact dataset (Wadden et al., 2020) as the corpora, which contains the abstract of 5,183 papers from basic science and medicine.
+The LLMs are obtained by finetuning the pretrained LLMs for 20,000 iterations, where AdamW is used as the optimizer (Loshchilov and Hutter, 2018).
+Then, we use 500 samples from the corpora to estimate the (i) uncentered covariance matrices of the activations and pre-activation pseudo-gradients $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$, and (ii) the variances of the projected pseudo-gradient $\boldsymbol{\Lambda}^{(l)}$ for each selected dense layer $l$, and cache them on the hard disk (details see Eqs.
+(16) and (20) in (Grosse et al., 2023)).
+We select the c_fc layer for GPT2-1.5B, and gate_proj layer for LLaMA2-7B, Mistral-7B, and LLaMA2-13B.
+For evaluation, we randomly select 200 samples to construct the test set, which we name SciFact-Inf.
+Specifically, for the $j$-th sample $z_{j}=x_{j}$, we use the first three sentences in $x_{j}$, i.e., $\hat{x}_{j}$, to generate a completion $\hat{y}_{j}$ with the finetuned LLM.
 Ideally, the $j$-th training sample $z_{j}$ itself should be the most influential sample w.r.t.
-the generation of $\hat{y}_{j}$ for test sample $\hat{z}_{j}$, which facilitates quantitative analysis of the effectiveness of Eq.
-(3).
-In our implementation, for each test sample $\hat{z}_{j}$, we first calculate the EK-FAC approximated HVP part of the influence $\mathcal{I}_{\text{EKFAC}}(z_{i},\hat{z}_{j})$, i.e., $\sum_{l}\nabla_{\theta^{(l)}}\mathcal{L}(\hat{z}_{j},\hat{\theta})^{\top}(\hat{\mathbf{G}}_{\hat{\theta}^{(l)}}+\lambda^{(l)}\mathbf{I})^{-1}$, which is shared for all training samples $z_{i}$.
-Specifically, we record the layer-wise gradient $\nabla_{\theta^{(l)}}\mathcal{L}(\hat{z}_{j},\hat{\theta})$ and calculate the HVP with the cached $\mathbf{Q}_{A}^{(l)}$, $\mathbf{Q}_{S}^{(l)}$ as Eq.
-(21) in Grosse et al.
-(2023).
-We then go through candidate training samples (1 positive and 99 negative), calculate the gradient $\nabla_{\theta^{(l)}}\mathcal{L}(z_{i},\hat{\theta})$, and take inner-product with the approximate HVP as the layer-wise influence.
-Finally, the layer-wise influences are summed up as Eq.
-(3) as the total influence $\mathcal{I}_{\text{EKFAC}}(z_{i},\hat{z}_{j})$.
-We rank the influence and calculate the top-$K$ hit rate of the positive training sample.
-4.2.2 Results and Analysis
-Table 4: Effectiveness of EK-FAC approximated influence function on the established SciFact-Inf dataset.
-Time (Pre.) stands for the time for precomputing the $\mathbf{Q}_{A}$, $\mathbf{Q}_{S}$, and $\bm{\Lambda}$.
-Time (Inf.) stands for the time for calculating the influence of 100 training samples per test sample.
-Experiments for GPT2-1.5B, LLaMA2-7B, Mistral-7B are done on 4 A100 GPUs, whereas experiments for LLaMA2-13B are done with 8 A100 GPUs.
+the generation of $\hat{y}_{j}$.
+4.2.2. Results and Analysis
+Table 2.
+Effectiveness of EK-FAC approximated influence function on the SciFact-Inf dataset.
+Time (Pre.) stands for the time for precomputing the $\mathbf{Q}_{A}$, $\mathbf{Q}_{S}$, and $\boldsymbol{\Lambda}$.
+Time (Inf.) stands for time for calculating the influence of 100 training samples per test sample.
 Strategy LLM Recall@5 Recall@10 Time (Pre.) Time (Inf.)
-Random - 0.0100 $0$
 Inf. Func. GPT2-1.5B 0.6368 0.7363 0h 27min 0min 28sec
@@
 LLaMA2-13B 0.7811 0.8940 3h 11min 3min 08sec
-The experimental results are summarized in Table 4.
-From Table 4 we can find that, the EK-FAC approximated influence function achieves a good accuracy in finding the training sample that has the greatest influence on the generation of a test sample, even if only the influences mediated by a small part of dense layers are considered.
-In addition, we find that the main computational bottleneck in calculating the EK-FAC-based influence is to estimate the covariances $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$ and variance $\bm{\Lambda}^{(l)}$, which can take hours when 500 training samples are used for the estimation.
-However, after the estimation, it is relatively cheap to calculate the influence of 500 training samples for each test sample, which takes around 3 minutes for a 13B LLaMA2 model on 8 A100 GPUs.
-This demonstrates the scalability of EK-FAC-based influence by assuming the independence of different dense layers and using EK-FAC to simplify the computation.
-4.3 Challenges
-Overall, explaining the generation of LLMs by tracing back to the training samples is still an emerging area.
-Open questions need to be addressed to further advance the field.
-In this section, we identify three main challenges as follows, which can serve as directions for future explorations.
-4.3.1 Strong Assumptions for Scalability
+The results are summarized in Table 2.
+We observe that the EK-FAC approximated influence function achieves good accuracy in finding the training sample with greatest influence on the generation of LLMs.
+In addition, we find that the main computational bottleneck in calculating the EK-FAC-based influence is to estimate the covariances $\mathbf{Q}^{(l)}_{A}$, $\mathbf{Q}^{(l)}_{S}$ and variance $\boldsymbol{\Lambda}^{(l)}$, which takes hours when 500 training samples are used for the estimation.
+However, after that, it is relatively cheap to calculate the influence training samples for each test sample.
+4.3. Challenges
+Overall, explaining the generation of LLMs by tracing back to the training samples is still an emerging area. Open questions need to be addressed to further advance the field.
+4.3.1. Strong Assumptions for Scalability
 The unprecedented number of parameters in modern LLMs causes severe scalability issues for sample-based explanation strategies.
@@
 (8) via first-order approximation.
-Schioppa et al.
-(2022) assume the Hessian to be low rank.
-Grosse et al.
-(2023) that assume that the weights from different layers of the LLMs are independent, as well as the tokens in different steps, such that EK-FAC can be appropriately applied to approximate the influence function.
-From the above analysis, we can find that while the method from Grosse et al.
-(2023) has the best scalability, it also has the strongest assumption, which may fail to hold in practice.
-While highly efficient to compute, embedding-based methods make the implicit assumption that semantics similarly implies explainability, which we have demonstrated may not always be the case.
+(Schioppa et al., 2022) assume the Hessian to be low rank.
+(Grosse et al., 2023) that assume that the weights from different layers of the LLMs are independent, as well as the tokens in different steps, such that EK-FAC can be appropriately applied to approximate the influence function.
+From the above analysis, we can find that while the method from (Grosse et al., 2023) has the best scalability, it also has the strongest assumption, which may fail to hold in practice.
 Therefore, how to improve the scalability with weak assumptions needs to be investigated in the future.
-4.3.2 Explainability v.s. Understandability
-Despite the advantage of influence/embedding similarity to provide a qualitative measurement of a specific training sample as the explanation for LLM generation, the understandability of the identified sample can still be weak, where the connection between the selected training samples and the generation may not be understandable to human beings.
-Specifically, Grosse et al.
-(2023) cautions that the sign of influence score of the training tokens may be difficult for humans to connect to the positive or negative influence on the generation results.
-This severely jeopardizes the usability of the identified training samples.
-In addition, Grosse et al.
-(2023) also found that, since LLMs are usually not trained to the minimum to avoid overfitting (and due to overparameterization, the number of local minimums may be large), the connection between influence defined in Eq.
-(1) with the counterfactual loss of removing the sample $z_{i}$ at $z$ is also weak.
-For the embedding-based methods, since most LLM models are black box transformer models, the similarity of embeddings can also be hard to interpret by human beings; therefore, it is imperative to improve the interpretability of the identified training samples, such that tracing back becomes more meaningful.
-4.3.3 LLM-Oriented Sample-based Explanations
+4.3.2. LLM-Oriented Sample-based Explanations
 Finally, we observed that both gradient-based and embedding-based methods are loosely connected to the LLM, as well as the backbone transformer networks.
 For example, algorithms like TracIn (Pruthi et al., 2020) are designed to scale up influence functions to large models, which are not specific for LLMs.
-Similarly, the embedding-based method proposed in Akyurek et al.
-(2022) is applicable to most machine learning models with latent representations.
-Grosse et al.
-(2023) considers the specialty of LLMs by utilizing the knowledge neuron assumption of the backbone transformers (Wang et al., 2023d) to simplify the influence function, where the weights considered are constrained to the MLP layers, which may not fully utilize the property of transformers.
+Similarly, the embedding-based method proposed in (Akyurek et al., 2022) is applicable to most machine learning models with latent representations.
+(Grosse et al., 2023) considers the specialty of LLMs by utilizing the knowledge neuron assumption of the backbone transformers (Wang et al., 2023d) to simplify the influence function, where the weights considered are constrained to the MLP layers, which may not fully utilize the property of transformers.
 Therefore, how to further utilize the property of the LLM and the backbone transformer to design LLM-tailored sample-based influence/similarity (either to reduce the computational/space overhead or to improve the explanation quality) is highly promising for future work.
-5 Explainability for Trustworthy LLMs and Human Alignment
+5. Explainability for Trustworthy LLMs and Human Alignment
 In previous sections, we explore the use of explanation techniques for assessing and improving the performance of LLMs.
 In this section, we shift the focus towards examining LLM trustworthiness.
-As LLMs are increasingly integrated into various applications of daily life, including high-stakes areas like healthcare, finance, and legal advice, it is crucial that their responses not only are accurate but also align with human ethical standards and safety protocols (Liu et al., 2023b; Li et al., 2023f).
-Thus, the need arises to extend the scope of explanations from merely evaluating the accuracy of LLMs to scrutinizing their trustworthiness.
-Herein, we delve into how explanation techniques, discussed in the previous sections, can be instrumental in assessing LLMs across key aspects of trustworthiness like security, privacy, fairness, toxicity, and honesty.
+As LLMs are increasingly integrated into high-stakes areas like healthcare, finance, and legal advice, it is crucial that their responses not only are accurate but also align with human ethical standards and safety protocols (Liu et al., 2023c; Li et al., 2023a; Hu et al., 2023).
+Herein, we examine how explanation techniques, discussed in the previous sections, can be instrumental in assessing LLMs across key aspects of trustworthiness like security, privacy, fairness, toxicity, and honesty.
 It is worth noting that while explainability itself is an aspect of trustworthiness, it holds the promise of serving as a foundational tool for addressing other trustworthiness concerns.
-5.1 Security
-LLMs are known to be vulnerable to attacks and exploitation, such as spreading misinformation, launching phishing attacks, and poisoning training data (Derner et al., 2023).
-For enhanced safety, LLMs are designed to reject certain types of prompts that may result in the generation of harmful content, e.g., by precluding prompts likely to elicit unsafe outputs from the training phase.
+5.1. Security
+LLMs are known to be vulnerable to attacks and exploitation, such as spreading misinformation, and poisoning training data (Derner et al., 2023).
+For enhanced safety, LLMs are designed to reject prompts that may result in harmful contents.
 However, jailbreak techniques can circumvent these restriction measures and manipulate LLMs into producing malicious contents.
-Malevolent users (i.e., attackers) can craft special prompts that compel or induce LLMs to prioritize instruction following over rejections (Liu et al., 2023c; Li et al., 2023a).
+Malevolent users can craft special prompts that compel LLMs to prioritize instruction following over rejections (Liu et al., 2023a; Li et al., 2023c).
 For example, through Prefix Injection, attackers can use out-of-distribution prompt prefixes that are less likely to be rejected (walkerspider, 2022; Wei et al., 2023).
-Another approach, called Refuse Suppression, involves directing or persuading models to ignore established safety protocols (Wei et al., 2023; Zeng et al., 2024), where the instruction following ability is then employed to perform the attack.
-Existing methods mainly rely on prompt engineering for attacking LLMs, but they usually have low attack success rates and significant time costs (Li et al., 2024c).
-Thus, by understanding and engineering latent representations of LLMs, explanation methods provide a viable way to design advanced attacks and discover the potential vulnerabilities of LLMs (Liu et al., 2021).
-For example, a recent work extracts “safety patterns” via explaining the latent space of LLMs using representation engineering.
-Specifically, these patterns can be captured from the activation differences between malicious queries and benign queries.
-The salient portion of difference vectors’ dimensions is localized and utilized to generate features of safety patterns.
-The safety patterns reflect the internal protection mechanisms within LLMs.
-Circumventing these patterns leads to novel attacks, which helps exploring potential vulnerabilities of LLMs (Li et al., 2024c).
+Another approach, called Refuse Suppression, involves directing or persuading models to ignore established safety protocols (Wei et al., 2023; Zeng et al., 2024a), where the instruction following ability is then employed to perform the attack.
+{forest}
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, rectangle, draw, font=, rounded corners, align=center, minimum width=2em, edge+=darkgray, line width=1pt, s sep=4pt, inner xsep=3pt, inner ysep=2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, text centered, , [Trustworthy LLMs & Human Alignment,ver,fill=gray!20,text width=10em,text centered [Security,ver,fill=orange!40,text centered,text width=2.5em [Adversarial Attack,fill=orange!10,text centered,text width=6.5em [(Liu et al., 2023a), (Li et al., 2023c), (Wei et al., 2023), (Zeng et al., 2024a), (walkerspider, 2022), (Liu et al., 2021), (Jain et al., 2023)., leaf1, align=left] ] [Adversarial Defense,fill=orange!10,text centered,text width=6.5em [(Wu et al., 2025c), (Yu et al., 2024b), (Zheng et al., [n.
+d.]), (Yousefpour et al., 2025)., leaf1, align=left] ] ] [Privacy,ver,fill=orange!40,text width=2.5em,text centered [Memorization Attack,fill=orange!10,text centered,text width=6.5em [(Nasr et al., 2023), (Li et al., 2023c), (Fu et al., 2024), (Bai et al., 2024)., leaf1, align=left] ] [Knowledge-based Defense,fill=orange!10,text centered,text width=6.5em [(Meng et al., 2022a), (Dai et al., 2022), (Hase et al., 2024), (Yin et al., 2024b), (Suri et al., 2025)., leaf1, align=left] ] ] [Fairness,ver,fill=orange!40,text width=2.5em,text centered [Bias Detection,fill=orange!10,text centered,text width=6.5em [(Adebayo et al., 2023), (Ma et al., 2023), (Zou et al., 2023a), (Qian et al., 2024)., leaf1, align=left] ] [Bias Mitigation,fill=orange!10,text centered,text width=6.5em [(Rakshit et al., 2024), (Ma et al., 2023), (Yu et al., 2023), (Thakur et al., 2023)., leaf1, align=left] ] ] [Toxicity,ver,fill=orange!40,text width=2.5em,text centered [Toxic Content Detection,fill=orange!10,text centered,text width=6.5em [(Balestriero et al., 2023), (Wang et al., 2024d), (Lee et al., 2024), (Li et al., 2024d)., leaf1, align=left] ] [Toxic Content Mitigation,fill=orange!10,text centered,text width=6.5em [(Lee et al., 2024), (Liu et al., 2024b), (Leong et al., 2023)., leaf1, align=left] ] ] [Truthfulness,ver,fill=orange!40,text width=2.5em,text centered [Hallucination Detection,fill=orange!10,text centered,text width=6.5em [(Azaria and Mitchell, 2023), (Campbell et al., 2023), (Duan et al., 2024), (Li et al., 2024b), (Marks and Tegmark, 2023), (Xu et al., 2023a), (Heo et al., 2024), (Park et al., 2025)., leaf1, align=left] ] [Hallucination Mitigation,fill=orange!10,text centered,text width=6.5em [(Li et al., 2024b), (Chen et al., [n.
+d.]), (Liu et al., 2024b), (Yu et al., 2024a), (Huang et al., 2024), (Zhang et al., 2024c), (He et al., 2024)., leaf1, align=left] ] ] ]
+Figure 4. Taxonomy of explanation techniques for trustworthiness aspects of LLMs.
+By engineering latent representations of LLMs, explanation methods provide a viable way to discover the potential vulnerabilities of LLMs by simulating advanced attacks (Liu et al., 2021).
+For example, a recent work extracts “safety patterns” via explaining the latent space.
+These patterns are captured from the activation differences between malicious and benign queries.
+They reflect the internal protection mechanisms within LLMs.
+Circumventing these patterns leads to novel attacks, which helps exploring potential vulnerabilities of LLMs (Li et al., 2024d).
 Besides, a deeper understanding of fine-tuning can shed light on the reliability of existing safety measures.
 In particular, Jain et al.
-(2023) use networking pruning, attention map activation, and probing classifiers to track the changes of model capabilities from pre-training to fine-tuning.
-These tools are helpful in finding salient weights that identify crucial neurons to restore pre-training capabilities.
-These neurons have demonstrated that the capabilities gained during fine-tuning can be removed easily through fine-tuning on other unrelated tasks.
+(2023) apply network pruning, attention maps, and probing classifiers to track the changes of model capabilities from pre-training to fine-tuning.
+These tools are helpful in demonstrating that the capabilities gained during fine-tuning can be easily removed through fine-tuning on other unrelated tasks.
 This finding casts doubt on the robustness of current safety alignments in LLMs.
-5.2 Privacy
-Recent studies have revealed that LLMs such as ChatGPT can leak extensive amounts of training data through a method known as divergence attacks.
-These attacks utilize specially crafted prompts to lead the model away from its standard chatbot-style generation (Nasr et al., 2023).
+Finally, by probing LLM representations, Zheng et al.
+([n.
+d.]) investigates the impact of safety prompts that motivates a novel safety prompt optimization method to safeguard LLMs.
+5.2. Privacy
+Recent studies have revealed that LLMs such as ChatGPT can leak extensive amounts of training data through divergence attacks (Nasr et al., 2023).
+These attacks utilize specially crafted prompts to lead the model away from its standard chatbot-style generation.
 The risk of private data exposure through such means poses a serious challenge to the development of ethically responsible models.
-This issue is compounded by strategies similar to jailbreak attacks, where misalignment is exploited to induce LLMs into operating in an unconventional “developer mode” via out-of-distribution prompts (Li et al., 2023a).
-Traditional data pre-processing techniques, such as data sanitization (Lison et al., 2021), are impractical as defenses for LLMs due to the enormous scale of training data.
-Enhancing LLM privacy involves two strategic approaches: (1) preventing the models from memorizing sensitive data, and (2) establishing safeguards against the release of sensitive information during content generation.
-The latter can employ techniques used in jailbreak defenses, treating prompts that solicit private information as potentially malicious.
-The former approach requires identifying whether LLMs possess specific knowledge, which is traditionally tested by crafting question-answering (QA) tasks to determine if LLMs can provide answers.
-However, this approach faces limitations due to LLMs’ sensitivity to the phrasing of QA prompts, while the optimal prompt is usually unknown.
-To tackle the challenge, explanatory techniques can serve as a tool to confirm whether LLMs have internalized certain knowledge.
-For instance, via explaining the relation between factual knowledge and neuron activations (Meng et al., 2022a; Dai et al., 2022; Hase et al., 2024), we may investigate whether and where a piece of factual knowledge is stored within transformers.
-In addition, Yin et al.
-(2024a) recently proposes the concept of “knowledge boundary” and develops a gradient-based method to explore whether LLMs master certain knowledge independent of the input prompt.
-5.3 Fairness
-Despite LLMs’ powerful generation capabilities, their widespread applications also bring concerns about exacerbating bias issues in society, as LLMs are able to learn social biases within human-generated corpus (Gallegos et al., 2023).
-For example, in a gender bias case, “[He] is a doctor” is much more likely than “[She] is a doctor” because of the gender stereotype.
-In this subsection, we focus on fairness issues that refer to biases related to race, gender, and age within human communities (Li et al., 2023g).
-There is a rich body of literature quantifying fairness issues within LLMs through various tests (Adebayo et al., 2023).
-Interpretation complements these methods by providing a unique perspective for bias mitigation, focusing on unraveling the mechanisms through which biases are embedded into LLMs.
-A research direction within this domain is the examination of biased attention heads.
-For instance, Ma et al.
-(2023) detect stereotype encodings through probing attention heads and evaluating their attributions via Shapley values (Lundberg & Lee, 2017).
-The results unveil that approximately 15% to 30% of attention heads across six transformer-based models are linked to stereotypes.
-These attention heads tend to specialize in maintaining various stereotypes, offering a potential pathway for developing effective de-biasing techniques.
-Moreover, scoring head biases with gradient-based metrics offers another way for identifying biased heads (Yang et al., 2023a).
+This issue is compounded by strategies similar to jailbreak attacks, where misalignment is exploited to induce LLMs into operating in an unconventional “developer mode” via out-of-distribution prompts (Li et al., 2023c).
+Enhancing LLM privacy involves two strategic approaches: (1) preventing LLMs from memorizing sensitive data, and (2) establishing safeguards against the release of sensitive information.
+The second strategy involves employing techniques from jailbreak defenses, treating prompts that solicit private information as potentially malicious.
+The first approach requires identifying specific knowledge within models, which is traditionally achieved with prompt engineering.
+However, prompt engineering faces limitations due to LLMs’ sensitivity to the phrasing of prompts.
+Explanatory techniques can serve as an auxiliary tool to confirm whether LLMs have internalized certain knowledge.
+For instance, factual knowledge is localized via explaining the relation between factual knowledge and neuron activations (Meng et al., 2022a; Dai et al., 2022; Hase et al., 2024).
+In addition, (Yin et al., 2024b) recently proposes the concept of “knowledge boundary” and develops a gradient-based method to explore whether LLMs master certain knowledge independent of the input prompt.
+5.3. Fairness
+The widespread applications of LLMs also bring concerns about exacerbating bias issues in society (Gallegos et al., 2023).
+For example, in a gender stereotype case, “[He] is a doctor” is much more likely than “[She] is a doctor”.
+In this subsection, we focus on fairness issues related to race, gender, and age (Li et al., 2023b).
+Prompt engineering is a major method to quantify fairness issues within LLMs (Adebayo et al., 2023).
+Interpretation complements this method by unraveling the mechanisms through which biases are embedded into LLMs.
+One direction is examining biased attention heads.
+Ma et al.
+(2023) unveil that approximately 15% to 30% of attention heads linking to specific stereotypes through probing attention heads.
 Furthermore, recent work has placed LLM representations under scrutiny (Zou et al., 2023a).
-Typically, specific templates relevant to a given concept or function are designed beforehand.
-Then, representations closely aligned with the concepts or functions are examined using principal component analysis (PCA).
-From this analysis, a vector is derived from the first principal component to predict a certain bias.
-To achieve fair model predictions, a diverse range of mitigation techniques have been proposed to debias models.
-One stream of work proposes to debias LLMs at the embedding level, refining embeddings through various approaches.
-For example, a recent work attempts to alter biased embeddings with minimal alterations to make them orthogonal to neutral embeddings (Rakshit et al., 2024).
-Additionally, some studies concentrate on removing biases at the level of attention heads.
+Typically, representations of crafted templates showing specific biases are examined to derive a vector that identifies a certain bias.
+To achieve fair model outputs, a diverse range of mitigation techniques have been proposed.
+One stream of work proposes to debias LLMs at the embedding level.
+For example, one work alters biased embeddings with minimal alterations to make them orthogonal to neutral embeddings (Rakshit et al., 2024).
+Additionally, some studies remove biases at the level of attention heads.
 Ma et al.
-(2023) address this by pruning attention heads that significantly contribute to certain biases.
-Similarly, another study adapts movement pruning (Sanh et al., 2020) to dynamically choose an optimal subset of low-bias attention heads (Joniak & Aizawa, 2022).
-Beyond modifying embeddings and pruning attention heads, another strategy involves targeting a specific group of neurons known to propagate biases.
+(2023) address this by pruning biased attention heads.
+Beyond modifying embeddings and pruning attention heads, another strategy targets a specific group of neurons known to propagate biases.
 It unlearns the biases by retraining weight vectors for these neurons (Yu et al., 2023).
-Besides, bias mitigation can also be approached from a data-centric perspective using a few training samples (Thakur et al., 2023).
-This work uses a pre-trained model to find the most biased training examples, and then modifies these examples to fine-tune the model.
-5.4 Toxicity
+Besides, bias mitigation can also be approached from a data-centric perspective using the most biased training samples, then modifying them to fine-tune the model (Thakur et al., 2023).
+5.4. Toxicity
 Toxicity is another form of harmful content that LLMs may produce.
-This issue arises because LLMs are often trained on extensive online text corpora that have not been thoroughly filtered, containing elements of toxicity that can hardly be fully eliminated.
+This issue arises from training documents containing elements of toxicity that can hardly be fully eliminated.
 Toxicity can be identified by interpreting LLM components like the feed-forward layers and attention heads.
-For instance, recent work reveals how toxicity is represented within LLMs by identifying multiple vectors promoting toxicity within the MLP layers and further recognizes relevant dimensions through singular value decomposition (Lee et al., 2024).
+For instance, recent work reveals how toxicity is represented within LLMs by identifying multiple vectors and relevant dimensions promoting toxicity within the MLP layers (Lee et al., 2024).
 Furthermore, the exploration of geometric structures in per-layer representations offers another way to detect toxicity.
-Balestriero et al.
-(2023) applied spline formulations within MLPs to extract seven characteristic input features, demonstrating their utility in describing the domain of prompts and classifying toxic remarks.
-The insights into how toxicity manifests within LLMs also shed light on mitigation strategies.
+(Balestriero et al., 2023) extract input features from MLPs to describe the domain of prompts and classifying toxic remarks.
+Aforementioned insights also shed light on mitigation strategies.
 Motivated by the finding that toxicity can be reduced by manipulating relevant vectors, Lee et al.
-(2024) develops a method called direct preference optimization (DPO), where paired toxic and non-toxic samples are utilized to fine-tune models so that non-toxic content is promoted.
+(2024) utilizes paired toxic and non-toxic samples to fine-tune models so that non-toxic content is promoted.
 By examining the changes in the parameter matrices during the fine-tuning process, it substantiates that even minor adjustments to these critical vectors can reduce toxicity.
 Built on the observation that LLMs’ representations are updated by outputs from attention layers (Elhage et al., 2021), another work attempts to reduce toxicity by identifying the “toxicity direction” and then adjusting representations in the opposite direction (Leong et al., 2023).
-5.5 Truthfulness
-One prominent drawback of LLMs is their tendency to confidently produce false statements.
-These statements fall into two main categories: 1) statements that contradict learned knowledge within models, a problem often related to models’ honesty; 2) statements that are factually incorrect and appear to be fabricated by models, a phenomenon commonly referred to as hallucination.
-In the following, we delve into various approaches that aim to understand aforementioned two behaviors by leveraging explainability tools.
-5.5.1 Honesty
-Honesty of LLMs describes models’ ability to produce true statements based on their learned information, where dishonest behaviors significantly compromise the perceived trustworthiness of LLMs.
-Numerous studies have been undertaken to understand how and why dishonest behaviors happen by looking inside LLMs.
-One notable work attempts to distinguish dishonesty by training a classifier to predict the accuracy of statements (Azaria & Mitchell, 2023).
-The classifier is simply trained on top of activations from the hidden layers of LLMs.
-These activations are generated from true and false statements.
-The classifier reaches an accuracy range between 60% and 80%, which suggests that LLMs may be internally aware of the truthfulness of their outputs (Azaria & Mitchell, 2023).
-Furthermore, research by Campbell et al.
-(2023) localizes dishonesty behaviors at the level of attention heads.
-This study employs specially crafted prompts to induce dishonest responses, and then trains logistic classifiers on models’ activations over true/false statements.
-It also employs activation patching to substitute lying activations with honesty ones.
-Intriguingly, both approaches have witnessed the importance of layer 23--29 in flipping dishonesty behaviors.
-Besides, another popular method tries to study the geometric structure of true/false statements(Marks & Tegmark, 2023).
-Typically, these structures are visualized by projecting representations of statements onto two principal components.
-A clear linear structure and the truth directions can be derived to mitigate the dishonest behaviors.
-5.5.2 Hallucinations
-Hallucinations in LLMs can arise due to poor data quality, biases, outdated knowledge, and the lack of explicit knowledge (Xu et al., 2024; Zhu et al., 2023b).
-However, whether LLMs are aware of their hallucination behaviors remains an open question.
+5.5. Truthfulness
+LLMs tend to confidently produce false statements that fall into two main categories: 1) statements that contradict learned knowledge, also known as dishonesty; 2) statements that are factually incorrect and fabricated by models, that is hallucination.
+In the following, we look into explainability tools that aim to understand the above two behaviors.
+5.5.1. Dishonesty
+Dishonesty of LLMs describes models’ ability to produce false statements compared to their learned information.
+Studies have been undertaken to understand how and why it happens.
+One notable work distinguishes dishonesty by training a classifier to predict the accuracy of statements, on top of activations from true/false statements (Azaria and Mitchell, 2023).
+The classifier reaches an accuracy range between 60% and 80% (Azaria and Mitchell, 2023).
+Furthermore,  Campbell et al.
+(2023) localize dishonesty behaviors at the level of attention heads.
+Despite adopting above probing-based method, it also employs activation patching to substitute lying activations with honesty ones.
+Both approaches have witnessed the importance of layer 23--29 in flipping dishonesty behaviors.
+Besides, another popular method tries to study the geometric structure of true/false statements(Marks and Tegmark, 2023).
+A linear structure and the truth directions are derived to mitigate the dishonest behaviors.
+5.5.2. Hallucinations
+Hallucinations in LLMs can arise due to poor data quality and the lack of explicit knowledge (Xu et al., 2024).
+However, whether LLMs are aware of their hallucination behaviors remains unknown.
 Recent work investigates this question by examining models’ hidden representation space (Duan et al., 2024).
-It examines three hidden states involving a question, its correct answer and its incorrect answer, which are used to compute an “awareness” score.
-This metric quantifies the uncertainty of LLMs regarding their own answers, finding that adversarially induced hallucination can increase models’ awareness.
+It defines an “awareness” score to quantify the uncertainty of LLMs’ answers, finding that adversarially induced hallucination can increase models’ awareness.
 Additionally, Li et al.
-(2024b) illustrates the major differences between models’ output and their inner activations, identifying these discrepancies as a potential source of hallucination.
-By training linear probing classifiers on each attention head’s activations, the most specialized attention head is identified.
-An orthogonal probe is subsequently trained with the identified specialized probe.
-The experiments reveal that “truth” might exist in a subspace instead of a single direction (Li et al., 2024b).
-Another work investigates the source of hallucination by analyzing patterns of source token contributions through perturbations (Xu et al., 2023c).
-Their findings suggest that hallucinations may stem from the models’ excessive dependence on a restricted set of source tokens.
-Besides, the static distribution of source token contribution, termed as “source contribution staticity”, can be used as another indicator of hallucinations.
-Building on the above insights into LLM hallucinations, Duan et al.
-(2024) apply PCA to derive the direction of the correct answer’s final hidden state, and enhance the hidden representations with this direction to reduce hallucinations.
-In contrast, Li et al.
-(2024b) adopts a different approach, by intervening on top-$K$ specialized attention heads, while minimizing the influence of the rest attention heads within models.
-Different from PCA that identifies a single principle direction, this work adopts two distinct techniques to find multiple directions of intervention.
-First, they use orthogonal vectors of each probe’s hyperplane, which is similar to PCA.
-Second, they leverage vectors that connect the mean of the true and false distributions (Li et al., 2024b).
-The vectors derived from mean shift has been demonstrated more effective than those from probe classifiers, which presents another feasible strategy for identifying directions of truth.
-5.6 Challenges
-We discuss the challenges in employing explanations to improve models’ trustworthiness and enhance alignment from two aspects: 1) limitations of existing detection techniques, and 2) shortcomings of mitigation strategies based on explanations.
-5.6.1 Challenges of Existing Detection Methods
-Current detection methods primarily focus on the level of LLM modules, such as layers, attention heads, and representations.
-However, we still lack a finer-grained understanding of how knowledge is encoded within LLMs, such as at the level of individual neurons and weight parameters.
-Furthermore, we lack general and robust strategies to identify these knowledge reliably.
-For instance, to identify gender biases, attention heads might be examined and then the related heads pruned (Li et al., 2024b).
-This approach requires analyzing each model individually, rather than adopting a general approach.
-Moreover, existing localization approaches rely either on probing classifiers or casual scrubbing, which might not be reliable.
-For probing classifiers, the pre-designed biases used to train these classifiers are crucial to their performance.
-On the other hand, casual cleaning usually introduces new variables that complicate the analysis.
-5.6.2 Challenges of Mitigation Strategies
-Since LLMs are trained on vast corpora of data, it is impractical to mitigate the aforementioned trustworthiness issues from the data perspective.
-It is widely believed that XAI, i.e.
-understanding the inner mechanisms of LLMs, is a fundamental way to solve these issues.
-Thus, mitigation methods for LLMs are typically developed based on explanations.
+(2024b) identifies differences between models’ output and their inner activations as a potential source of hallucination.
+By training two orthogonal probes to mitigate hallucinations, they reveal that “truth” might exist in a subspace instead of a single direction (Li et al., 2024b).
+Another work investigates the source of hallucination by analyzing patterns of source token contributions through perturbations (Xu et al., 2023a).
+They find that hallucinations may stem from the model’s excessive dependence on a restricted set of source tokens, and the static distribution of source token contributions.
+5.6. Challenges
+We discuss challenges in employing explanations to improve models’ trustworthiness and enhance alignment: 1) limitations of existing evaluation techniques, and 2) shortcomings of mitigation strategies based on explanations.
+5.6.1. Challenges of Existing Evaluation Methods
+Current detection methods primarily focus on layers, attention heads, and representations.
+However, we still lack robust understanding of how a certain trustworthiness concept (e.g., fairness, harmlessness) is comprehended by LLMs.
+Furthermore, we lack general strategies to identify this knowledge.
+Existing methods often rely on building an evaluation dataset, but it is difficult to guarantee the its comprehensiveness.
+Meanwhile, the definition of what constitutes a trustworthiness concept often varies among different stakeholders and contexts, complicating the evaluation process.
+Moreover, existing localization approaches rely on probing classifiers or casual scrubbing, which might not be reliable, since the performance of probing classifiers depends on the comprehensiveness of pre-defined biases, while casual scrubbing usually introduces new variables that complicate the analysis.
+5.6.2. Challenges of Mitigation Strategies
+Currently, mitigation methods for LLMs are typically developed based on explanations.
 Existing explanations are implemented using techniques from mechanistic interpretability and representation engineering (Zhao et al., 2024).
-While both streams of methods can alleviate these issues, they fail to fully address them.
-For example, principal component analysis (PCA) is often used to find geometric structures related to these issues, but cannot fix these issues.
-Another popular mitigation method is casual scrubbing, which introduces certain "positive" activations to replace those demonstrated to be responsible for specific issues.
-However, the identified directions and patched activations can only mitigate issues to a certain extent.
+However, neither stream of methods can fully address the issues.
+For example, the identified directions from representation engineering and patched activations from mechanistic interpretability can only mitigate issues to a certain extent.
 Moreover, the changes to either representations or activations could also influence other aspects of models’ capabilities, which we are yet unable to evaluate.
-6 LLM Enhancement via Explainable Prompting
-A key distinction between LLMs and traditional machine learning models lies in the LLMs’ ability to accept flexibly manipulated input data, namely prompts, during model inference (Liu et al., 2023a).
-LLMs generally give precedence to the information presented in these prompts when generating outputs.
-Therefore, to mitigate the opacity issue in LLM predictions, we can enhance prompts with understandable content, which is then prioritized over the LLMs’ inherent and implicit knowledge.
-These enriched prompts can include domain-specific insights, contextual information, or a step-by-step reasoning chain.
-In response, LLMs might reveal their decision-making processes during inference, which improves the explainability of their behaviors.
-6.1 Chain of Thoughts (CoT) Prompting
-The Chain of Thought (CoT) approach significantly enhances LLMs in tackling complex tasks (Wei et al., 2022).
-While LLMs are adept at generating human-like responses, they often lack transparency in their reasoning processes.
-This limitation makes it difficult for users to assess the credibility of the responses, especially for questions that require detailed reasoning.
-To bridge this gap, recent efforts have incorporated in-context learning with human-crafted explanations directly into prompts (Wei et al., 2022; Huang et al., 2023a; Yao et al., 2023b; Besta et al., 2023).
-Among these approaches, the CoT method stands out by employing explicit knowledge to guide the reasoning process.
+6. Effective LLM Inference via Explainable Prompting
+A key distinction between LLMs and traditional machine learning models lies in the LLMs’ ability to accept flexibly manipulated input data, namely prompts, during model inference (Liu et al., 2023d).
+To mitigate the opacity issue in LLM predictions, we can enhance prompts with explainable content.
+In response, LLMs can reveal their decision-making processes during inference, which improves the explainability and even rationality of their behaviors.
+6.1. Explainable Prompting Paradigms and Effects
+To better understand and control how LLMs generate outputs, many researchers directly prompt LLMs to provide the rationales along with their predictions (Wei et al., 2022; Huang et al., 2023a; Yao et al., 2023b; Besta et al., 2023; Kang et al., 2024).
+Chain-of-Thought (CoT) prompting (Wei et al., 2022) stands out by employing explicit reasoning to guide the generation process.
 Formally, we define the language model as $f_{\theta}$, and input prompt as $X=\{x_{1},y_{1},x_{2},y_{2},...,x_{n}\}$, where $x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1}$ denote the example question-response pairs for in-context learning, and $x_{n}$ is the actual question.
-In a standard question-answering scenario, we have the model output as $y_{n}=\operatorname*{arg\,max}_{Y}p_{\theta}(Y|x_{1},y_{1},x_{2},y_{2},...,x_{n})$.
-This approach, however, does not provide insights into the reasoning process behind the answer $y_{n}$.
-Therefore, the CoT method proposes to include human-crafted explanations $e_{i}$ for the $i$-th in-context example, resulting in a modified input format $X=\{x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}\}$.
-Given the input, the model will output not only $y_{n}$ but also the generated explanation $e_{n}$:
-$e_{n},y_{n}=\operatorname*{arg\,max}_{Y}p_{\theta}(Y|x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}).$ (5)
-Besides allowing for a more transparent and understandable interaction with LLMs, the CoT approach is also practically useful as it augments LLMs’ functionality by opening a window for users to control the models’ thought processes.
-Specifically, the usefulness of CoT methods lies in several key aspects:
-  • Reducing Errors in Reasoning: By breaking down complex problems into a series of smaller tasks, CoT reduces errors in complex calculations and logic-oriented tasks, leading to a more precise resolution of intricate problems (Wei et al., 2022; Qin et al., 2023; Zhang et al., 2023; Wang & Zhou, 2024b).
+In a standard prompting scenario, we have the model output as $y_{n}=\arg\max_{Y}p_{\theta}(Y|x_{1},y_{1},x_{2},y_{2},...,x_{n})$.
+However, this approach does not provide insight into how the answer $y_{n}$ is generated.
+Therefore, CoT proposes including human-crafted explanations $e_{i}$ for the $i$-th in-context example, resulting in a modified input format $X=\{x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}\}$.
+Given the input, the model will output not only prediction $y_{n}$ but along with an explanation $e_{n}$: $e_{n},y_{n}=\arg\max_{Y}p_{\theta}(Y|x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},...,x_{n}).$ Specifically, the usefulness of CoT methods lies in several key aspects:
+  • Reducing Errors in Reasoning: By breaking down complex problems into a series of smaller tasks, CoT reduces errors in logic-oriented tasks, leading to a more precise resolution of intricate problems (Wei et al., 2022; Qin et al., 2023; Zhang et al., 2023; Wang and Zhou, 2024).
   • Providing Adjustable Intermediate Steps: CoT enables the outlining of traceable intermediate steps within the problem-solving process.
 This feature enables users to trace the model’s thought process from inception to conclusion, and to adjust the prompts if undesirable model behaviors are observed (Lyu et al., 2023; Wang et al., 2023d).
-  • Facilitating Knowledge Distillation: The step-by-step reasoning processes derived from larger LLMs can serve as a specialized fine-tuning dataset for smaller LLMs.
-It allows smaller models to learn complex problem-solving by following explanations, effectively teaching them to tackle intricate questions with enhanced reasoning capabilities (Magister et al., 2022).
-6.2 Extended Methods of Explainable Prompting
-Advanced techniques beyond CoT methods have been developed to broaden the range of reasoning paths available to LLMs towards enhancing the transparency and understandability of the decision-making process (Yao et al., 2023b; Besta et al., 2023; Yao et al., 2023a; Dhuliawala et al., 2023; Lyu et al., 2023).
-We introduce several notable examples below.
-Tree-of-Thoughts (ToT).
-Proposed by Yao et al.
-(2023b), ToT advances beyond the traditional linear Chain of Thought reasoning, offering a more versatile structure that allows models to navigate through multiple reasoning paths.
-ToT makes the reasoning process of LLMs more interpretable by closely aligning it with human thought processes, as humans naturally consider multiple options and possible outcomes in both forward planning and retrospective analysis to reach conclusions (Sloman, 1996; Stanovich, 1999).
-This capability enhances the capacity of LLMs to tackle complex challenges that require the ability to consider and reevaluate different strategies, such as devising game strategies or generating creative content.
-By simulating the way humans think and make decisions, ToT not only makes their thought process more understandable to human users, but also improves the models’ effectiveness in handling complex tasks.
-Graph of Thoughts (GoT).
-Proposed by Besta et al.
-(2023), GoT transforms the output of LLMs into a graph format.
-This format visualizes information pieces as nodes and their connections as edges, enabling a more intricate and connected form of reasoning compared to previous methods CoT and ToT.
-By organizing data into nodes (individual concepts or pieces of information) and edges (relationship between these concepts), GoT makes the logical connections within complex systems more understandable (Yao et al., 2023a).
-This graphical representation brings several benefits for understanding complex information.
-Firstly, it enables dynamic modification of relationships between concepts, offering a clear visualization of how changing one element affects the others.
-This is crucial in fields like legal reasoning (Cui et al., 2023; Boche et al., 2024), scientific research (Ding et al., 2023; Choudhury et al., 2023), and policy analysis (Chen et al., 2023c), where the inter-dependencies between various factors can be intricate and subtle.
-Secondly, GoT enables an assessment of the significance of each node within the graph, providing insights into which pieces of information are most critical to the task.
-This level of adaptability and clarity makes GoT exceptionally powerful for analyzing and navigating complex information networks.
-6.3 Case Study: Is CoT Really Making LLM Inferences Explainable?
-6.3.1 Background and Experimental Settings
+  • Facilitating Knowledge Distillation: The step-by-step reasoning derived from larger LLMs can serve as a specialized fine-tuning dataset for smaller LLMs.
+It allows smaller models to learn complex problem solving by following explanations, effectively teaching them to tackle intricate questions with enhanced reasoning capabilities (Magister et al., 2022).
+In addition, techniques beyond simple CoT have been developed to broaden the range of reasoning paths available to LLMs (Zelikman et al., 2022; Yao et al., 2023b; Besta et al., 2023; Yao et al., 2023a; Chen et al., 2024a; Zelikman et al., 2024).
+For example, Tree-of-Thoughts (ToT) (Yao et al., 2023b) advances beyond the traditional linear chain-of-thought reasoning, offering a more versatile structure that allows models to navigate through multiple reasoning paths.
+Similarly, Graph of Thoughts (GoT) (Besta et al., 2023) transforms the outputs into a graph, where information pieces are nodes and their connections are edges, enabling a more intricate and connected form of reasoning paths.
+By organizing data into nodes (individual concepts or pieces of information) and edges (relationships between these concepts), GoT makes the logical connections within complex systems more understandable (Yao et al., 2023a).
+However, the above prompting-based methods may not always trigger LLMs to provide explanations.
+Thus, many efforts (Cheng et al., 2024; Zhang et al., 2024a; Pang et al., 2024; Yang et al., 2024) have been made to empower LLMs to acquire the inherent ability of providing their reasoning paths.
+DeepSeek-R1 (Guo et al., 2025) is one of the remarkable checkpoints, where they train LLMs to explain before making final predictions, and only give a positive reward if it provides a correct prediction.
+Even though there are no human-annotated explanations as ground truth to guide the LLM training, the researchers found that LLM can still learn to provide human-understandable explanations and achieve high accuracy in solving reasoning problems.
+To ensure the efficiency and faithfulness of LLM-generated explanations, recent works focus on encouraging LLMs to provide a shorter but necessary explanation (Luo et al., 2025; Hou et al., 2025; Yang et al., 2025a).
+6.2. Case Study: Is CoT Really Making LLM Inferences Explainable?
+6.2.1. Background and Experimental Settings
 Despite the apparent intuitiveness of the CoT prompt design, a critical question remains unanswered: Does CoT really make LLM inferences explainable?
@@
 In QA systems, answering multi-hop questions remains a significant challenge.
-Instead of leveraging a single information source, multi-hop questions require synthesizing information from multiple pieces or sources of data into a coherent and logical sequence.
-While LLMs show good performance in single-hop QA tasks (Radford et al., 2019), their efficacy significantly declines in multi-hop situations (Tan et al., 2023; Kim et al., 2023a; Zhong et al., 2023).
-This discrepancy highlights the need for more advanced methods to effectively handle the intricacy of multi-hop reasoning.
+Instead of leveraging a single information source, multi-hop questions require synthesizing information from multiple pieces or sources of data into a coherent and logical sequence (Tan et al., 2023; Kim et al., 2023b; Zhong et al., 2023).
+Figure 5. CoT Prompts for Multi-hop QA.
+Figure 6. CoT Faithfulness for Explanation.
 CoT Prompts for Multi-hop QA.
-To address the above challenge, our case study applies the CoT technique.
-CoT relies on high-quality templates as the in-context learning prompt, where we give an example as below.
-Here, $[x]$ denotes the test question.
+To address this, our case study applies the CoT technique with in-context examples.
+We provide an example as follows, where $[x]$ denotes the question.
 The “Thoughts” following each “Question” are step-by-step problem-solving statements for the multi-hop questions.
-The thoughts in the templates align the generation process of LLMs with human cognitive problem-solving patterns.
-Question: What is the capital of the country where Plainfield Town Hall is located?
-Thoughts: Plainfield Town Hall is located in the country of the United States of America. The capital of the United States is Washington, D.C.
-Answer: Washington, D.C.
-Question: Who has ownership of the developer of the Chevrolet Corvette (C4)?
-Thoughts: The developer of Chevrolet Corvette (C4) is Chevrolet. Chevrolet is owned by General Motors.
-Answer: General Motors
-Question: [x]
-CoT Faithfulness for Explanation: To quantitatively measure the faithfulness of CoTs, we select fidelity as the corresponding metrics (Zhao et al., 2023b; Wachter et al., 2017):
-$Fidelity=\frac{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)-\mathbbm{1}\left(\hat{\mathbf{y}}_{i}^{mislead}=\mathbf{y}_{i}\right)\right)}{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)\right)}\times 100\%,$ (6)
-where $\mathbf{y}_{i}$ denotes the ground truth label, $\hat{\mathbf{y}}_{i}$ denotes the original model output with CoT, while $\hat{\mathbf{y}}_{i}^{mislead}$ denotes the model output with misleading information inserted in the "Thoughts" section.
+CoT Faithfulness for Explanation: To quantitatively measure the faithfulness of CoTs, we select fidelity as the corresponding metrics (Zhao et al., 2023a; Wachter et al., 2017):
+$\enspace\enspace\enspace\enspace\enspace\enspace\enspace\enspace Fidelity=\frac{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)-\mathbbm{1}\left(\hat{\mathbf{y}}_{i}^{mislead}=\mathbf{y}_{i}\right)\right)}{\sum_{i=1}^{N}\left(\mathbbm{1}\left(\hat{\mathbf{y}}_{i}=\mathbf{y}_{i}\right)\right)}\times 100\%,$
+where $\mathbf{y}_{i}$ denotes the ground truth label, $\hat{\mathbf{y}}_{i}$ denotes the original model output with CoT, while $\hat{\mathbf{y}}_{i}^{mislead}$ denotes the model output with misleading information inserted.
 In the following, we give an example.
@@
 On the other hand, if it generates an answer corresponding to the incorrect thoughts, then we claim the thoughts are faithful.
-Question: What is the capital of the country where Plainfield Town Hall is located?
-Thoughts: Plainfield Town Hall is located in the country of the United States of America. The capital of United States is Washington, D.C.
-Answer: Washington, D.C.
-Question: Who has ownership of the developer of the Chevrolet Corvette (C4)?
-Thoughts: The developer of Chevrolet Corvette (C4) is Chevrolet. Chevrolet is owned by General Motors.
-Answer: General Motors
-Question: [Who is the head of state of the country where Ellie Kemper holds a citizenship?]
-Thoughts: Ellie Kemper is a citizen of Croatia. The head of state in Croatia is Zoran Milanović.
-Answer:
 Experimental Settings.
-We evaluate the performance on the MQUAKE-CF dataset Zhong et al.
-(2023), which includes 1,000 cases for each $K$-hop questions, $K\in\{2,3,4\}$, which totally consists of 3,000 questions.
-Our evaluation applies various language models, including GPT-2 (Radford et al., 2019) with 1.5 billion parameters, GPT-J (Wang & Komatsuzaki, 2021) with 6 billion parameters, LLaMA (Touvron et al., 2023a) with 7 billion parameters, Vicuna-v1.5 (Chiang et al., 2023) with 7 billion parameters, LLaMA2-chat-hf (Touvron et al., 2023b) with 7 billion parameters, Falcon (Almazrouei et al., 2023) with 7 billion parameters, Mistral-v0.1 (Jiang et al., 2023) with 7 billion parameters, and Mistral-Instruct-v0.2 (Jiang et al., 2023) with 7 billion parameters.
+We evaluate the performance on the MQUAKE-CF dataset (Zhong et al., 2023), which includes 1,000 cases for each $K$-hop questions, $K\in\{2,3,4\}$, which totally consists of 3,000 questions.
+Our evaluation applies various language models, including GPT-2 (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021) with 6 billion parameters, LLaMA (7B) (Touvron et al., 2023a), Vicuna-v1.5 (7B) (Chiang et al., 2023), LLaMA2-chat-hf (7B) (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023) with 7 billion parameters, Mistral-v0.1 (7B) (Jiang et al., 2023), and Mistral-Instruct-v0.2 (7B) (Jiang et al., 2023).
 These models have demonstrated proficiency in both language generation and comprehension.
-6.3.2 Experiment Results
-Performance Improvement.
-The performance reported in Table 5 for multi-hop question answering highlights the effectiveness of the CoT across various models.
-GPT-J exhibits the most dramatic improvement, particularly in 3-hop questions with over a 200% increase in accuracy, suggesting that CoT’s coherent reasoning greatly enhances LLMs’ question-answering ability.
-While GPT-2 shows modest gains, the performance of GPT-J and other models, such as LLaMA, Vicuna-v1.5, LLaMA2, Falcon, Mistral-v0.1, and Mistral-v0.2, indicates that the CoT method may be more beneficial for more advanced models.
-LLaMA2 and Mistral-v0.2 display considerable improvements of 78.4% and 63.8%, respectively, in 3-hop questions, further supporting this observation.
-Overall, the results underscore the potential of coherent reasoning techniques, i.e., CoT, in advancing the question-answering capabilities of LLMs across different model architectures and sizes.
-Table 5: Multi-hop question answering performance on MQUAKE-CF.
+6.2.2. Experiment Results
+Table 3. CoT Faithfulness Evaluation on MQUAKE-CF.
 Datasets MQUAKE-CF
 Question Type 2-hops 3-hops 4-hops
-Edited Instances Base Enhanced Improve Base Enhanced Improve Base Enhanced Improve
-GPT-2 (1.5B) 13.6 15.9 16.9% $\uparrow$ 11.6 8.9 23.3% $\downarrow$ 7.0 8.4 20.0% $\uparrow$
-GPT-J (6B) 23.1 51.9 124.7% $\uparrow$ 10.1 30.5 202.0% $\uparrow$ 21.3 49.8 133.8% $\uparrow$
-LLaMA (7B) 47.7 65.1 36.4% $\uparrow$ 29.6 39.3 32.8% $\uparrow$ 52.4 62.9 20.0% $\uparrow$
-Vicuna-v1.5 (7B) 41.3 56.3 36.3% $\uparrow$ 22.7 29.7 30.9% $\uparrow$ 31.6 53.1 68.2% $\uparrow$
-LLaMA2 (7B) 36.7 58.7 60.0% $\uparrow$ 17.0 30.3 78.4% $\uparrow$ 29.2 49.1 68.1% $\uparrow$
-Falcon (7B) 42.3 61.7 45.7% $\uparrow$ 23.2 31.7 35.7% $\uparrow$ 33.3 48.6 45.7% $\uparrow$
-Mistral-v0.1 (7B) 49.0 69.3 41.5% $\uparrow$ 30.0 42.3 41.1% $\uparrow$ 48.7 63.2 29.9% $\uparrow$
-Mistral-v0.2 (7B) 44.0 56.3 28.0% $\uparrow$ 23.0 37.7 63.8% $\uparrow$ 32.9 56.2 70.9% $\uparrow$
+Edited Instances C M F C M F C M F
+GPT-2 (1.5B) 15.9 5.2 67.3% 8.9 2.9 67.4% 8.4 1.3 84.5%
+GPT-J (6B) 51.9 7.3 85.9% 30.5 1.8 94.1% 49.8 2.0 96.0%
+LLaMA (7B) 65.1 9.9 84.8% 39.3 6.1 84.5% 62.9 6.0 90.5%
+Vicuna-v1.5 (7B) 56.3 21.7 61.5% 29.7 12.7 57.3% 53.1 16.1 69.7%
+LLaMA2 (7B) 58.7 17.0 71.0% 30.3 8.3 72.5% 49.1 12.0 75.6%
+Falcon (7B) 61.7 24.0 61.1% 31.6 15.0 52.6% 48.6 23.1 52.4%
+Mistral-v0.1 (7B) 69.3 24.0 65.4% 42.3 13.0 69.3% 63.2 18.4 70.8%
+Mistral-v0.2 (7B) 56.3 47.9 14.8% 37.7 22.0 41.6% 56.2 37.3 33.6%
+C: CoT; M: Misleaded CoT; F: Fidelity Score.
 Faithfulness Evaluation of CoT.
@@
 GPT-J and LLaMA exhibit high fidelity scores across different question types, indicating a strong adherence to the given reasoning paths.
-Conversely, models such as Vicuna-v1.5, LLaMA2, Falcon, Mistral-v0.1, and Mistral-v0.2 show relatively high mislead accuracy scores with lower fidelity scores.
+Conversely, other models show relatively high mislead accuracy scores with lower fidelity scores.
 In the experiments, we observe that these models usually rely on their own generated thoughts instead of using incorrect information provided in the CoT.
@@
 As a result, they are more likely to generate their own correct reasoning paths.
-In conclusion, the generated thoughts can be generally viewed as faithful explanations of their output answer.
-While high fidelity scores generally indicate a model’s adherence to the provided CoT, low fidelity scores do not necessarily imply a lack of faithfulness, especially when the model demonstrates the ability to reject misleading information.
-Further research on CoT faithfulness and the development of more sophisticated evaluation metrics could contribute to the advancement of interpretable and reliable language models.
-Table 6: CoT Faithfulness Evaluation on MQUAKE-CF.
-Datasets MQUAKE-CF
-Question Type 2-hops 3-hops 4-hops
-Edited Instances Correct Mislead Fidelity Correct Mislead Fidelity Correct Mislead Fidelity
-GPT-2 (1.5B) 15.9 5.2 67.3% 8.9 2.9 67.4% 8.4 1.3 84.5%
-GPT-J (6B) 51.9 7.3 85.9% 30.5 1.8 94.1% 49.8 2.0 96.0%
-LLaMA (6B) 65.1 9.9 84.8% 39.3 6.1 84.5% 62.9 6.0 90.5%
-Vicuna-v1.5 (7B) 56.3 21.7 61.5% 29.7 12.7 57.3% 53.1 16.1 69.7%
-LLaMA2 (7B) 58.7 17.0 71.0% 30.3 8.3 72.5% 49.1 12.0 75.6%
-Falcon (7B) 61.7 24.0 61.1% 31.6 15.0 52.6% 48.6 23.1 52.4%
-Mistral-v0.1 (7B) 69.3 24.0 65.4% 42.3 13.0 69.3% 63.2 18.4 70.8%
-Mistral-v0.2 (7B) 56.3 47.9 14.8% 37.7 22.0 41.6% 56.2 37.3 33.6%
-6.4 Challenges
-Within machine learning, explanation faithfulness refers to the degree to which an explanation accurately reflects the decision-making process of the model it is intended to elucidate (Li et al., 2022c).
-An explanation is considered as faithful if it causes the model to make the same decision as the original input.
+6.3. Challenges
+An explanation is considered as faithful if it causes the model to make the same decision as the original input (Li et al., 2022c).
 In this context, the challenge faced by explainable prompting (e.g., CoT prompt) lies in two aspects: (1) directing language models to generate explanations that are genuinely representative of the models’ internal decision-making processes, and (2) preventing language models from depending on potentially biased CoT templates.
@@
 Therefore, these rationales do not accurately represent the decision-making process within these models.
-Some efforts have been made to bolster the CoT capabilities of smaller language models by implementing instruction tuning with CoT rationales (Kim et al., 2023b; Ho et al., 2022).
-These methods can help improve the explanation faithfulness of CoT for small language models, thereby addressing this issue to some extent.
+Some efforts have been made to bolster the CoT capabilities of smaller language models by implementing instruction tuning with CoT rationales (Kim et al., 2023a; Ho et al., 2022).
 Nevertheless, it remains a challenging problem of how to ensure the generated explanations (i.e., “what the model says”) are faithful to the internal mechanism (i.e., “what the model thinks”) of language models.
 Regarding the second challenge, recent research shows that explanations in the CoT can be heavily influenced by the introduction of biasing prompt templates into model input (Turpin et al., 2024).
-This is because existing CoT requires carefully designed templates to prompt language models to produce explanations.
 If incorrect or biased information is encoded in such templates, the generated explanations could be misleading.
-Recently, Wang & Zhou (2024a) propose a novel decoding strategy to implement CoT with prompting, which could mitigate this issue.
-However, how to effectively help language models get rid of the template reliance still remains to be underexplored.
-7 LLM Enhancement via Knowledge-Augmented Prompting
+Recently, (Wang and Zhou, 2024) propose a novel decoding strategy to implement CoT with prompting, which could mitigate this issue.
+However, how to effectively help language models get rid of the template reliance still remains underexplored.
+7. Effective LLM Inference via Knowledge-Augmented Prompting
 Enhancing models with external knowledge can significantly improve the control and interpretability of decision-making processes.
 While LLMs acquire extensive knowledge through pre-training on web-scale data, this knowledge is embedded implicitly within the model parameters, making it challenging to explain or control how this knowledge is utilized during inference.
-Additionally, LLMs may not always encompass the unique knowledge specific to certain domains, nor keep pace with the constantly evolving information in the world.
-To address these limitations, this section discusses Retrieval-Augmented Generation (RAG) for the explicit integration of external knowledge into the mechanism of LLM-based AI systems, aiming to yield more interpretable predictions.
-7.1 Preliminaries: Retrieval-Augmented Generation
-By fetching relevant information from external databases or the internet, RAG ensures that LLM outputs are accurate and up-to-date.
-It addresses LLMs’ limitation of relying on fixed and potentially outdated knowledge bases.
-RAG operates in two steps: (1) Retrieval: It locates and fetches pertinent information from an external source based on the user’s query; (2) Generation: It incorporates this information into the model’s generated response.
-Given an input query $x$ and the desired output $y$, the objective function of RAG can be formulated as (Guu et al., 2020):
-$\max_{\phi,\theta}\,\log p(y|x)=\max_{\phi,\theta}\,\log\sum_{z\in\mathcal{K}}p_{\phi}(y|x,z)\cdot p_{\theta}(z|x),$ (7)
-where $z$ stands for the external knowledge retrieved from a knowledge base $\mathcal{K}$.
+To address these limitations, this section discusses Retrieval-Augmented Generation (RAG) for the explicit integration of external knowledge into LLMs, aiming to yield more interpretable predictions.
+7.1. Retrieval-Augmented Generation with Knowledge
+Generally, RAG operates in two steps: (1) Retrieval: It locates and fetches pertinent information from an external source based on the user’s query; (2) Generation: It incorporates this information into the model’s generated response.
+Given an input query $x$ and the desired output $y$, the objective function of RAG can be formulated as (Guu et al., 2020): $\max_{\phi,\theta}\,\log p(y|x)=\max_{\phi,\theta}\,\log\sum_{z\in\mathcal{K}}p_{\phi}(y|x,z)\cdot p_{\theta}(z|x),$ where $z$ stands for the external knowledge retrieved from a knowledge base $\mathcal{K}$.
 Thus, the target distribution is jointly modeled by a knowledge retriever $p_{\theta}(z|x)$ and an answer reasoning module $p_{\phi}(y|x,z)$.
@@
 In this case, RAG can be treated as a data-centric problem:
-$\max_{z\in\mathcal{K}}\,\log p(y|z,x)=\max_{z\in\mathcal{K}}\,\frac{p(z|x,y)}{p(z|x)}p(y|x),$ (8)
+(12) $\max_{z\in\mathcal{K}}\,\log p(y|z,x)=\max_{z\in\mathcal{K}}\,\frac{p(z|x,y)}{p(z|x)}p(y|x),$
 where the goal is to find appropriate knowledge that supports the desired output.
 The interpretability of RAG-based models comes from the information in $z$: (1) $z$ usually elucidates or supplements the task-specific information in $x$; (2) $z$ could explain the generation of output $y$.
-Unlike other deep models that directly estimate $p(y|x)$ in an end-to-end manner, where the decision process is not comprehensible, the RAG process provides justification or rationale $z$ that supports the result.
-Existing Retrieval-Augmented Generation (RAG) approaches can be categorized based on when they integrate external knowledge into the model’s workflow.
-The first category incorporates external knowledge at the inference stage.
-For instance, Karpukhin et al.
-(2020) employ dense vectors to identify related documents or text passages, enhancing the data retrieval step of RAG.
-Similarly, Lewis et al.
-(2020) refine the data retrieval process to ensure only the most pertinent information influences the model’s output.
-The second category integrates external knowledge during the model tuning stage.
-Some representative approaches include Guu et al.
-(2020); Borgeaud et al.
-(2022); Nakano et al.
-(2021).
-Generally, these methods embed a retrieval mechanism into the model’s training phase, enabling the model to utilize external data more efficiently from the outset.
-7.2 Enhancing Decision-Making Control with Explicit Knowledge
-The incorporation of explicit external knowledge through RAG enhances the precision and controllability of decision-making in LLMs.
-This method leverages real-time information from external databases to produce responses that are not only accurate but also tailored to the specific requirements of each query.
-Below, we explore the mechanisms by which RAG achieves a more controllable and directed content generation process, with references to key papers that have contributed to these advancements.
-7.2.1 Reducing Hallucinations in Response
-“Hallucination" in the context of LLMs refers to instances where these models generate information that, while coherent and contextually appropriate, is not based on factual accuracy or real-world evidence (Huang et al., 2023c).
+Unlike other models that estimate $p(y|x)$ in an end-to-end manner, where the decision process is not comprehensible, the RAG process provides justification or rationale $z$ that supports the result.
+7.2. Enhancing Decision-Making Control with Explicit Knowledge
+7.2.1. Reducing Hallucinations in Response
+“Hallucination” in the context of LLMs refers to instances where these models generate information that, while coherent and contextually appropriate, is not based on factual accuracy or real-world evidence (Huang et al., 2023d).
 This issue can lead to the production of misleading or entirely fabricated content, posing a significant challenge to the reliability and trustworthiness of LLMs’ outputs.
@@
 This process significantly enhances the factual basis of the model’s outputs, thereby reducing the occurrence of hallucinations.
-Shuster et al.
-(2021) applies neural-retrieval-in-the-loop architectures to knowledge-grounded dialogue, which significantly reduces factual inaccuracies in chatbots, as confirmed by human evaluations.
-Siriwardhana et al.
-(2023) introduces RAG-end2end, which joint trains retriever and generator components together.
+(Shuster et al., 2021) applies neural-retrieval-in-the-loop architectures to knowledge-grounded dialogue, which significantly reduces factual inaccuracies in chatbots, as confirmed by human evaluations.
+(Siriwardhana et al., 2023) introduces RAG-end2end, which joint trains retriever and generator components together.
 Their method demonstrates notable performance improvements across specialized domains like healthcare and news while reducing knowledge hallucination.
-7.2.2 Dynamic Responses to Knowledge Updating
-RAG empowers LLMs with the ability to incorporate the most current information, keeping their decision-making processes aligned with the latest developments.
+7.2.2. Dynamic Responses to Knowledge Updating
+RAG empowers LLMs by incorporating the most up-to-date information, keeping their decision-making aligned with the latest developments.
 This feature is especially vital in fast-evolving fields such as medicine and technology, where the need for timely and accurate information is paramount (Meng et al., 2022b).
-For example, research by (Izacard & Grave, 2020) demonstrates significant enhancements in output relevance and accuracy through real-time information retrieval.
-Similarly, Han et al.
-(2023) suggest using retrieved factual data to correct and update the knowledge within pre-trained LLMs efficiently.
-Additionally, Wang et al.
-(2023e) introduce a method for integrating newly retrieved knowledge from a multilingual database directly into the model prompts, facilitating updates in a multilingual context.
-7.2.3 Domain-specific Customization
+For example, research by (Izacard and Grave, 2020) demonstrates significant enhancements in output relevance and accuracy through real-time information retrieval.
+Similarly, (Han et al., 2023) suggest using retrieved factual data to correct and update the knowledge within pre-trained LLMs efficiently.
+Additionally, (Wang et al., 2023b) introduce a method for integrating newly retrieved knowledge from a multilingual database directly into the model prompts, facilitating updates in a multilingual context.
+7.2.3. Domain-specific Customization
 RAG enhances LLMs by incorporating knowledge from specialized sources, enabling the creation of models tailored to specific domains.
-Research by Guu et al.
-(2020) illustrates how integrating databases specific to certain fields into the retrieval process can empower models to deliver expert-level responses, boosting their effectiveness in both professional and academic contexts.
-Shi et al.
-(2023) have applied this concept in the medical domain with MedEdit, utilizing an in-context learning strategy to merge relevant medical knowledge into query prompts for more accurate medical advice.
+Research by (Guu et al., 2020) illustrates how integrating databases specific to certain fields into the retrieval process can empower models to deliver expert-level responses, boosting their effectiveness in both professional and academic contexts.
+(Shi et al., 2023) have applied this concept in the medical domain with MedEdit, utilizing an in-context learning strategy to merge relevant medical knowledge into query prompts for more accurate medical advice.
 Moreover, recent research finds that LLMs struggle to capture specific knowledge that is not widely discussed in the pre-training data.
-Specifically, Mallen et al.
-(2023) observe that LLMs often fail to learn long-tail factual knowledge with relatively low popularity, finding that simply increasing model size does not significantly enhance the recall of such information.
-However, they note that retrieval-augmented LLMs surpass much larger models in accuracy, particularly for questions on well-known subjects, suggesting that this method can effectively bridge knowledge gaps.
-Similarly, Kandpal et al.
-(2023) highlights LLMs’ challenges with acquiring rare knowledge and proposes that retrieval augmentation offers a viable solution, minimizing reliance on extensive pre-training for capturing nuanced, less common information.
-7.3 Challenges
+Specifically, (Mallen et al., 2023) observe that LLMs often fail to learn long-tail factual knowledge with relatively low popularity, finding that simply increasing model size does not significantly enhance the recall of such information.
+However, they note that retrieval-augmented LLMs surpass much larger models in accuracy.
+Similarly, (Kandpal et al., 2023) highlights LLMs’ challenges with acquiring rare knowledge and proposes that retrieval augmentation offers a viable solution, minimizing reliance on extensive pre-training for capturing nuanced, less common information.
+7.3. Challenges
 We discuss the challenges in RAG that are relevant to its explainability aspects: (1) In the retrieval stage $p_{\theta}(z|x)$, does the retrieved information $z$ always elucidate the task-specific information contained in the input $x$?
 (2) In the generation stage $p_{\phi}(y|x,z)$, does $z$ effectively serve as an explanation for the generation of output $y$?
-Please note that our goal is not to exhaustively discuss all the limitations of RAG in this paper as RAG itself is a broad topic in NLP research.
-For a more detailed examination of the broader limitations of RAG, we direct readers to other reviews (Gao et al., 2023).
-7.3.1 Retrieval Accuracy Bottlenecks
-Existent RAG methods typically rely on similarity search to pinpoint relevant information (Lewis et al., 2020; Gao et al., 2023), which represents a substantial improvement over basic keyword searches (Robertson et al., 2009).
+7.3.1. Retrieval Accuracy Bottlenecks
+Existent RAG methods typically rely on similarity search to pinpoint relevant information (Lewis et al., 2020; Gao et al., 2023b), which represents a substantial improvement over basic keyword searches (Robertson et al., 2009).
 However, these methods may struggle with complex queries that demand deeper comprehension and nuanced reasoning.
-The recent “lost-in-the-middle” phenomenon (Liu et al., 2024) has revealed that an ineffective retrieval can result in the accumulation of extraneous or conflicting information, negatively affecting the generation quality.
-To address this challenge, recent RAG approaches have integrated adaptive learning processes (Asai et al., 2023).
-This advancement enables the retrieval system to refine their performance over time through feedback, adapting to evolving language use and information updates, ensuring their responses remain relevant and accurate.
-Nonetheless, efficiently handling intricate and multi-hop questions remains a significant challenge, highlighting the need for ongoing research to enhance the capabilities of RAG systems.
-7.3.2 Controllable Generation Bottlenecks
-In-context learning stands out as the premier method for incorporating external knowledge to boost the capabilities of LLMs such as GPT-4 (Asai et al., 2023; Gao et al., 2023).
-Despite its effectiveness, there’s no surefire way to ensure that these models consistently leverage the provided external knowledge within the prompts for their decision-making processes.
-In practice, to achieve thorough coverage, commonly used dense retrieval usually returns a large volume of content, including both relevant and redundant information to the input question.
-Unfortunately, redundant information in the model prompt raises the computational cost and can mislead LLMs to generate incorrect answers.
-Recent research shows the retrieved information can degrade the question-answering task performance (Yoran et al., 2023; Petroni et al., 2020; Li et al., 2022a).
-Some recent work proposes to fine-tune the LLM to improve resilience to noise and reduce hallucinations.
-However, such approach still cannot prevent oversized retrieval information decrease the system interpretability (Yoran et al., 2023; Xu et al., 2023b).
+The recent “lost-in-the-middle” phenomenon (Liu et al., 2024a) has revealed that an ineffective retrieval can result in the accumulation of extraneous or conflicting information, negatively affecting the generation quality.
+Efficiently handling intricate and multi-hop questions remains a significant challenge, highlighting the need for ongoing research to enhance the capabilities of RAG systems.
+7.3.2. Controllable Generation Bottlenecks
+In-context learning stands out as the premier method for incorporating external knowledge to boost the capabilities of LLMs such as GPT-4 (Asai et al., 2023b; Gao et al., 2023b).
+Despite its effectiveness, there’s no surefire way to ensure that these models consistently leverage the provided external knowledge within the prompts for their decision-making processes (Wu et al., 2024b; Yoran et al., 2023; Petroni et al., 2020; Li et al., 2022b).
 The challenge of optimizing the use of external explanations to achieve more precise and controlled decision-making in LLMs is an ongoing issue that has yet to be fully addressed.
-8 Training Data Augmentation with Explanation
-This section explores the generation of synthetic data from explanations using large language models, a technique poised to enhance various machine learning tasks.
-In machine learning, limited data availability often constrains model performance, presenting a significant challenge across many domains.
-A viable solution is data augmentation, where LLMs, with their robust generative capabilities, can be utilized for text data augmentation (Whitehouse et al., 2023), such as transforming existing text samples into new variants (Dai et al., 2023).
-Nevertheless, there are several challenges to be tackled for effective text augmentation.
-First, for utility, the generated samples need to exhibit diversity compared to the original data.
-Second, these samples should be exhibit useful patterns relevant to the downstream tasks.
-To address these challenges, explanation methods offer a valuable tool to guide data augmentation by providing supplemental contexts and useful rationales (Carton et al., 2021).
-Using LLMs for explanation-guided data augmentation is a nascent but promising field.
-In this section, we aim to outline feasible frameworks and discuss potential applications, offering directions for future research in this field.
-Explanations can be particularly beneficial in data augmentation within two scenarios.
-In the first scenario, explanations are used to delineate desired model behaviors or to identify existing deficiencies, which effectively guides the data augmentation process of LLMs.
-The second scenario involves employing LLMs to directly produce explanatory texts, which serve as supplementary information to enrich the dataset.
-8.1 Explanation-guided Data Augmentation for Mitigating Shortcuts
-Machine learning models are prone to make predictions with spurious correlations, also known as shortcuts (Geirhos et al., 2020), which are misaligned with human reasoning processes.
-This dependency on shortcuts underlies various challenges in machine learning, notably diminishing a model’s ability to generalize to out-of-distribution samples and weakening its resilience against adversarial attacks (Xu et al., 2020).
-The extent to which a model depends on shortcut features is intrinsically linked to its interpretability.
-Essentially, if a model’s predictions are predominantly based on such unreliable features, it indicates that the underlying mechanisms of the model are not fully comprehensible or interpretable from a human perspective.
-Therefore, post-hoc explanation techniques usually play a crucial role in detecting undesirable correlations between input and predictions within deep models (Liu et al., 2018; 2021).
+8. Training Data Augmentation with Explanation
+This section explores the generation of synthetic data from explanations using LLMs, a technique poised to enhance various machine learning tasks.
+In machine learning, limited data availability often leads to data imbalance and scarcity, which in turn constrains model performance in many domains.
+A viable solution is data augmentation, where LLMs, with advanced generative capabilities, can be utilized (Whitehouse et al., 2023; Tan et al., 2024).
+Nevertheless, several challenges exist for effective text augmentation.
+First, for utility, the generated samples should exhibit diversity compared to the original data.
+Second, these samples should exhibit useful patterns relevant to downstream tasks.
+To this end, explanation tools could guide augmentations by providing supplemental context and rationales (Carton et al., 2021).
+Explanations can be particularly beneficial in two scenarios.
+In the first scenario, explanation tools are used to to identify existing deficiencies, which effectively guides the data augmentation.
+The augmented data include samples and related label annotations.
+The second scenario employs LLMs to directly produce explanatory texts, such as rationales, as supplementary information to enrich the dataset.
+8.1. Label-based Data Augmentation
+For learning tasks, the training dataset includes a large volume of data points, defined as $\mathcal{D}=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\}$.
+We categorize learning tasks into classification tasks and generation tasks.
+In traditional classification tasks, each data point $x_{i}$ has a label $y_{i}$ representing a class.
+Models are trained to predict the labels for new data points.
+However, these models are prone to make predictions with spurious correlations also known as shortcuts (Geirhos et al., 2020).
+Thus, model’s predictions are predominantly based on shortcut features, and the underlying mechanisms of the model are not interpretable from a human perspective.
+Post-hoc explanation techniques play a crucial role in detecting undesirable correlations between input and predictions (Liu et al., 2018, 2021).
 For example, Du et al.
-(2021) adopt Integrated Gradient (IG) to attribute a model’s predictions to its input features, showing that the model tends to treat functional words, numbers, and negation words as shortcuts and strongly rely on these shortcut features for prediction in natural language understanding tasks.
-Data augmentation can be applied to training better models that are less sensitive to shortcut features.
-Explanatory information such as counterfactuals (Wang & Culotta, 2021) has been incorporated in data augmentation to improve model robustness.
-It generates counterfactual samples by first identifying critical features (e.g., word tokens) and then replacing these features with their antonyms, along with reversing their associated labels.
-Subsequently, the generated samples are combined with the original ones to train downstream models.
-Furthermore, these techniques can be extended to enhance the out-of-distribution performance of smaller models (Sachdeva et al., 2023; Wen et al., 2022).
-Namely, large language models could serve as an effective tool to augment data.
-For example, LLMs are able to synthesize examples that represent outlier cases or rare occurrences, which helps smaller models generalize better on unseen data (Xu et al., 2023a).
-This could be helpful in building robust models in scenarios where data are scarce or confidential (Tang et al., 2023a).
-Besides, LLMs are promising in improving models’ safety by generating adversarial examples that are more valid and natural compared to conventional approaches (Wang et al., 2023f).
-First, the most vulnerable words are identified with attribution-based methods.
-Then, these words are replaced by LLMs in a manner that maintains the integrity of the original text.
-The quality of these examples can be examined with an external classifier.
-Subsequently, these adversarial examples are employed to train downstream models, effectively fortifying them against potential attacks and boosting their security.
-Similarly, LLMs are also helpful in mitigating biases such as fairness issues in models (He et al., 2023).
-This work claims that it automatically identifies underrepresented subgroups, and chooses an optimal augmentation strategy that avoids hurting other groups.
-New group examples are generated using LLMs with human-providing labels.
-The experiments observe improvements on both underrepresented groups and overall model performance.
-This approach prevents skewed datasets from leading to unfair outcomes for a certain demographic, thereby potentially promoting fairness in society.
-8.2 Explanation-enhanced Data Enrichment
-As a powerful generative tool, LLMs have been leveraged to directly generate natural language explanations as augmented data.
-This strategy relies on LLMs’ understanding abilities to assist smaller models in their learning tasks.
-One objective in such work is to add natural language explanation generated by LLMs to training data, so as to enhance the performance of small models.
+(2021) adopt Integrated Gradient (IG) to attribute a model’s predictions to its input features, showing that the model tends to treat functional words, numbers, and negation words as shortcuts and rely on them for prediction in natural language understanding tasks.
+LLMs can be applied to resolve data set issues by augmenting various data.
+For example, explanatory information such as counterfactuals has been incorporated to improve model robustness (Wang and Culotta, 2021) and out-of-distribution performance (Sachdeva et al., 2023).
+LLMs can also synthesize rare examples, which helps models to generalize better on unseen data (Xu et al., 2023b).
+Besides, another work utilizes attribution-based interpretability tool to identify vulnerable words and leverage LLMs to synthesize natural adversarial examples to enhance model safety (Wang et al., 2023c).
+Similarly, LLMs are also helpful in mitigating biases such as fairness issues in models.
+He et al.
+(2023) claims that it automatically identifies underrepresented subgroups, and use LLMs to augment theese subgroups while avoids hurting other groups.
+In text generation tasks, each data point consists of a prompt $x$ and a response $y$.
+The data used for model training is usually text including both prompt and response.
+Generative models are designed to predict next token based on existing input, which can be formulated as $y_{m}:\arg\max p_{\theta}(y_{m}|x,y_{(i\ :\ m-1)})$.
+To make models generate desired output, well-aligned models are indispensable.
+Recently, a few work employ LLMs to augment data and enhance model alignment.
+For example, Wang et al.
+(Wang et al., 2024b) identifies the weakness of existing datasets and instructs LLMs to generate desired data to improve model safety.
+Another work use attribution-based method to identify key information from privacy-relevant data and replace the privacy information with synthetic information, which provides a way to avoid privacy leakage (Zeng et al., 2024b).
+8.2. Rationale-based Data Enrichment
+LLMs have been leveraged to directly generate natural language explanations as augmented data.
+This strategy makes use of LLMs’ chain-of-thought reasoning abilities (Kojima et al., 2022).
+The training datasets contain data with annotated information, denotes as $\mathcal{D}=\{(x_{1},r_{1},y_{1}),\ \cdots,(x_{m},r_{m},y_{m})\}$.
+Each data point includes not only prompt $x$ and response $y$, but reasoning process or annotated information defined as $r$.
+As state-of-art LLMs are excellent in interpreting context, many studies have utilized them to provide annotations and enrich data.
+One line of research uses generated explanations to assist and guide model training.
 Li et al.
-(2022b) introduce explanations from LLMs that facilitate the training of smaller models to enhance their reasoning capabilities and acquire explanation generation abilities.
-Specifically, three explanation generation approaches from LLM are utilized to train smaller models, including (1) explanations generated through chain of thought prompting, (2) explanations by rationalization prompting conditioned on golden labels and (3) hybrid approach that combines the first two approaches.
-With the best case scenario in a reasoning case, a 9.5% improvement on accuracy compared to 60x larger GPT-3 has been seen (Li et al., 2022b).
-It is worth noting that LLMs including ChatGPT and GPT-4 are competitive in generating coherent and logical explanations, except for limited languages (Whitehouse et al., 2023).
-Explanations from LLMs have also been utilized to mitigate spurious correlations in aspect-based sentiment analysis tasks (Wang et al., 2023c).
-This study proposes using LLMs to generate explanations for the sentiment labels of aspects in sentences (Wang et al., 2023c).
-These explanations provide reasoning grounded in contextual semantics rather than relying on spurious correlations between words and labels.
-The explanations are integrated into the training of aspect-based sentiment analysis models through two methods: augmenting the training data with the explanations or distilling knowledge from the explanations through mimicking behaviors.
-By focusing on the explanations instead of superficial textual cues, the models can better learn the true associations between text and sentiment and become more robust, improving both in-domain performance and generalization ability (Wang et al., 2023c).
-Another line of work involves integrating LLM rationales as additional supervision to guide the training of smaller models.
-Experiments have shown that this approach not only requires fewer training data but also outperforms traditional fine-tuning and distillation methods (Hsieh et al., 2023).
-In addition to the existing application of augmentation techniques summarized above, we envision that incorporating additional text information can also be practical and efficient in enhancing the performance of various models.
-For example, one promising application lies in the realm of guiding the parameter learning process of small models by using automatically generated explanations.
-Previous research has investigated this avenue by directing the attention of natural language inference models towards human-crafted explanations (Stacey et al., 2022).
-As human-crafted explanations are both arduous and non-transferable, utilizing LLMs as generators presents a more economic and versatile alternation.
-Another potential application is to enhance model performance on complex tasks using natural language explanations from LLMs.
-For instance, code translation generation tasks incorporate explanations as an intermediate step, improving model performance by 12% on average (Tang et al., 2023b).
+(2022a) introduce LLM explanations to facilitate the training of smaller models, which improves their reasoning capabilities and acquires explanation generation abilities.
+Hsieh et al (Hsieh et al., 2023) integrate LLM rationales as additional supervision to guide the training of smaller models.
+Experiments have shown that this approach not only requires fewer training data but also outperforms traditional fine-tuning and distillation methods.
+Another line of study employ annotated information to boost models’ performance on specific tasks.
+For example, LLM explanations have also been utilized to generate explanations for sentiment labels of aspects in sentences to mitigate spurious correlations in aspect-based sentiment analysis tasks (Wang et al., 2023a).
+Besides, Code translation generation tasks incorporate explanations as an intermediate step, improving model performance by 12% on average (Tang et al., 2023).
 The result shows that explanations are particularly useful in zero-shot settings.
-Apart from assisting in smaller models, LLMs have demonstrated their ability to improve their own reasoning abilities by generating reliable rationales (Huang et al., 2022).
-Further, Krishna et al.
-(2023) embed post-hoc explanations, attributing scores to all input features, into natural language rationales.
-This approach improves model accuracy by 10-25% across multiple tasks via in-context learning.
-Another study explicitly investigates LLMs’ ability to generate post-hoc explanations in natural language.
-The experiments show that LLMs spot key features with 75% accuracy (Kroeger et al., 2023).
-These studies present a novel strategy for leveraging post-hoc explanations, beyond traditional natural language explanations, to enrich training data so as to bolster model performance.
-8.3 Challenges
-8.3.1 Computational Overhead
+Wei et al.
+(Wei et al., 2024a) instruct LLMs to explain the retrieval process from documents to produce rationales.
+Then they can be used to train models and further improve models ability on retrieval-augmented generation tasks.
+Similarly, Lupidi et al.
+(Lupidi et al., 2024) prompt LLMs to generate data explanations from real-world sources such as articles, which are then used to fine-tune models and improve their performance on multi-hop reasoning and information retrieval.
+8.3. Challenges
+8.3.1. Computational Overhead
 Conventional post-hoc explanations, built on well-trained models, are often resource-intensive tasks.
@@
 This process typically requires multiple rounds of model training and applying interpretability methods to develop fair and robust models.
-Consequently, the crafting process can be both time and energy-consuming.
-Given these challenges, exploring the development of data-centric evaluation metrics is crucial.
-These metrics can offer a more efficient way to assess data issues, bypassing traditional, cumbersome explanation methods.
-By focusing on these data-centric measurements, data issues can be diagnosed and fixed before training.
-The number of training rounds needed is then significantly reduced.
-This shift not only streamlines model development but also helps reduce computational overhead, making the whole process more practical and efficient.
-8.3.2 Data Quality and Volume
+Consequently, the crafting process can be both time- and energy-consuming.
+8.3.2. Data Quality and Volume
 Despite their advanced capabilities, LLMs still have limitations when dealing with highly specialized or niche contexts.
@@
 This could adversely affect the quality of augmented data, potentially introducing more biases to which LLMs are also vulnerable.
-Another challenge is controlling the relevance of LLM-generated content.
-That is, the explanations or data points may seem reasonable but often lack factual accuracy or nuances specific to a domain.
-Currently, we lack robust metrics to effectively measure the quality and relevance of these generated data relative to the original tasks.
 Determining the precise amount of data required is also challenging, often leading to new dataset imbalances.
-Managing the quality of LLM-generated data is an immense challenge, as augmented data can introduce other biases (Zhao & Vydiswaran, 2021).
+Managing the quantity of LLM-generated data is an immense challenge, as augmented data can introduce other biases (Zhao and Vydiswaran, 2021).
 This stems from LLMs’ limited ability to accurately control the quantity and distribution of generated data.
-Moreover, crafting effective prompts is more of an art than a science, adding uncertainty around generated data quality.
 Together, these factors underscore the complexities and challenges in fully harnessing LLMs’ potential for data augmentation and related tasks.
-9 Generating User-Friendly Explanation for XAI
+9. Generating User-Friendly Explanation with LLMs
 Previous sections mainly focused on quantitative explanations with LLM via numerical values.
-For example, sample-based explanation discussed in Section 4 aims to assign each training sample an influence score (see Eqs.1-4) that measures the confidence that we can use that training sample to explain the prediction of a test sample.
-However, using numerical values for explanations is not intuitive, which can be difficult to understand by practitioners with little domain knowledge (Latif & Zhai, 2024; Lee et al., 2023; Li et al., 2020).
-User-friendly explanations, on the contrary, aim to generate human-understandable explanations, e.g., natural language-based descriptions, regarding certain data, the reason why a model makes certain predictions, or what role a neuron plays in the network, such that the generated explanations can be well-understood by both researchers and practitioners.
-Given an explainee $e$, which can be a data sample $(x_{i},y_{i})$, a neuron $\theta_{i}$ from a pretrained model $f_{\theta}$, or a prediction result $\hat{y}$ based on the input $x$, generating user-friendly explanation aims to map the explainee $e$ to a sequence of natural language tokens as the explanation for the explainee $e$, such that the generated explanations can be easily comprehended by human beings.
-9.1 User-friendly Data Explanation with LLMs
-Data explanation refers to the process of translating difficult materials (e.g., program codes, long documents) into concise and straightforward language so that they are easy to understand by humans.
-Language models have long been used to generate explanations for textual data (Dai & Callan, 2019).
-Since modern LLMs are trained on corpora composed of codes, math, and papers, they can be leveraged to explain data beyond pure textual content.
-For example, Chen et al.
-(2021) have demonstrated that pretrained GPT models possess the ability to understand and generate codes, where explanatory comments are generated simultaneously that facilitate the understanding of programmers.
-In addition, Welleck et al.
-(2022) propose to explain math theorems by providing detailed derivations, so that the theorems are easier to understand.
-Recently, LLMs have also been used to elucidate academic papers (Castillo-González et al., 2022), making difficult content to be easily understood by individuals with little domain knowledge.
-9.2 Explaining Small Models with LLMs
+For example, sample-based explanation discussed in Section 4 aims to assign each training sample an influence score (see Eqs.8-11) that measures the confidence that we can use that training sample to explain the prediction of a test sample.
+However, using numerical values for explanations is not intuitive, which can be difficult to understand by practitioners with little domain knowledge (Latif and Zhai, 2024; Lee et al., 2023; Li et al., 2020; Chen et al., 2024c).
+User-friendly explanations, on the contrary, aim to generate human-understandable explanations, e.g., natural language-based descriptions, regarding why a model makes certain predictions or what role a neuron plays in the network, such that the explanations can be well-understood by both researchers and practitioners.
+9.1. Explaining Small Models with LLMs
 Recently, there has been growing interest in leveraging LLMs to generate free-text explanations for small models.
-For example, to explain black-box text classifiers, Bhattacharjee et al.
-(2023) propose a prompting-based strategy to identify keywords $K=\{k_{1},k_{2},...,k_{n}\}$ in the input texts $x$ with pretrained LLMs that are informative for the label $y$, and ask LLMs to substitute them with another set of keywords $K^{\prime}=\{k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{n}\}$, such that changed text $x^{\prime}$ changes the label prediction to $y^{\prime}$.
+For example, to explain black-box text classifiers, (Bhattacharjee et al., 2023) propose a prompting-based strategy to identify keywords $K=\{k_{1},k_{2},...,k_{n}\}$ in the input texts $x$ with pretrained LLMs that are informative for the label $y$, and ask LLMs to substitute them with another set of keywords $K^{\prime}=\{k^{\prime}_{1},k^{\prime}_{2},...,k^{\prime}_{n}\}$, such that changed text $x^{\prime}$ changes the label prediction to $y^{\prime}$.
 They view the textual mapping rule “if we change $K$ into $K^{\prime}$ in $x$, then $y$ will be classified as $y^{\prime}$” as the counterfactual explanation for the model.
@@
 The explaining ability of LLMs is not necessarily limited to text models.
-For example, Zhao et al.
-(2023a) propose using pretrained vision-language models to generate explanations for a neuron $\theta_{i}$ of an image classification model.
+For example, (Zhao et al., 2023b) propose using pretrained vision-language models to generate explanations for a neuron $\theta_{i}$ of an image classification model.
 Specifically, for each class $y=y_{c}$, they first find regions in images with label $y_{c}$ that have maximum activation of the neuron $\theta_{i}$ as the surrogate explainees for $\theta_{i}$, and prompt LLMs such as ChatGPT to generate candidate explanations (words, short phrases) for the class label $y_{c}$.
 Then, they use the pretrained vision-language model CLIP (Radford et al., 2021) to match the candidate explanations with the surrogate explainees as the explanations for the neuron $\theta_{i}$.
-Recently, LLMs have also found applications in explaining recommender systems Zhu et al.
-(2023a).
-Specifically, Yang et al.
-(2023c) found that LLMs can well interpret the latent space of sequential recommendation model after alignment, whereas Lei et al.
-(2023) propose to align user tokens of LLMs with the learned user embeddings of small recommendation model to generate explanations of user preferences encoded in the embeddings.
-Recently,  Schwettmann et al.
-(2024) propose a unified framework to explain all models where inputs and outputs can be converted to textual strings.
+Recently, LLMs have also found applications in explaining recommender systems (Zhu et al., 2023).
+Specifically, (Yang et al., 2023b) found that LLMs can well interpret the latent space of sequential recommendation model after alignment, whereas (Lei et al., 2023) propose to align user tokens of LLMs with the learned user embeddings of small recommendation model to generate explanations of user preferences encoded in the embeddings.
+Recently,  (Schwettmann et al., 2024) propose a unified framework to explain all models where inputs and outputs can be converted to textual strings.
 Specifically, the explainer LLM is used as an agent to interact with the explainee model by iteratively creating inputs and observing outputs from the model, where the textual explanations are generated by viewing all the interactions as the context.
-9.3 Self-Explanation of LLMs
-Due to the black-box nature of LLMs, it is promising to generate user-friendly explanations for the LLMs themselves, such that the operational mechanics and the predictions of LLMs can be well-understood by human experts.
+9.2. Self-Explanation of LLMs
+Due to the black-box nature of LLMs, it is promising to generate user-friendly explanations for the LLMs themselves, such that their operational mechanics and the predictions can be understood by human experts.
 Based on whether the LLM needs to be retrained to generate explanations for themselves, the self-explanation of LLM can be categorized into two classes: fine-tuning based approach and in-context based approach, which will be introduced in the following parts.
@@
 One exemplar method is P5 (Geng et al., 2022), which fine-tunes the pre-trained language model T5 (Raffel et al., 2020) on both the rating and explanation data to generate an explanation alongside the recommendations.
-Recently, several works have improved upon P5 (Cui et al., 2022; Zhu et al., 2024), which fine-tunes different LLMs such as GPT2, LLaMA, Vicuna, etc., and propose different prompt learning strategies (Li et al., 2023d) with generating explanation as the auxiliary task.
+Recently, several works have improved upon P5 (Cui et al., 2022; Zhu et al., 2024), which fine-tunes different LLMs such as GPT2, LLaMA, Vicuna, etc., and propose different prompt learning strategies (Li et al., 2023e) with generating explanation as the auxiliary task.
 With explanations introduced as additional supervision signals to fine-tune pretrained LLMs for recommendations, the performance can be improved with good explainability.
@@
 The application of in-context based self-explanation of LLMs is broad.
-For example, Huang et al.
-(2023d) explore generating zero-shot self-explanation of sentiment analysis with LLMs by directly asking them to generate explanations alongside the predictions.
-In addition, Huang et al.
-(2023a) propose a chain-of-explanation strategy that aims to explain how LLMs can detect hate speech from the textual input.
-Lu et al.
-(2022) find that CoT can generate well-supported explanations for question answering with scientific knowledge.
-9.4 Challenges
-9.4.1 Usability v.s. Reliability
+For example, (Huang et al., 2023c) explore generating zero-shot self-explanation of sentiment analysis with LLMs by directly asking them to generate explanations alongside the predictions.
+In addition, (Huang et al., 2023a) propose a chain-of-explanation strategy that aims to explain how LLMs can detect hate speech from the textual input.
+(Lu et al., 2022) find that CoT can generate well-supported explanations for question answering with scientific knowledge.
+9.3. Challenges
+9.3.1. Usability v.s. Reliability
 Many existing methods rely on prompts to generate user-friendly explanations, which are not as reliable as numerical methods with good theoretical foundations.
-Ye & Durrett (2022b) find that the explanations by CoT may not be factually grounded in the inputs.
-Therefore, they believe that these explanations are more suitable as post-hoc explanations regarding why the LLM makes certain predictions (no matter whether the predictions are right or wrong).
-However, the validity of viewing CoT explanations as post-hoc justifications has been questioned by recent findings from Turpin et al.
-(2024), which uses biased datasets (e.g., the few-shot examples in the prompt always answer “A” for multiple choice questions) to show that the generated explanations may be plausible, but systematically unfaithful to represent the true reasoning process of the LLMs.
-This issue arises from biased reasoning steps which inadvertently skew the predictions.
-Thus, there’s a growing need for more theoretical scrutiny of user-friendly explanations to ensure faithfulness and credibility.
-9.4.2 Constrained Application Scenarios
+(Ye and Durrett, 2022) find that the explanations by CoT may not be factually grounded in the inputs.
+Therefore, they believe that these explanations are more suitable as post-hoc explanations regarding why the LLM makes certain predictions.
+However, the validity of viewing CoT explanations as post-hoc justifications has been questioned by recent findings (Turpin et al., 2024), which uses biased datasets (e.g., the few-shot examples in the prompt always answer “A” for multiple choice questions) to show that the generated explanations may be plausible, but systematically unfaithful to represent the true reasoning process of the LLMs.
+Therefore, there’s a growing need for more theoretical scrutiny of user-friendly explanations to ensure faithfulness and credibility.
+9.3.2. Constrained Application Scenarios
 Currently, the utilization of LLMs to explain smaller black-box models is mainly limited to those that deal with data with rich textual information (Bhattacharjee et al., 2023; Lei et al., 2023).
-Although Zhao et al.
-(2023b) propose a strategy to explain image classifiers, the ability to match candidate textual explanations with image patterns still relies on the pretrained vision-language model CLIP.
-This method may not be applicable to other domains such as graph machine learning (e.g., graph neural networks), or time-series analysis (recurrent neural networks), where large pretrained models have demonstrated little progress compared to natural language processing and computer vision.
-Therefore, there is a compelling need to devise more versatile strategies for explaining models across a wider range of fields.
+Although (Zhao et al., 2023a) propose a strategy to explain image classifiers, the ability to match candidate textual explanations with image patterns still relies on the pretrained vision-language model CLIP.
+Therefore, there is a compelling need to development more versatile strategies for explaining models across a wider range of fields.
 This endeavor could depend on the fundamental research on combining LLM with other domain-specific tasks, such as the development of Graph-Language Models that are applicable to unseen graphs in a zero-shot manner.
-10 LLMs for Interpretable AI System Design
+10. Interpretable AI System Design with LLMs
 An intriguing but challenging problem in XAI is creating model architectures or even AI systems that are inherently interpretable (Rudin, 2019), where different model components represent clear and comprehensible concepts or functionalities that are easily distinguishable from one another.
-Machine learning models such as support vector machines (Hearst et al., 1998) and tree-based models (Song & Ying, 2015) were classical techniques for achieving model interpretability.
-In the deep learning era, typical research areas in this context include concept-bottleneck models (Koh et al., 2020; Yuksekgonul et al., 2022), disentangled representation learning (Denton et al., 2017; Higgins et al., 2016), and network dissection (Bau et al., 2017; 2018).
+Machine learning models such as support vector machines (Hearst et al., 1998) and tree-based models (Song and Ying, 2015) were classical techniques for achieving model interpretability.
+In the deep learning era, typical research areas in this context include concept-bottleneck models (Koh et al., 2020; Yuksekgonul et al., 2022), disentangled representation learning (Denton et al., 2017; Higgins et al., 2016), and network dissection (Bau et al., 2017, 2018).
 Nevertheless, under the traditional deep learning setting, the usability of these techniques remains limited because of two major challenges.
@@
 By leveraging the common-sense knowledge embedded within them, foundation models can design interpretable architectures by providing cues that encourage creating and using the features or procedures within AI workflows.
-This is different from traditional deep learning pipelines where the deep models automatically discover the features during the training process, which may not end up with model components with clear meanings.
+This is different from traditional deep learning pipelines, where the deep models automatically discover the features during the training process, which may not end up with model components with clear meanings.
 Furthermore, LLMs can decompose complex tasks into simpler and collaborative sub-tasks, enhancing both the system’s interpretability and its overall performance.
-10.1 Designing Interpretable Network Architectures with LLMs
-Representative methods for developing interpretable deep architectures include Generalized Additive Models (GAMs) (Zhuang et al., 2021; Lou et al., 2012) and Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022).
+10.1. Designing Interpretable Network Architectures with LLMs
+Representative methods for developing interpretable deep architectures include Generalized Additive Models (GAMs) (Agarwal et al., 2021; Zhuang et al., 2021; Lou et al., 2012) and Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022; Shang et al., 2024).
 These models map inputs into a human-understandable latent space, and then apply a linear transformation from this space to the target label.
@@
 One promising strategy is to utilize LLMs to provide comprehensible concept candidates.
-Menon & Vondrick (2022) use human language as an internal representation for visual recognition, and create an interpretable concept bottleneck for downstream tasks.
+(Menon and Vondrick, 2022) use human language as an internal representation for visual recognition, and create an interpretable concept bottleneck for downstream tasks.
 By basing the decision on those comprehensible concepts, the model architecture itself is provided with better transparency.
-Similarly, a recent approach Labo (Yang et al., 2023b) constructs high-performance CBMs without manual concept annotations.
+Similarly, a recent approach Labo (Yang et al., 2023a) constructs high-performance CBMs without manual concept annotations.
 This method controls the concept selection in bottlenecks by generating candidates from the LLMs, which contain significant world knowledge (Petroni et al., 2019) that can be explored by prompting a string prefix.
@@
 Besides the concept-based models, another promising strategy is to employ LLMs to enhance the conventional architectures that are inherently interpretable, such as GAMs and Decision Trees (DTs).
-Singh et al.
-(2023) leverages the knowledge captured in LLMs to enhance GAMs and DTs, where LLMs are only involved during the augmented model training instead of the inference process.
+(Singh et al., 2023) leverages the knowledge captured in LLMs to enhance GAMs and DTs, where LLMs are only involved during the augmented model training instead of the inference process.
 For GAMs training, LLMs can provide decoupled embeddings for enhancement.
@@
 With the extra information from LLMs, augmented GAMs and DTs are capable of achieving better generalization performance compared with non-augmented ones.
-10.2 Designing Interpretable AI Workflows with LLM Agents
+10.2. Designing Interpretable AI Workflows with LLM Agents
 Traditional deep models are usually designed in an end-to-end manner.
 The internal workflows are not quite understandable to general users.
-By utilizing common-sense world knowledge, LLMs can break down complex problems into smaller ones and organize the workflows among them, leading to more interpretable design of AI systems (Feng et al., 2023).
-A recent example on interpretable AI workflow design comes from  Shen et al.
-(2024), where an LLM-powered agent leverages ChatGPT to integrate various off-the-shelf AI models (e.g., from Hugging Face (Jain, 2022)) to handle different downstream application tasks.
-In order to handle the complicated tasks in a transparent workflow, LLMs serve as a pivotal role in coordinating with external models with language mediums to harness their powers.
+By utilizing common-sense world knowledge, LLMs can break down complex problems into smaller ones and organize the workflows among them, leading to a more interpretable design of AI systems (Feng et al., 2023).
+A recent example of interpretable AI workflow design comes from  (Shen et al., 2024), where an LLM-powered agent leverages ChatGPT to integrate various off-the-shelf AI models (e.g., from Hugging Face (Jain, 2022)) to handle different downstream application tasks.
+In order to handle complicated tasks in a transparent workflow, LLMs play a pivotal role in coordinating external models with language mediums to harness their powers.
 By planning the target task, selecting candidate models, executing decomposed subtasks and summarizing responses, LLMs can help disassemble tasks based on user requests, and assign appropriate models to the tasks based on the model descriptions.
-Similarly, to transparentize the workflow, Liu et al.
-(2023d) introduces a task decomposer to analyze the user prompts and break it down into a number of subtasks for solving using LLMs.
+Similarly, to transparentize the workflow, (Liu et al., 2023b) introduces a task decomposer to analyze the user prompts and break it down into a number of subtasks for solving using LLMs.
 Each subtask is well managed and attributed with description, domain, inputs, and outputs.
 In this way, the AI systems are then capable of handling intricate user prompts with a step-by-step understandable workflow.
-Under the prompting paradigm, Khot et al.
-(2022) also employs LLMs to solve complex tasks by decomposition.
+Under the prompting paradigm, (Khot et al., 2022) also employs LLMs to solve complex tasks by decomposition.
 Drawing inspiration from software libraries where the workflows are trackable, the decomposer and shared subtasks are designed in a modular manner.
-One step further, Wang et al.
-(2024) introduces an interactive planning approach for complex tasks, which enhances the error correction on initial LLM-generated plans by integrating plan execution descriptions and providing self-explanation of the feedback.
-Such interactive nature enables better workflow transparency in long-term planning and multi-step reasoning task scenarios.
-10.3 Challenges
-10.3.1 Planning Feasibility in Complicated Scenarios
-Despite the task planning capability of LLMs, it is still challenging to be applied to certain scenarios in real-world applications due to the feasibility issues.
-One typical scenario is the few-shot planning cases (Guo et al., 2023), where acquiring large datasets for training is either impractical or cost-prohibitive thus making feasible planning about unseen cases from sparse exemplars extremely challenging.
+One step further, (Wang et al., 2024a) introduces an interactive planning approach for complex tasks, which enhances the error correction on initial LLM-generated plans by integrating plan execution descriptions and providing self-explanation of the feedback.
+Such an interactive nature enables better workflow transparency in long-term planning and multi-step reasoning task scenarios.
+10.3. Challenges
+10.3.1. Planning Feasibility in Complicated Scenarios
+Despite the task planning capability of LLMs, it is still challenging to apply to certain scenarios in real-world applications due to feasibility issues.
+One typical scenario is the few-shot planning cases (Guo et al., 2023), where acquiring large datasets for training is either impractical or cost-prohibitive, thus making feasible planning about unseen cases from sparse exemplars extremely challenging.
 To better assist the interpretable designs, LLM planning needs to generalize well without extensive supervision and is expected to have the ability to integrate information from prior experiences as well as knowledge.
@@
 Dynamic scenarios urgently and frequently involve high computational costs resulting from the iterated invocations of LLMs, and still face challenges in dealing with the limits of the context window and recovering from hallucinations on planning.
-10.3.2 Assistance Reliability with Knowledge Gaps
+10.3.2. Assistance Reliability with Knowledge Gaps
 LLMs exhibit remarkable proficiency in encapsulating real-world knowledge within their parameters, but they resort to hallucinations and biases with high confidence when certain knowledge is missing or unreliable.
-Although a growing number of techniques has been proposed, such as retrieval augmentation (Guu et al., 2020), searching integration (Nakano et al., 2021) and multi-LLM collaboration (Feng et al., 2023), to expand LLM knowledge, such discrepancy in knowledge may perpetually exist owing to the continuously evolving character of human understanding (Ji et al., 2023).
 As a result, a crucial research challenge keeps rising, i.e., how to effectively detect and mitigate the LLM knowledge gaps from humans when employing LLMs for designs.
 We will need further research on evaluating and developing robust LLM mechanisms to address the knowledge-gapping problems, with the goal of helping improve LLM reliability, reducing hallucinations and mitigating biases.
-Furthermore, the intersections between the knowledge gaps and the safety aspects are also of great challenges to be solved, which may pose some security concerns especially when using LLMs for downstream models or workflow designs.
-11 Emulating Humans with LLMs for XAI
-This section discusses how LLMs can be leveraged to serve XAI by playing the role of humans.
+Furthermore, the intersections between the knowledge gaps and the safety aspects are also a great challenge to be solved, which may pose some security concerns, especially when using LLMs for downstream models or workflow designs.
+11. Emulating Humans with LLMs for XAI
+This section discusses how LLMs can play the role of humans in serving XAI development.
 Building explainable models requires two main steps where humans are in the loop: (1) collecting a dataset with human-annotated rationales to train the models; (2) collecting human feedback on the quality of explanations produced by the models for evaluation.
@@
 LLMs emerge as a promising solution to this challenge, thanks to their capability to emulate human reasoning and produce responses that closely resemble human-generated content.
-In the following, we introduce the methods that demonstrate LLMs’ ability to generate human-like annotations and feedback, contributing to the creation of explainable models.
-11.1 Emulating Human Annotators for Training Explainable Models
-Incorporating human-understandable rationales into model development has shown its effectiveness in enhancing both the transparency and performance of the system for various NLP tasks, such as question answering (Li et al., 2018; Wu et al., 2020a), sentiment analysis (Du et al., 2019b; Antognini & Faltings, 2021), and common sense reasoning (Rajani et al., 2019; Camburu et al., 2021).
+In the following, we introduce the methods that demonstrate LLMs’ ability to generate human-like annotations and feedback for XAI.
+11.1. Emulating Human Annotators for Training Explainable Models
+Incorporating human-understandable rationales into model development has shown its effectiveness in enhancing both the transparency and performance of the system for various NLP tasks, such as question answering (Li et al., 2018; Wu et al., 2020a), sentiment analysis (Du et al., 2019b; Antognini and Faltings, 2021), and common sense reasoning (Rajani et al., 2019; Camburu et al., 2021).
 We use the term rationales to describe supportive evidence that justifies the connection between inputs and outputs (Gurrapu et al., 2023).
@@
 These studies found that advanced LLMs show comparable annotation qualities against average crowd human annotators on most tasks with a lower cost, pointing out the scalability of using machine-emulated annotators.
-Inspired by these works, some studies (Huang et al., 2023b; a) attempt to leverage advanced LLMs to collect rationales by applying the chain-of-thought technique.
+Inspired by these works, some studies (Huang et al., 2023b, a; Pang et al., 2024; Yang et al., 2024; Wu et al., 2024a; Zhang et al., 2024b) attempt to leverage advanced LLMs to collect rationales by applying the chain-of-thought prompting.
 Specifically, researchers provide several input-rationale-output demonstrations within the input text to prompt the LLMs to generate rationale and output for an unlabeled input instance.
@@
 Compared with other methods, this approach balances the annotation quality and the cost budget in developing explainable models by using LLM-emulated annotators.
-11.2 Emulating Human Feedback for Evaluating Explainable Models
+11.2. Emulating Human Feedback for Evaluating Explainable Models
 The explanations generated by the explainable models could be classified into two categories: extractive and abstractive (Gurrapu et al., 2023).
@@
 In contrast, abstractive explanations are generated in a free-form text manner, such as chain-of-thought (CoT) responses (Wei et al., 2022), offering a more nuanced interpretation.
-The quality of extractive explanations is typically assessed through their agreement with annotated rationales (DeYoung et al., 2020), such as accuracy, recall, and precision.
-However, evaluating abstractive explanations presents a significant challenge, as it is impractical to exhaustive all reasonable abstractive results comprehensively.
-To automatically assess abstractive explanations, early studies first collect some free-text rationales, and then apply LLMs to estimate the similarity between the explanation and the rationales (Cheng et al., 2023; Li et al., 2023b).
+The quality of extractive explanations is typically assessed through their agreement with human-annotated rationales (DeYoung et al., 2020), such as accuracy, recall, and precision.
+However, evaluating abstractive explanations presents a significant challenge, as it is impractical to exhaustively evaluate all reasonable abstractive results comprehensively.
+To automatically assess abstractive explanations, early studies first collect some free-text rationales, and then compute the cosine similarity between embeddings of explanations and rationales (Cheng et al., 2023; Li et al., 2023d).
 A higher similarity between the abstraction explanation and the annotated rationales indicates a more transparent model.
-Recently, some researchers directly use LLMs to check the rationality of the model explanations without referring to human-annotated rationales (Miao et al., 2023; Bills et al., 2023), emphasizing the potential of emulating human feedback with advanced LLMs.
-11.3 Challenges
-11.3.1 Uncontrollable Credibility of Emulation
-While LLMs can assist in rationale collection and explanation evaluation, their behaviors of collected results may not always match human annotators, primarily due to hallucinated responses in their unfamiliar domains (Ji et al., 2023).
-This issue leads to unreliable annotations or feedback, as LLMs confidently generate factually incorrect conclusions.
+Recently, some researchers have followed the LLM-as-a-judge framework to ask LLM to judge the model explanations without referring to any human-annotated rationales (Miao et al., 2023; Bills et al., 2023; Zheng et al., 2023; Saha et al., 2025; Wei et al., 2024b), emphasizing the potential of emulating human feedback with advanced LLMs.
+11.3. Challenges
+11.3.1. Uncontrollable Credibility of Emulation
+While LLMs could assist in rationale collection and explanation evaluation, their provided results can be biased and may not always match human annotators, due to hallucinated responses in their unfamiliar domains (Ji et al., 2023) and that LLMs are still immature (Harding et al., 2023).
 The quality of data gathered from this process is compromised, impacting the development of XAI systems.
@@
 These methods could enhance the reliability of LLM outputs, making them more comparable to human-generated content in the context of XAI development.
-11.3.2 Ethical Considerations in LLM Annotation
+11.3.2. Ethical Considerations in LLM Annotation
 When LLM annotators keep human annotators away from subjective scenarios, such as hate speech detection (Huang et al., 2023b), LLMs also have a chance to inject unethical opinions into their annotated datasets.
@@
 Ensuring LLM annotators follow ethical guidelines is worth further exploration.
-12 Discussion and Conclusion
-XAI research is undergoing a significant transformation and experiencing rapid expansion in the era of large models.
-In previous sections, we have introduced XAI methodologies with an emphasis on their usability.
-In this final section, we provide a high-level overview of the overarching challenges that persist in the field and suggest directions for future endeavors.
-  • Circumvent the interpretability-accuracy tradeoff.
-The advent of modern LLMs (e.g., ChatGPT) has a significant impact on this tradeoff.
-Traditionally, in many applications, people are willing to sacrifice a certain degree of performance for better transparency.
-A corresponding XAI strategy is to train and deploy an inherently interpretable model that mimics the black-box model (Che et al., 2016).
-However, applying this strategy to LLMs presents a challenge due to the difficulty in identifying an interpretable model that can match the performance levels of LLMs.
-This requires the creation of XAI strategies that can circumvent this tradeoff, where enhanced interpretability can contribute to improved accuracy.
-This is consistent with the goal of Usable XAI discussed in this paper.
-  • Data-driven AI vs.
-XAI.
-Data-driven AI refers to developing AI models that operate based on large volumes of training data.
-This approach often leads to “black-box” models, as it emphasizes results over the clarity of decision-making pathways.
-Currently, the development of XAI techniques lags behind the advancement of LLMs because the latter easily scale up with data-driven methods – they ingest gigantic amounts of texts from the Internet to train.
-However, we believe that XAI might still catch up because of several opportunities.
-(1) We may run out of data.
-It was predicted that “we will have exhausted the stock of high-quality language data before 2026” ***https://www.livemint.com/mint-top-newsletter/techtalk20102023.html.
-Should the accumulation of more data cease to yield substantial improvements, the focus might shift towards enhancing model interpretability to leverage existing data more effectively.
-(2) The model is relatively stable.
-As the Transformer architecture of LLMs is pretty mature and stable, it will attract more attention to interpret their inner workings.
-(3) Leveraging LLMs for XAI.
-The advancement of XAI research can be accelerated if it can properly use the knowledge and human-like capabilities of LLMs.
-  • The objective matters for explanation.
-During the transition from the era of classical machine learning (when SVMs and decision trees dominate) to the deep learning era (when convolutional and recurrent neural networks became popular), XAI techniques put a strong emphasis on achieving complete transparency within models, as if “anything less than fully transparent is not transparent”.
-However, as LLMs begin to match or even exceed human capabilities across various tasks, the importance of certain XAI problems shifts.
-For example, when Recurrent Neural Networks (RNNs) were widely adopted for text generation, we are interested in how the output is linguistically derived because RNNs often produce nonsensical sentences.
-Nowadays we are less interested in this for LLMs, as they are proficient in generating coherent text.
-Nevertheless, our focus may shift to explaining how LLMs construct output with factual information, as LLMs are prone to producing hallucinations.
-Similar observations exist in human cognition, which can be categorized into system-1 and system-2 styles: system-1 handles intuitive and unconscious tasks that are less explainable, while system-2 encompasses logical thinking, planning, and reasoning (Goyal & Bengio, 2022).
-Given the vast scale and complexity of LLMs, achieving absolute transparency across all aspects of these models appears increasingly unfeasible in the immediate future.
-Thus, prioritizing meaningful and feasible objectives of explanation, customized for specific tasks, becomes essential in enhancing the utility of AI systems in practical applications.
-  • Evaluation remains challenging for XAI in LLMs.
-Traditional XAI has developed a comprehensive taxonomy of explanation problems and formats, accompanied by clear definitions for each category (Han et al., 2022; Doshi-Velez & Kim, 2017; Rudin et al., 2022).
-However, the established taxonomy cannot be simply grafted into the study of LLM because of two reasons.
-First, certain XAI challenges lose their prominence in the context of LLMs, while some approaches become too complex for practical application.
-Second, while XAI is becoming a common pathway of solving problems for LLMs, the exploration of LLMs’ inner mechanics has branched into various directions.
-For instance, there has been a notable trend towards leveraging insights from human behavior and limitations to interpret LLMs, such as whether LLMs can lie (Azaria & Mitchell, 2023), can LLMs keep secrets (Mireshghallah et al., 2023), the impact of politeness in prompts on LLMs (Yin et al., 2024b), and even how they can be "hypnotized" (Li et al., 2023e).
-These diverse approaches have not converged to unified methodologies in interpreting LLM behaviors, which makes it challenging for evaluation.
-A potential risk is the resultant explanations might give users a false sense that they accurately understand the model, especially when users attempt to shoehorn certain human knowledge or concepts to explain LLMs (Schut et al., 2023).
-Conclusion.
-In this paper, we hope to guide readers through a crucial yet frequently underappreciated aspect of Explainable AI (XAI) – usability.
+12. Conclusion
+In this paper, we discuss a crucial yet frequently underappreciated aspect of Explainable AI (XAI) – usability.
 To this end, we present 10 strategies for advancing Usable XAI within the LLM paradigm, including (1) leveraging explanations to reciprocally enhance LLMs and general AI systems, and (2) enriching XAI approaches by integrating LLM capabilities.
@@
 We further discuss open challenges at the end of each strategy, suggesting directions for future work in this evolving area.
-Acknowledgement
-The work is, in part, supported by NSF (#IIS-2223768, #IIS-2223769, #IIS-2310261, #DRL-2101104).
-The views and conclusions in this paper are those of the authors and should not be interpreted as representing any funding agencies.